### v1

# TermSight: Making Service Contracts Approachable

Abstract.
Terms of Service (ToS) are ubiquitous, legally binding contracts that govern consumers’ digital interactions. However, ToS are not designed to be read: they are filled with pages of ambiguous and complex legal terminology that burden potential users. We introduce TermSight, an intelligent reading interface designed to make ToS more approachable. TermSight offers visual summaries that highlight the relevance and power balance of information in a ToS. TermSight also categorizes and simplifies information within the ToS into concise plain-language summaries. To aid in reading the original text, TermSight offers contextualized definitions and scenarios for unfamiliar phrases. Our within-subjects evaluation of TermSight (N=20) revealed that TermSight significantly reduced the difficulty of reading ToS and increased participants’ willingness to do so. We also observed emerging strategies that participants took when interacting with AI-powered features that highlight the diverse ways that TermSight assisted ToS reading.
large language models, legal documents, reading support tools
copyright: acmlicensedjournalyear: 2025ccs: Human-centered computing Interactive systems and tools
Figure 1. The TermSight system provides guidance and reading support at multiple levels of ToS reading. At the contract level, TermSight visualizes the relevance and power balance of content in each policy (1). At the document level, TermSight chunks, summarizes, and categorizes content into 1-sentence plain-language summaries (Summary Snippets) highlighted with colors that reflect power and relevance (2). Readers can click the Summary Snippet (2) or Highlight Bar (3) to navigate between the summary snippets on the left and the original text on the right. At the phrase level, TermSight offers phrase definitions and scenarios for unfamiliar and vague phrases (4).
1. Introduction
Terms of Service (ToS)—the ubiquitous, rarely-read but legally binding contracts that govern the use of digital services—are widely incomprehensible to potential consumers. Despite consumers’ self-reported concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007), and their eventual regret when they encounter issues related to this information (e.g., the selling of their data)  (Good et al., 2005; Obar and Oeldorf-Hirsch, 2020), almost no one reads ToS content fully  (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020). There is good reason not to do so: ToS are not designed to be read. ToS are filled with legal jargon that can derail general-audience (also referred to as ‘lay-audience’) readers with confusing or under-defined terms (e.g., a ToS stating the service will collect certain ‘information’)  (Ibdah et al., 2021; Tang et al., 2021; Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). ToS also contain more than one policy (e.g., a privacy policy) that in turn each contain reams of boilerplate or semi-boilerplate text. Within these pages of text, there might be information relevant to a reader (e.g., an arbitration clause or copyright policy) but such information is lost among the many pages of text  (Srinath et al., 2021; Obar and Oeldorf-Hirsch, 2020). Legal scholars have argued that these ambiguous, jargoned, and exceptionally long contracts are designed not to cultivate a shared understanding, but to mitigate liability and establish unilateral control (Kar and Radin, 2019; Benoliel and Becher, 2019; Karanicolas, 2021; Palka, 2023).
Summarizing and simplifying text is a first step in making ToS more approachable, but resolves only part of the challenge these contracts pose. Prior interfaces have leveraged crowdsourcing (TOS, 2024; pri, 2024) and, more recently, AI-based approaches like large language models (LLMs) to extract and process text automatically from ToS (Pan et al., 2023). However, these approaches primarily focus on presenting information from pre-defined categories (e.g., user privacy) or on single policies within a ToS (e.g., a privacy policy). In reality, ToS contain many policies that cover different information (§3.3.1), and what information a reader needs is dependent on their planned usage of the site (e.g., a seller vs. buyer on an e-commerce site).
We envision a new reading experience for ToS. Our goal is to enable readers to find information important to them anywhere within a ToS and scaffold their ability to understand and contextualize this information within their potential use of a service. We start by developing a more nuanced understanding of readers’ information needs and challenges when reading ToS through a formative observational study (N=20). We found that participants cared about different information based on their envisioned usage of a service (e.g., a content creator will have different needs than a passive consumer on a social media website) and their own personal values (e.g., people had different levels of comfort when it came to privacy and data sharing). When participants sought to resolve these information needs, they encountered barriers at every level of granularity within a ToS. Within the full ToS contract, participants struggled to navigate nested policies and decide which policies to read. Within a specific policy of the ToS (e.g., the privacy policy), participants struggled to surface and interpret relevant information from extended, visually dense, and difficult text. When reading the text itself, participants struggled to interpret jargon and phrases that were ambiguous or vague. Interestingly, existing affordances (e.g., policy names, section headers with table of contents, and summaries) were often not helpful and at times misleading. For example, participants often found that policy names and section headers were unclear and that overview summaries used deceptive friendly language to hide unfavorable clauses.
We address the specific challenges surfaced by our formative work with a new intelligent reading interface, TermSight. TermSight provides multi-level reading guidance and support through contract-level overviews (Power Meter), document-level plain-language summaries (Summary Snippets), and phrase-level definitions with scenarios of potential implications (Phrase Scope). To evaluate TermSight, we conducted a counterbalanced within-subject study (N=20) comparing TermSight and a baseline HTML reader during a timed reading task. Participants found reading ToS to be significantly easier with TermSight and were more willing to read and spend time on ToS. Participants reported that features of TermSight provided guidance and reading support at all granularities of ToS, resolving challenges surfaced in our formative study. We also observed interesting and unexpected user behaviors centered around trade-offs between TermSight’s AI-generated features.
In this paper, we make the following contributions:
  1. (1)
    We characterize the challenges readers face when trying to make sense of ToS contracts. Readers face challenges at all granularities of a ToS, from determining which policies to read to being able to resolve individual ambiguous terms within the text.
  2. (2)
    We design and develop an intelligent reading interface, TermSight, for helping readers make sense of ToS contracts at the contract, document, and phrase levels.
  3. (3)
    We collect empirical evidence from our user study (N=20) demonstrating the potential value of the interaction affordances of TermSight.
2. Background and Related Work
2.1. Service Contracts and The Duty to Read
A Terms of Service (ToS) is a type of standard form contract consisting of a body of policies and conditions outlined by a service provider, dictating the rules and expectations for both the service provider and the consumer. These contracts have evolved to be more legally comprehensive compared to earlier ToS  (Amos et al., 2021), encompassing a wide range of provisions such as copyright, privacy, returns, acceptable use, and service-specific terms. By clicking agree to a ToS, a legally binding contract is formed  (Dickens, 2007). U.S. contract law is premised on the doctrine of the ‘duty to read’ where contracting parties are legally responsible for reading and understanding the contract  (Benoliel and Becher, 2019). However, the ‘duty to read’ was established in the era when contracts were short and based on mutual negotiation  (Kar and Radin, 2019). Modern ToS often have important terms buried within a large body of text and documents  (Srinath et al., 2021). The language used to write ToS can also require advanced reading capability  (Srinath et al., 2021; Amos et al., 2021) and contain unfamiliar (Ibdah et al., 2021; Tang et al., 2021) or ambiguous terminology (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). Due to the cognitive load associated with reading ToS, most consumers do not read modern ToS (Obar and Oeldorf-Hirsch, 2020; Bakos et al., 2014), despite the binding nature of ToS and consumers’ self-reports indicating concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007). ToS can also vary in content (Liu et al., 2021a; Fiesler et al., 2016) and structure (Ibdah et al., 2021) across services, leading to false consumer expectations. This widespread disengagement highlights what Kar and Radin described as a ”paradigm slip” in contract law, where agreements once premised on shared understanding and mutual consent are now reduced to unilateral boilerplate ”pseudo-contracts”, eroding fundamental principles of contract law and personal autonomy  (Kar and Radin, 2019; Palka, 2023; Benoliel and Becher, 2019). Our formative study complements prior findings by aggregating and uncovering consumers’ varying information needs and additional challenges encountered when navigating and reading policies within a ToS.
2.2. Augmenting Terms of Service
There are two major threads of work on improving the readability of ToS that we cover below: changing how ToS are designed or presented (§2.2.1) and providing summaries of ToS (§2.2.2).
2.2.1. Stylistic Manipulations
To help readers better understand the implications of using a service, prior work has investigated design recommendations for service providers to make documents within ToS more accessible  (Kay and Terry, 2010; Habib et al., 2021; Taber et al., 2020). Kay and Terry (2010) proposed Textured Agreements that use typographic manipulation, pull quotes, vignettes, and iconic symbols to improve reader attention and comprehension on privacy policy compared to plain text. Habib et al. (2021) investigated how icons and linked text could be designed to better convey privacy choices. Taber et al. (2020) proposed crowdsourced sentiment highlighting of sentences in ToS, providing visual cues to guide readers toward sentences in ToS with negative sentiment. Design recommendations on styling user agreements have also been proposed  (Kay and Terry, 2010; Habib et al., 2021). While helpful in navigating contracts, these design changes do not simplify the content itself and do not take into consideration the fact that readers have different information needs when reading a ToS.
2.2.2. Summarization
Work has also proposed helping consumers gain an overview of specific policies within a ToS, often the privacy policy (PP), without requiring reading the original document. Platform for Privacy Preferences (P3P) is an early attempt to standardize privacy policies by allowing service providers to submit privacy policies in a machine-readable format  (Reagle and Cranor, 1999). Tools such as Privacy Bird build atop the P3P format to provide warnings when a site’s PP does not match consumers’ privacy preferences  (Cranor et al., 2006). However, P3P is rarely adopted by service providers  (Cranor et al., 2008) and no longer supported  (Microsoft, 2016). Work has also proposed the use of a single-page summary for end-user license agreements (EULAs)  (Good et al., 2005, 2007), comic-based summaries for privacy policies  (Tabassum et al., 2018), and Privacy Policy Nutrition Labels (PPNL) that label privacy policies in a table format  (Kelley et al., 2009; Reinhardt et al., 2021) to help users gain an overview of a policy.
Many earlier tools required buy-in from service providers to design better ToS. Other work, though, has focused on providing an overview summary of a ToS automatically or via crowdsourcing. For example, ToS;DR leverages volunteers to label and summarize terms in the privacy policy and the home page of ToS  (TOS, 2024). Works like CLAUDETTE automatically detect a predefined set of potentially unfair clauses from ToS  (Lippi et al., 2019; Guarino et al., 2021; Braun and Matthes, 2021; Braun et al., 2017). PrivacyCheck (Zaeem et al., 2018; Nokhbeh Zaeem et al., 2020, 2022), PrivacyGuide (Tesfay et al., 2018), and Polisis  (Harkous et al., 2018) extract terms relevant to a list of privacy-related questions and criteria and generate a grading for each. Sathyendra et al. extract opt-out choices from privacy policies  (Sathyendra et al., 2017; Bannihatti Kumar et al., 2020). Relatedly, some works evaluate the completeness of privacy policies  (Costante et al., 2012) and the GDPR compliance of extracted clauses (Liu et al., 2021b). More recently,  Pan et al. (2023) investigated generating privacy nutrition labels from a privacy policy. Work has also attempted to generate an aggregated summary based on the extracted information from a privacy policy  (Tomuro et al., 2016; Keymanesh et al., 2020). TermSight extends this rich prior work by combining the benefit of plain language summaries with interactive affordances that enable readers to navigate the original ToS text, helping readers navigate and read the original body of policies within ToS beyond privacy policies.
2.3. Intelligent Reading Interfaces
TermSight takes inspiration from a rich body of work on intelligent reading interfaces that support readers in navigating documents and clarifying information. Prior work has explored helping readers navigate documents and locate relevant information in scholarly documents. For example, work has highlighted key information in research papers, such as objectives, results, and methods, to support skimming  (Fok et al., 2022). Also in research papers, ScholarPhi identifies position-sensitive definitions of terms and symbols, enabling readers to track their meaning across contexts  (Head et al., 2021). Talk to Papers allows readers to ask questions about research papers directly, while Qlarify supports clarification by letting readers expand text spans in paper abstracts to generate explanations  (Zhao and Lee, 2020; Fok et al., 2024a). In medical research papers, Paper Plain provides a list of guiding questions to help readers navigate important information in a medical paper (August et al., 2023). In the business domain, Marco allows readers to search and ask questions across collections of business documents  (Fok et al., 2024b). In the news domain, Chen et al. (2023) introduced Marvista, a reading interface that identifies the most summative portions of news articles based on a reader’s time constraints or topic in the news article.
Compared to prior reading interfaces that have focused on research papers, news articles, or business documents, TermSight focuses on helping general audience readers make sense of legal documents, namely, ToS. Legal documents, and specifically ToS, present challenges different from prior reading contexts. In contrast to other knowledge-intensive settings (e.g., scientific or medical text), general audience readers in the legal setting are faced with language not designed to communicate information. Readers then struggle to (1) determine what text to read within and across policies, (2) define and contextualize vague language, and (3) navigate existing affordances that are themselves ambiguous or incomplete. Prior intelligent reading interfaces and context often assume that there is a single document (August et al., 2023; Fok et al., 2022; Chen et al., 2023), readers have consistent information needs  (August et al., 2023), readers know what they need to read or ask  (Fok et al., 2022; Zhao and Lee, 2020; Fok et al., 2024b), or readers can make sense of term definitions  (Head et al., 2021; August et al., 2023; Fok et al., 2024a). These existing affordances, while powerful, do not fully support readers’ needs and challenges encountered when reading digital contracts such as ToS. TermSight takes inspiration from prior works to explore the design of reading interfaces to help readers surface and interpret relevant information from legal documents.
3. Formative Study
To better understand the challenges readers face when trying to make sense of the Term of Service and their information needs, we conducted a formative study to answer the following research questions:
RQ1: What information do readers want from a ToS?
RQ2: What are the challenges readers encounter when trying to get this information from ToS?
3.1. Methodology
3.1.1. Study Procedure:
The entire study took 40 minutes and participants were compensated 10\$. The study was approved by the Institutional Review Board (IRB) office at our organization. Participants initially completed a demographics form and reported their past experiences interacting with ToS. Following the form, participants were asked to imagine as if they were a first-time user of the assigned service registering for the service and were provided 15 minutes to go through the ToS. Participants were free to navigate to different policies within the ToS. While reading, participants were instructed to speak aloud any challenges they faced or thoughts about the ToS. After reading the ToS, we conducted semi-structured interviews with participants about their experiences and challenges with reading ToS. Lastly, we asked the participants to envision how computational supports could aid their reading. All interview sessions were transcribed. The semi-structured interview questions are listed in Appendix E.
3.1.2. Analysis
We conducted a reflexive thematic analysis on the interview transcripts to identify common challenges and themes. We followed the six phases of reflexive thematic analysis suggested by Braun and Clarke  (Braun and Clarke, 2006). The lead researcher thoroughly explored the data, noted interesting features, systematically coded the data, and iteratively compared the codes to generate the initial themes. Then, the lead researcher discussed with other members of the research team to make sure the themes fit the data and further developed the themes.
3.1.3. Materials
Participants were randomly assigned one of the 10 ToS in the study. We selected ToS from 10 services spanning social media, e-commerce, video streaming, and internet services (Social Media: Reddit, Facebook, Instagram, Twitter; E-commerce: Amazon, eBay; Video: Youtube, Netflix; General: Google, Yahoo). These 10 services were selected from the top 20 visited sites on semrush.com (2024/05) where we selected the top 4 social media services, top 2 E-commerce services, top 2 video platforms, and top 2 internet services. We over-sampled social media services because they are the most represented service category in the top 10 visited services. In the study, participants were first directed to the main ToS and were allowed to navigate across any other policies.
3.1.4. Participants
We recruited 20 participants through Prolific. Each participant was randomly assigned one of the 10 selected ToS. All participants were over 18, fluent in English, and located in the US. 14 participants self-identified as female and 6 as male. 3 participants were between the ages of 18-25, 9 were between 26-35, 4 were between 36-45, 3 were between 46-55, and 1 was 67. 3 participants had never read or skimmed a ToS before and 17 participants had read or skimmed at least 1 ToS before. We did not require participants to have prior experience in reading the ToS. All the participants either had not read or did not remember reading the ToS they were assigned.
3.2. RQ1: What information do readers want from a ToS?
3.2.1. Participants have diverse information needs based on usage of the service and personal value
All 20 participants described the need to know what control and rights the service had over the user, what the service was allowed to do, and how the user could be negatively impacted. Additionally, 12 participants highlighted the need to know their (as opposed to the service’s) rights and control. Furthermore, 4 participants expressed wanting to know what they were allowed or not allowed to do on the platform. Participants also noted more specific and diverse information that they cared about such as data collection and usage (11), intellectual property right over user content (8), purchasing and returns (7), account deletion by the service (5), arbitration and liabilities (3), content moderation (2), and exposure to offensive content or misinformation (1).
10 participants explicitly mentioned how their information needs and priorities depended on usage and personal values. For example, P1 described that “I don’t think any information is irrelevant. I just think some parts are more of a priority of knowing. If I happen to become a person who wants to create graphic designs, I would definitely utilize the Copyright portion of the ToS to ensure that I’m being compliant with [the platform’s] expectations.” On the other hand, P3, with a background in software, raised privacy concerns when using services for professional communication: “For Microsoft Teams and Gmail, I was using them for a collaborative project with a small company. I wanted to ensure that whatever data or information was shared would remain proprietary to us.”
3.3. RQ2: What are the challenges readers encounter?
Our findings revealed that participants encountered challenges at the contract (§3.3.1), document (§3.3.2), and phrase (§3.3.4) level. In addition, participants highlighted how existing navigational and reading affordances were often ineffective, misleading, and failed to support meaningful comprehension (§3.3.3).
3.3.1. Contract Level: Participants struggled to navigate nested policies and decide which policy to read.
11 participants mentioned that it’s unclear to them what sub-policies are included when they are navigating the ToS because sub-policies are often hyperlinked throughout the documents. There often lack a centralized list of what policies are included as part of ToS. 7 participants described the challenge of navigating across these nested policies as getting ‘lost in a whole tree of information’ (P18).
All 20 participants described relying on policy names to help decide whether to explore a policy. Yet, 17 participants described that the names of the policies alone did not provide a clear mental model of what was in the sub-policies and whether there was important information they should know about. For example, P14 mentioned that: “When it’s hyperlinked like ‘our rules and policies’ and ‘privacy policy’, it’s difficult to determine if I need to know something from the link.” 8 participants explicitly expressed their reluctance to click into the linked polices due to the lack of awareness of relevant information and the challenge of finding relevant information from a hyperlinked document.
3.3.2. Document Level: Participants lacked guidance and struggled to surface relevant information from extended, visually dense, and obfuscating text within a policy.
17 out of 20 participants described their reading behavior as skimming the ToS to see if any information caught their eye. 6 participants noted that they were not sure what they should look for in the ToS and might have unintentionally missed important information. P12 further commented how it’s not always obvious if the information was relevant at first sight: “A lot of it seemed irrelevant until I started reading in a little bit more.”
18 participants highlighted the overwhelming length of the document, especially when considering the nested policies. These participants expressed concerns about missing important terms buried in the text. For example, P7 pointed out that “When all the important information gets hidden behind walls of text, it doesn’t really incentivize anyone to read it, and it gets hidden.” Reflecting P7’s comment, 13 participants described the text as visually dense and difficult to parse when skimming. These participants further noted how important information can be positioned at the very bottom of an extended policy that was hard to discover. For example, P10 noticed how information about returns and cancellations of an E-commerce platform was positioned under the 13th section of the ToS with the title ‘Additional Terms’.
Additionally, 3 participants described how the use of friendly language actually obfuscated the ToS and made skimming more challenging. For example, P15 noticed language such as ‘We only use data to make [the service] a better place’ when reading a privacy policy. Yet, she later found out how user information can be used for targeted ads. As a result, P15 described that it’s easy for lay people to misinterpret the language and miss important terms: “If someone’s just skimming the terms of service and is not skeptical at all. Then, you could read over that and think that means they’re not using your data for targeted ads or stuff.”
3.3.3. Document Level: Participants found existing affordances to be ineffective and misleading
Some of the ToS participants read contained navigational and reading affordances like an interactive table of contents with section headers and overview summaries. However, 12 participants noted that, similar to policy names, section headings within policies were vague and not descriptive of the actual content. As a result, it was unclear what information was contained in a section by only looking at the section header. For example, when skimming the ToS, P10 skipped the ‘Content’ section but later realized that the section was about intellectual property rights over user-generated content: “It just says ‘content’ which is pretty vague when they’re talking about getting exclusive intellectual rights. I feel like it’s misleading. So people would skip right past it. I did the first time.”
In addition to ineffective navigational affordances such as vague policy names (17/20) and section headers (12/20), 3 participants accessed a sub-policy with a summary at the top. All three participants (P4, P14, P15) described using the summaries to gain an overview at the start. Yet, after reading the policy, participants noticed that the summary was missing key details: “The summary is kind of misleading. They gave you a more palatable version of [ToS]. If you scroll down, it will say we can terminate your account for no reason. They don’t say that in the summary” (P4). They also expressed concern that readers might only glance at the summary and mistakenly think it covers the entire policy.
3.3.4. Phrase Level: Participants struggled to contextualize unfamiliar or ambiguous terminology
When reading individual sections in a policy, 12 participants pointed out unfamiliar legal terminology (e.g., Arbitration, Indemnity, or class lawsuit). These participants described how the terminology hindered their understanding and motivation to read the section. For example, P8 described skipping sections that were full of unfamiliar terms. Additionally, participants noted how challenging it was to interpret the meaning of a word or phrase in the context of their situation: “I can recognize these are legal terms, but I don’t necessarily know what that means for me” (P10). P6 also described how it’s hard to envision scenarios in which things go wrong beforehand from the ToS.
8 participants noticed ambiguous or vague terminologies that made it challenging for them to understand the implications of signing. For example, P15 mentioned that “I really think the big difficulty is the vagueness of the language and the constancy of exceptions that are vague which gives them a lot more leeway.” More specifically, 5 participants noticed vague terms around the service’s information collection and sharing practices such as: ‘we share you data with 3rd parties’, ‘track information about your interaction’, and ‘retain certain information about you’. 4 participants noted vague terms around how the service can use user-generated content. For example, P3 and P14 pointed out that it’s unclear what “use, distribute, modify, and copy user content” implicates and what are examples of what the service can do with user-generated content.
3.4. Design Guidelines:
We propose four design guidelines based on our observations that readers faced interrelated barriers at all levels of ToS reading:
DG1: Contract level: Help readers form mental models of each policy of a ToS to determine important documents to interrogate.
DG2: Document level: Help readers identify potentially relevant information within a policy.
DG3: Phrase level: Help readers read and interpret original text with technical language, jargon, and vague phrases.
DG4: All summaries should be tightly coupled with the text used to generate them. When reading any generated overviews, readers should be able to retrieve the original document text in at most one click.
4. The TermSight System:
We reify these design guidelines in TermSight, an augmented reading interface that supports user sense-making of ToS with features at each level of granularity of a ToS:
Power Meter: A contract-level visualization of the relevance of information within a document and the distribution of power between the user and the platform that information represents.
Summary Snippets: 1-sentence plain language summaries of information chunks at the document-level. Snippets are rendered with color and saturation that represent power and relevance to the user, similar to the Power Meter.
Phrase Scope: In-situ phrase-level definitions and scenarios via a tooltip that contextualizes the meaning of the phrase while allowing users to ask clarification questions.
We adopted an iterative design approach in developing TermSight. 8 participants evaluated an early prototype of TermSight, and their feedback informed the final design. Additional details about the iterative design study, results, and design changes can be found in Appendix  A.
4.1. Design Abstractions:
The design of TermSight is grounded in two design abstractions: 1) Information Snippets, and 2) Power and Relevance classification. These abstractions serve as foundational building blocks for TermSight’s features that support contract and document-level sensemaking as illustrated in Figure  6.
4.1.1. Information Snippets
Our formative study reveals that the information readers cared about in ToS is often buried within extended and visually dense text. Consequently, prior works that support document understanding through document-level summaries (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023) can fall short in helping readers surface pieces of information buried in large sections. To address this limitation, we use Information Snippets as a basic building block for TermSight. An information snippet is a continuous span of original text that shares the same topic and can be summarized in a single sentence of up to 12 words (detailed in Section 5.2). These information snippets form the basis of the Summary Snippets (§4.3.1). Information Snippets are non-overlapping and together reconstruct the original document’s content (Examples in Figure  2). Figure  6 illustrates the process of how TermSight divides a paragraph into five Information Snippets, each paired with its Summary Snippet. The idea of designing interactions and sense-making around Information Snippets rather than entire sections was inspired by prior works that modularize and ‘objectify’ tools (Bier et al., 1993; Bederson et al., 1996), attributes (Xia et al., 2016), and AI agent’s memory (Huang et al., 2023) to enable interaction flexibility, interaction specificity, and direct manipulation.
Figure 2. Design conceptualization of Power and Relevance: Power refers to the degree to which a snippet grants control to the service provider or the user. Relevance refers to whether or not the snippet is relevant to the user’s intended usage of the service and personal value. Examples of Information Snippets, TermSight classified under each category of power, are shown.
4.1.2. Power and Relevance Classification
We define Power and Relevance as two qualities of an information—and thus a summary—snippet. Power is the degree to which a snippet’s text grants control to the Service Provider or the User (Categories: Service, Neutral, or User). Relevance is whether the snippet’s text is directly relevant to the user’s intended usage of the service or their values (Categories: High, Low). For example, a snippet related to selling would have low relevance for a user more focused on buying. As shown in Figure  2, we visually encode Power and Relevance in TermSight with a combination of hue (Red, Yellow, Green) and saturation (High, Low).
Figure 3. Power Meter: a visualization of the distribution of the power and relevance of Information Snippets within a document (a). Users can gain an overview of the policies within the ToS (b). When hovering over the Power Meter, the list of Summary Snippets of the corresponding policy is shown (c). Clicking the blue backquote would allow users to view the referenced Information Snippet (d).
4.2. Power Meter: Policy level visualization of power and relevance
While prior works have focused on a single policy such as the Privacy Policy  (Nokhbeh Zaeem et al., 2022; Tesfay et al., 2018) or the main ToS page (Lippi et al., 2019; Guarino et al., 2021), Terms of Services include many policies (e.g., content policy, community guidelines, return policy, and more). To help users form mental models of each sub-policy, TermSight provides information scent and preview to help users gain an overview of the policies within ToS to decide which policies to read (DG1).
4.2.1. Providing information scent of a policy
In TermSight, the policies that are part of the ToS agreement are placed in the top navigation panel where users can navigate to different policies. Each policy in the top navigation panel or hyperlinked within the document is accompanied by a Power Meter: a horizontal bar visualization representing the distribution of Information Snippets in a document, with colors denoting power and relevance. Shown in Figure 3, Power Meter provides information scent (Pirolli and Card, 1999) of the power distribution of Information Snippets within the policy and the fraction of Information Snippets that are potentially relevant to the user.
4.2.2. Providing a preview of the document
Hovering over a Power Meter reveals a preview popup containing a list of Summary Snippets, allowing users to skim and gain an overview of a policy’s Information Snippets without having to be redirected to a new page (Figure 3c). Each Summary Snippet in the popup includes a clickable quote icon that allows users to view the referenced Information Snippet (DG4).
4.3. Summary Snippets: Plain language summaries of discrete information snippets
To help users surface and interrogate snippets of information buried in a wall of text (DG2) and make original text more accessible (DG4), TermSight provides Summary Snippets accompanied with Highlight Bars for seamless navigation between the summary and original text.
4.3.1. Summary Snippets
Unlike prior methods that rely on document-level overviews (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023), TermSight supports interaction and sensemaking for finer-grained snippets of information within dense sections. TermSight features Summary Snippets - one-sentence plain language summaries of Information Snippets (§4.1.1). Each Summary Snippet is rendered with color based on the Power and Relevance of the corresponding Information Snippet (§4.1.2). To enable users to interrogate original text and verify AI-generated summaries (DG4), clicking on a Summary Snippet jumps a user to the corresponding Information Snippet in the document (Figure  4a).
4.3.2. Highlight Bar
To help users surface and identify Information Snippets while reading or skimming the original text (DG2), TermSight introduces Highlight Bars. Highlight bars are positioned to the left of each Information Snippet in the original text, visually segmenting dense sections into Information Snippets. For paragraphs containing multiple Information Snippets, additional line breaks are added after each snippet to add visual separation. Similar to the Summary Snippets, Highlight Bars are color-coded based on the Power and Relevance of the corresponding Information Snippet (§4.1.2). To support navigation from the original document to the summary, users can click on a Highlight Bar to jump to the corresponding Summary Snippet, as shown in Figure  4b.
Figure 4. Viewport Scrolling: (a) By clicking the Summary Snippet, the viewport of the original passage will scroll to the referenced Information Snippet; (b) By clicking the Highlight bar, the left viewport will scroll to the corresponding Summary Snippet.
Figure 5. Phrase Scope: A tooltip that provides term definition and scenario that contextualizes the meaning of the phrase while allowing users to ask clarification questions. Users can hover over the references to view the referenced text in the ToS.
4.4. Phrase Scope: Context-dependent definition, personalized scenario, and clarification question
To help users read original text with unfamiliar or vague terminology (DG3), TermSight extends prior work focusing on offering dictionary definitions (e.g.,  (August et al., 2023)) by introducing Phrase Scope: a set of features that support the identification and interpretation of unfamiliar or vague phrases (Figure  5). First, TermSight identifies words and phrases that might be unfamiliar or vague to general audience readers and underlines these phrases in blue. This aims to provide a starting point for users to explore potentially unfamiliar phrases without relying on users to identify them. Clicking on the identified phrase opens Phrase Scope, a tooltip with the phrase’s definition in context, a personalized scenario to envision potential implications, and an option to ask clarification questions (examples in Table  2). In addition to the suggested phrases, TermSight allows users to highlight any arbitrary span of text to request Phrase Scope.
5. Implementation details
TermSight is rendered as a web interface implemented with Next.js. Below, we explain the input, output, processing, and overall performance of TermSight.
5.1. System Input and Pre-processing:
TermSight assumes a set of HTML or markdown source files representing the policies included in the Terms of Service (ToS). For our user study, the research team manually copied the text from documents in ToS on the internet into markdown files with annotated hyperlinks and headers. We assumed clean source files because the focus of TermSight is not on developing web scraping and cleaning technologies but rather on investigating meaningful navigational and reading affordances for ToS. Each document is segmented into text ”chunks”, composed of one or more paragraphs, which are further vectorized and stored in a vector database (detailed in Appendix  B.1).
For the TermSight features that depend on user preferences, such as classifying the relevance of information snippets (§5.2) and generating personalized scenarios (§5.3.3), the system uses a text-based persona that includes users’ intended usage of the service (e.g., content consumers vs. content creators) and their values or concerns (e.g., privacy, copyright, etc.). The user personas used for the study are further explained in Section  6.1.2 and displayed in Appendix  F.1.
Figure 6. A flowchart of the implementation for (1) obtaining Summary Snippets and Information Snippets and (2) classifying Power and Relevance for each Information Snippet.
5.2. Obtaining Summary Snippets and Information Snippets:
For each chunk of text obtained from the document pre-processing pipeline, we prompted GPT-4o to generate a list of 1-sentence summaries (Summary Snippets) each referenced to a span of the input text chunk (Information Snippet) as illustrated in Figure  6. While the length of the referenced text for each summary is unrestricted, each summary is constrained to a maximum of 12 words based on prior works (Cutrell and Guan, 2007; Sweeney and Crestani, 2006) and prompt engineering. Then, we prompted GPT-4o to classify each Information Snippet along two dimensions: Power and Relevance (to user persona). We detail our prompt design in Appendix  B.3.
5.3. Phrase Scope
Below, we detail the implementation of each feature in Phrase Scope.
5.3.1. Identifying unfamiliar or vague phrases
Prior works have demonstrated the capabilities of LLM in personalized jargon identification (Guo et al., 2024) and detecting phrases that can potentially be elaborated (Fok et al., 2024a). Inspired by this work, we employed GPT-4o with two few-shot prompts to identify potentially unfamiliar and vague phrases within the document chunks produced by the pre-processing pipeline (Prompts in Appendix  B.4.1). The input chunks are also referred to and used as the ”phrase context” for each identified phrase.
5.3.2. Generating phrase definitions or responses to user questions
To generate in-context definitions for the identified phrases, we used a retrieval-augmented question answering approach  (Fok et al., 2024a). We first query the vectorized document chunks (§5.1) to retrieve chunks that might define the phrase. The database query is framed as a question: “What does {input phrase} refer to in the sentence: {phrase context}. Here, phrase context refers to the chunk of text containing the input phrase as defined in §5.3.1. Then, the query is embedded using OpenAI’s text-embedding-3-small model, and the top 15 chunks from the vector database are retrieved based on cosine similarity. These chunks are collectively referred to as ”retrieved chunks”. Using the phrase context and the retrieved chunks, we prompted GPT-4o to generate an in-context definition with references to the retrieved chunks. When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked. We detail our prompts in Appendix  B.4.2.
5.3.3. Generating scenarios for phrases
Scenarios intend to complement phrase definitions, helping users interpret and envision the potential implications of a phrase in context. In addition, customizing scenarios based on user personas intends to ensure that the generated examples are relevant to the user’s intended usage of the service (e.g., content creator or consumer) and personal value (e.g., privacy, copyright). Building on prior works that demonstrated the effectiveness of large language models (LLMs) in generating unintended consequences (Wang et al., 2024) and stories to teach legal concepts (Jiang et al., 2024), we employed GPT-4o with zero-shot prompting (Figure  27) to generate the scenarios with the phrase context, phrase definition, and user persona.
5.4. Evaluation of System Output
Before reporting on our user study, we randomly sampled and conducted a manual evaluation of TermSight’s core outputs. Our goal was not to assess advances in system performance, but to verify that the system can produce meaningful outputs that could support ToS reading. Across 116 sampled information snippets, we found 5 imperfect classifications of power or relevance due to the lack of full context in the input information snippet. Similarly, we reviewed 113 generated definitions and scenarios each. One scenario was found to be factually incorrect due to hallucination. All the definitions were correct except for 4 overly general definitions for service-specific phrases not explicitly defined anywhere in the ToS. More details are provided in Appendix C.
6. User Study
We conducted a counterbalanced within-subject experiment with 20 participants to evaluate TermSight. We asked the following research questions:
RQ1: How did participants perceive TermSight and its features?
RQ2: How did participants read with TermSight and its features?
RQ3: How did TermSight influence comprehension and recall?
6.1. Study Design
6.1.1. Study Conditions/Interface Variants
To evaluate TermSight, we conducted a within-subject experiment where each participant used both TermSight and a baseline interface to read 2 services’ ToS contracts, one with each interface variant. The order of the interface variant and service type were counterbalanced to reduce ordering effects. The baseline interface intended to mimic a standard ToS. The baseline interface had the same layout as TermSight without the Power Meter, Summary Snippet, Highlight Bar, and Phrase Scope. In place of the Summary Snippets on the left was a table of contents. Participants could click a section header in the table of contents to navigate to the corresponding section.
6.1.2. Materials: ToS and User Persona
For the user study, we used the ToS of one social media site (Reddit) and one e-commerce site (Poshmark) because social media and e-commerce platforms are among the most visited digital services and are representative of the standard long-form contract used in prior studies  (Kay and Terry, 2010; Taber et al., 2020; Good et al., 2005; Tabassum et al., 2018; Obar and Oeldorf-Hirsch, 2020). For each service, we manually collected policies that are linked or hyperlinked as part of the ToS. We did not include policies that are location-specific (e.g., a California Privacy Notice). For Reddit, we also did not include policies for advertisers, brand&press, or governmental agencies as they were not relevant to a lay user and to keep the number of policies for both services similar. We collected a total of 14 policies for Reddit and 15 for Poshmark. The service names for Reddit and Poshmark were anonymized as ServiceX and ServiceY. Features of TermSight rely on a user persona to determine relevance. For the user study, we designed two personas based on users’ information needs identified in the formative study: a content consumer who posts personal content on social media sites (Figure  17) and a buyer who rarely posts reviews on E-commerce sites (Figure  18).
Figure 7. The experiment procedure for the within-subject study. At the top, we show the procedure for the entire study. At the bottom, we show the process diagram for a reading session with either TermSight or Baseline. The orders for the reading session were counterbalanced in the study.
6.1.3. Study Procedure
The entire study took 90 minutes and participants were compensated 25\$. The study was approved by the Institutional Review Board (IRB) office at our organization. The study was composed of 2 reading sessions, 45 minutes each (Figure  7). After obtaining consent and before the reading sessions, we asked participants about their past experiences with ToS. Within each reading session, participants first completed a pre-survey and reported their past experiences interacting with social media or e-commerce platforms. Then, an anonymized description of the service and user persona was given. Participants were given 10 minutes to read the ToS without having to speak aloud. Reminders were given at the 5 and 1-minute mark. After the reading session, participants completed seven 5-point Likert-style rating questions about their reading experiences followed by one free-response recall question and six multiple-choice comprehension questions. We also conducted semi-structured interviews with participants about their experiences and reading strategies for 10 minutes. The same procedure was repeated for the second reading session. The study materials are included in Appendix  F and the supplemental materials.
Figure 8. Participants’ self-ratings of their general familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service. Most participants were familiar with both social media and e-commerce platforms in general and found the given persona to align with their personal usage of the service and personal value.
6.2. Participants and Setting
We recruited 22 participants through Prolific. The recruitment required participants to be over 18, fluent in English, located in the US, not have color deficiencies, and not be legal professionals. We did not require participants to have prior experience in reading ToS. Two participants encountered technical difficulties, and their data were removed from the analysis. Of the remaining 20 participants, 13 self-identify as female and 7 as male. 1 participant had a high school degree, 3 had an associate degree, 12 had a bachelor’s degree, and 4 had a master’s degree. 4 participants were between the ages of 18-25, 5 were between 26-35, 5 were between 36-45, and 6 were between 46-55. When asked about how many ToS they’ve read, 4 participants read none, 5 read between 1—3, 5 read between 4—6, 1 read between 7—9, and 5 read greater than 10. The median of participants’ self-rated familiarity with ToS was 3.5 ($\sigma=1.5$). Most participants were generally familiar with social media and e-commerce platforms and found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  8). None of the participants had read or remembered the ToS for Reddit and Poshmark prior to the study.
6.3. Measures
6.3.1. Perceived reading experiences
We collected 5-point Likert-style ratings (1=“Not at all,” 5=“Very”) of participants’ reading experiences (Appendix  F.3). The experience measures included ease of reading, perceived understanding, and confidence in obtaining relevant information  (August et al., 2023). In addition, we included two 5-point Likert-style rating questions related to the ease of deciding which policy to read and what text to read within a policy, both of which are challenges identified in our formative study. Lastly, we included two 5-point Likert-style questions about their willingness to (1) read other ToS with the interface and (2) spend more time on their given ToS with the interface and get a link to the ToS after the study  (August et al., 2024).
6.3.2. Reading comprehension and recall
For reading comprehension, we wrote 10 multiple-choice questions for each service and selected 6 questions that were relevant for the given user persona for each service, similar to prior works on evaluating reading comprehension for ToS and privacy policy (Tabassum et al., 2018; Vu et al., 2007; August et al., 2023). The questions are different for the two services to minimize learning effects. Comprehension was measured as the number of questions participants got right out of 6. Examples of comprehension questions and their answers are listed in Table  1.
For recall, participants were asked to write down information in ToS they found interesting, surprising, or any detail they remembered: “What did you learn from reading the ToS? Recall one or more interesting things you learned from the ToS.” To analyze these free-form responses, we manually broke the response into single references to a clause in the ToS. Then, we labeled whether each reference was correct based on the ToS. Recall is measured as the number of correct facts participants wrote in the response  (Taber et al., 2020; Good et al., 2005).
6.3.3. Feature usage
To understand users’ interaction behavior, we logged all viewport scrolls and feature usage during the reading session along with timestamps. This allows us to measure the frequency of feature usage of TermSight and the baseline interface.
6.4. Analysis
In this section, we introduce our analysis framework for the user study. We used a causal framework (§6.4.1) to understand the effects of the treatment and a Bayesian analysis (§6.4.2) to estimate the treatment effect. Then, we discuss the thematic analysis (§6.4.3) used to analyze the qualitative responses from the semi-structured interviews.
6.4.1. Structural Causal Framework
We used a structural causal framework popularized by Pearl (2009) to better understand the effects of the treatment. A structural causal model framework comprises exogenous variables $U$, endogenous variables $V$ and a set of functions $F$ that represent the physical process that results in the values of the endogenous variables. For example, the question $v_{i}=f(v,u)$ implies that the values of $v$ and $u$ completely determine the value of $v_{i}$ through the function $f$. The structural causal model can be represented as a Directed Acyclic Graph (DAG) where the nodes represent the variables and the edges represent the causal relationships among the variables. Thus, in the preceding example, $v$ and $u$ are the parents of $v_{i}$ in the DAG. The structural causal model framework is non-parametric, but frequently used with generalized linear models to estimate the functions $f$. The structural causal framework has connections with structural equation models (Wright, 1934) used in the Economic Sciences, though the latter were developed with linear relationships in mind. Figure 9 shows the DAG for the study.
[…]
Figure 9. The causal graph (a DAG) has three types of nodes: Treatment ($T$) colored blue, pre-treatment covariates $D$ (demographics), $K$ (knowledge of Terms of Service) colored gray, and measured outcomes $Q$ (comprehension), $R$ (recall of correct facts), and $E$ (experience of the interface) colored orange. Since we use randomized assignment, including the interface type (treatment vs. control), type of ToS (social media vs. e-commerce website), and order of presentation (first vs. second), the treatment $T$ is independent of the pre-treatment covariates $D$ and $K$. Thus, to estimate the causal effect of the treatment $T$ on the outcomes $Q$, $E$, and $R$, we can use a simple regression model, without conditioning on the pre-treatment covariates $D$ and $K$ since there is no backdoor path from the treatment $T$ to the outcome variables $Q$, $E$, and $R$.
6.4.2. Bayesian Analysis
We used Bayesian inference to estimate the treatment effect of TermSight on the outcomes of interest. While the use of Bayesian estimates is growing in HCI (Kay et al., 2016; Koshy et al., 2023), we briefly justify its use over traditional non-Bayesian methods. Standard statistical methods including t-tests of significance can be powerful, but they require us to test for the conditions under which a test is valid (e.g., Normality for the t-test). Prior work in Biological and Psychological Sciences has shown that Null Hypothesis Statistical Testing (NHST) including $p$-values are problematic, and must be used with caution. They depend on the researchers’ intentions, when they stop collecting samples, and pre-selection of hypothesis (Ioannidis, 2005, 2012; Button et al., 2013).
We believe that Bayesian techniques offer several advantages over non-Bayesian methods. First, as Kay et al. (2016) point out, a Bayesian framework leads to an accumulation of knowledge within HCI, where the posterior of the parameters in a prior work can serve as the prior in the current experiment. Second, a Bayesian model is transparent—the researcher will foreground all the assumptions in the analysis via their model, allowing for a principled critique of the model. Third, use of a Bayesian framework shifts the discussion from “did it work” to “effect size of the intervention” (Kay et al., 2016). While NHST techniques can compute the confidence interval, these intervals are susceptible to misinterpretation (Belia et al., 2005; Hoekstra et al., 2014) and importantly, underemphasized in favor of $p$-values. Finally, as  McElreath (2015) points out, a Bayesian model with use of maximum entropy priors for the parameters (e.g., the normal distribution) is the most conservative given the evidence to estimate of the effect of the treatment.
6.4.3. Thematic analysis
To analyze the qualitative responses from the semi-structured interviews, we adopted the six phases of reflexive thematic analysis suggested by Braun and Clarke  (Braun and Clarke, 2006), following the same procedure as our formative study (§3.1.2).
7. Findings
7.1. RQ1: How did participants perceive TermSight and its features?
When asked about which version of the interface they prefer, 19 participants preferred TermSight. One participant preferred the baseline interface, citing difficulty navigating TermSight’s two scrollable columns on a small screen with a trackpad that lacks scroll functionality. Below we present our quantitative findings on the experience outcomes (§7.1.1), followed by qualitative findings organized around recurring themes about features of TermSight.
Figure 10. 5-Point Likert Ratings of participants’ reading experiences: Participants found it to be easier to find and understand relevant information. Participants were also more willing to spend more time on the ToS with TermSight and read other ToS with TermSight. The 7 experience questions are specified in Appendix F.3.
7.1.1. Quantitative Findings: TermSight improved user experience measures
Figure 10 shows participants’ ratings of the 7 reading experience measures for both interfaces. Our Bayesian model analysis (model details in Appendix D.1) reveals a significant effect of the TermSight interface on the reading experience measures. The forest plots shown in Figure 19 in the Appendix demonstrate that the 94% Highest Posterior Density Interval (HPDI [^1] In Bayesian analysis, the 94% HPDI refers to that smallest interval (this interval is unique) of the posterior distribution that has 94% of the probability mass. It is common in Bayesian analysis to use intervals distinct (97% and 89% HPDI intervals are also used) from the typical 95% value to avoid the confusion of the frequentist confidence intervals.) for each treatment–control pair for every user experience measure does not overlap with each other. This indicates a significant positive effect of the treatment on all 7 user experience measures.
In Figure 11, we show the contrast between the treatment and control conditions averaged across 7 user experience measures as well as the posterior of the effect size. The posterior distribution (Figure 11(b)) of the effect size (Cohen’s $d$) of the treatment on the user experience is centered around 2.7, with an HPDI of [2.3, 3.2], indicating a large effect size [^2] This is the effect size measured not on the outcome Likert scale (1–5) but on latent scale of the model coefficients.. Because the HPDI doesn’t overlap with the ROPE, the effect of TermSight on user experience is significant. Below, we further present recurring themes related to participants’ perceptions and feedback on features of TermSight.
(a) Posterior distributions of the coefficients for the treatment and control conditions.
(b) Posterior distribution of the effect size of the treatment on the user experience.
Figure 11. Figure 11(a) shows the posterior distributions of the coefficients for the treatment and control conditions. Note that the posterior distribution of the treatment condition is shifted to the right, indicating a positive effect of the treatment on the user experience. Figure 11(b) shows the posterior distribution of the effect size of the treatment on the user experience centered around 2.7, suggesting a large effect size. Additionally, the HPDI doesn’t overlap with the ROPE, suggesting a significant effect of the treatment on user experience.
7.1.2. Power Meter provided information scent of the policies and reduced navigation cost
All 20 participants described that the Power Meter helped them gain an intuition of the content of a policy and provided guidance. Participants described how the saturated colors, especially red and green, “piqued interest”(P4). For example, P12 described how Power Meter provided guidance and supported decision-making on which policy to pay attention to.
“The color system was really guiding a lot of my decisions on which policies to read. The vibrant red was always what I was gonna try to get to 1st and the things that I cared the most about. So, finding which terms I’m agreeing to that I have the least amount of agency.” (P12)
P6 used the TermSight system first. However, he lamented the fact that when using the baseline interface without the Power Meter, deciding which policy to read becomes a guessing game.
“Policies like terms of service or privacy policy are kind of broad. Does that apply to something I would actually care about? I don’t know. So I’d have to click each one and have to go through it. It was more like a guessing game.” (P6)
8 participants described how the document preview that appears when hovering over the Power Meter complemented and provided a more detailed overview of the policy to help them decide what to read. Additionally, 9 participants mentioned how the document preview allowed them to access information from linked pages or policies without “clicking on the link and possibly losing where I am on” (P1).
7.1.3. Summary Snippet and Color provided guidance and helped surface power and relevance within documents
18 Participants described how the color of the Summary Snippets helped surface where power lies in the contract, which is challenging to ”see” when presented in plain text with the baseline interface.
“The color coding gave me a better understanding of who’s benefiting more, how is it benefiting them, and how it affects me. Versus [in the baseline interface] it feels like it only really benefits the company itself.” (P18)
Moreover, all participants described how the color and summary allowed them to make decisions and prioritize what to focus on strategically.
“I really liked having the summary with colors. I had a hierarchy of how I was going to read stuff. I made sure I read all the [saturated] red ones, and then the [saturated] green. If I had time, I would skim the [saturated] yellow and see if anything stuck out. It would take me a lot longer to filter for information otherwise.” (P6)
While participants described focusing on the more saturated colors, especially red and green, 6 participants also described skimming the less saturated ones to validate and make sure they didn’t miss information.
7.1.4. Summary Snippet simplified and broke down dense text to support sensemaking
All participants highlighted how the Summary Snippets “broke down the entire section into a few bullet-pointed summaries” (P11), which helped them absorb the information and surface relevant pieces of information hidden in dense text. 9 participants further described how the Summary Snippets simplified the text and served as a scaffold to encourage reading the original text.
“Not only does it [TermSight] bring me up to things that I would normally miss. but then that would be an introduction or a guide, or a push towards reading the whole 3 paragraphs. … It is really guiding me through this entire document, and it’s simplifying it.” (P2)
Similarly, 4 participants mentioned how the highlight bars and line breaks added for each information snippet made it easier to consume the original text without having to “mentally separate the concepts” (P7). P20 described the experience of using Summary Snippets to make sense of the big ideas of the section while being able to dive deeper as forming “a web of thought and ideas: what everything kind of means together and piece by piece”. Not only do the Summary Snippets serve as an entry point to the original text, 4 participants described how the Summary Snippets helped them check their understanding after reading the original text.
7.1.5. Term definition and scenario help consume original text and envision implications
14 participants described that the underlined phrases attracted their attention, made them feel like these were phrases that “should be known and understood” (P18), and prompted them to check on their understanding of the phrase.
In addition, participants commented on how the definitions and scenarios made legal language more approachable (14/20).
“It really just made it a lot easier to understand, because most of the time they write it in legal terms that most people don’t know. People don’t speak legal language. So this broke it down easier for most people like the Layperson.” (P17)
12 participants highlighted how the scenarios helped interpret the meaning and envision potential implications of the phrase in the context of the user’s intended usage of the service and was “easier to resonate”(P17).
“I like that it provides a real-world example of when and how [the phrase] would be relevant to someone. I think it actually does better explaining the concept than the definition. Because it’s not just the definition. It’s the definition in context.” (P11)
Moreover, one participant (P1) noted how the scenarios helped her come up with more questions, which she asked with the Ask function. Through this process, TermSight helped P1 to quickly obtain information from correlated policies.
7.2. RQ2: How did participants read with TermSight and its features?
Figure 12. Usage of features during the reading session. Each dot represents the number of times a feature is used by one participant during a reading session. Participants used most features of TermSight. In contrast to the baseline, participants when using TermSight more frequently clicked the Summary Snippets as opposed to the section titles to navigate to the original text.
Figure 13. Minute-by-minute usage of features of TermSight during the ten-minute reading task. Participants used these features throughout the reading session, rather than being limited to the beginning or end. The Summary Snippet is the most frequently used feature throughout.
7.2.1. Feature Usage
In the baseline interface, participants on average navigated to 7 policies ($\sigma=5$). Participants’ reading behaviors were similar to our findings from the formative study (§3.3.2). Participants described how they were trying to skim everything from top to bottom. When they encounter sections that were hard to read, they tended to skip. Only 11 participants clicked the section heading in the table of contents at least once to navigate to targeted sections.
As shown in Figure  12 and  13, participants used most features of TermSight throughout the session, demonstrating that TermSight provided value throughout the session. When using TermSight, participants on average navigated to 6 policies ($\sigma=3$). 19 participants used the document preview of Power Meter to gain an overview of documents in ToS for more than 1 second. 13% (17/135) of the total usage was for previewing documents hyperlinked inline, demonstrating that the Power Meter provided value also for inline hyperlinks.
Summary Snippet was the most frequently used feature, 19 participants clicked the Summary Snippet to navigate and read the original text at least once. On average, they clicked 17 Summary Snippets ($\sigma=13$). In contrast, participants rarely clicked the section headings to navigate to the original text (4 clicks in total across 3 participants). This demonstrates that Summary Snippets were potentially more useful as navigational affordances to surface and interrogate relevant information from the original text compared to section headers. We also noticed how 10 participants clicked the Highlight Bar to read the summary snippet after reading the original text, suggesting how these participants were back-checking their understanding using the summary snippets.
For Phrase Scope, 16 participants viewed the term definitions and 12 participants viewed the term scenarios. Participants on average used term definition 3 times ($\sigma=4.3$) and term scenario 1 time ($\sigma=2.7$). Out of the 77 instances when Phrase Scope was accessed, 72 were suggested by TermSight. Participants rarely manually highlighted phrases to activate phrase scope, asked questions, or checked the references provided for the definitions. While only 1 participant asked questions using Phrase Scope, P1 asked 8 questions in a 10-minute reading session.
While TermSight provided multiple reading affordances, participants were in control of deciding whether or not to rely on these features. As opposed to a consistent preference for one interaction strategy over others, we observed diverse ways that participants relied or did not rely on TermSight (Figure  14). Below, we provide case studies of these non-exclusive emergent reading strategies.
Figure 14. Scrolling behavior of participants that employed each non-exclusive reading strategy. TermSight features two scrollable columns: Summary Snippets on the left (red lines) and Original Text on the right (blue lines). Clicking a Summary Snippet (orange X) auto-scrolls the Original Text panel to the referenced text, while clicking the Highlight Bar (light blue X) auto-scrolls the Summary panel to the corresponding Summary Snippet. (a) Summary-driven reading occurs when participants primarily read and scroll the Summary Snippets, occasionally clicking on them to interrogate the original text (19/20). (b) Original text-driven reading occurs when participants first read the original text and use the Summary Snippets as a supplement (9/20). (c) The calibration of trust mainly occurs at the beginning of the session where participants would compare and contrast the original text to the Summary Snippets with color (6/20).
7.2.2. Summary driven reading
19 out of 20 participants at times relied on the summary snippets as a primary entry point for engaging with the text. One notable example of this behavior was P3, who mainly read the Summary Snippets and clicked 11 of them to interrogate the original text (Figure  14a). P3 described how colors were driving his attention to read. P3’s decision to click the Summary Snippets and interrogate the original text relied on the color coding and evaluation of the information loss of the Summary Snippets.
“I mainly stay on the left side. I read them if it was like a saturated red or green. I clicked on them because I felt like there was more information for me [behind the Summary], and I want to learn more about it.” (P3)
7.2.3. Original text driven reading
At other times participants centered their interactions around the original text. 9 participants went through at least one policy focusing mainly on the original text. These participants described how they wanted to first read the actual text before relying on AI-generated summaries due to the potential imperfections of AI. These participants explained, “I didn’t want my opinion to be affected by AI” (P5) and “sometimes summaries don’t have all of the information from my other experiences” (P18).
When reading the Privacy Policy, P8 focused on the original text as shown by the frequent scrolling of the original text in Figure  14b. In addition, P8 clicked the Highlight Bar 11 times to refer back to the Summary Snippets. P8 described how he knew AI was not perfect from his prior professional experiences in AI. As a result, P8 took the colors and summary snippets as a guide rather than a guarantee.
“I’m not gonna just go and trust it without doing my due diligence. I would take it as a guide. I won’t take it as a guarantee, because AI is not perfect. I think it [AI] would try to label it the best it can and I think it did, to be honest.” (P8)
7.2.4. Calibration of trust
6 participants described that they went through a calibration process with AI-generated features such as the Summary and Color at the start of the reading session. For example, P6 described how his reading behavior was different during the calibration process and after. P6 described reading the original text on the right side to evaluate the colors and summaries at the beginning of the reading session (original text driven reading). Afterward, P6 relied on the summaries a lot more by scrolling and skimming the left panel and clicking on the summary snippets to dive to the original text as seen in Figure  14c (summary driven reading).
“So like early on in the reading, I kind of read the whole thing on the right first. Then, I looked at the summaries, and once I realized that it did do a good job summarizing it with colors. I trusted it enough to trust it for the rest of the time.” (P6)
7.2.5. Switching strategy based on time and document.
ToS includes multiple documents, and in the study, there were 14-15 policies with varying lengths. We noticed how participants’ reading behavior can vary based on document length (P5, P19). For example, P19 mainly read the summary for longer policies such as User Agreement and Privacy Policy (Summary driven reading). For shorter policies such as Buyer Policy, P19 only read the original text (Original text driven reading).
We also noticed how time can change participants’ reading strategy (P7, P8). For example, P7 spent 8.5 minutes and was very interrogative when reading the User Agreement policy. When she realized that she was running out of time, she skimmed through the summary snippets for 5 more policies without interrogating the original text.
7.3. RQ3: How did TermSight influence comprehension and recall?
Below we present the results of the comprehension quiz (§7.3.1) and recall task (§7.3.2).
7.3.1. Comprehension
For the comprehension quiz, out of 6 questions, participants on average scored 2.00 ($\sigma=1.03$) when using TermSight and 2.15 ($\sigma=1.23$) when using the baseline interface. Our Bayesian model analysis for the comprehension quiz (model details in Appendix  D.2) reveals no significant differences in the comprehension scores across interface conditions. As shown in Figure 15, the posterior distributions for both conditions were nearly identical and centered around zero. When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair as shown by the forest plots in Figure 20 in the Appendix, suggesting no significant differences.
Figure 15. The figure shows the posterior distributions of the coefficients for the treatment and control conditions. Both posterior distributions are nearly evenly split around zero, indicating no effect of the treatment on the comprehension quiz questions.
7.3.2. Recall
For the recall task, participants using TermSight recalled 39 correct facts and 3 incorrect facts in total. On the other hand, participants using the baseline interface totally recalled 26 correct facts and 7 incorrect facts. Our Bayesian model analysis for the recall task (model details in § D.3) reveals no significant differences in the recall scores across interface conditions. Figure 16(b) shows the posterior distribution of the effect size of the treatment on recall ($\mu=0.6$, HPDI = [-0.39, 1.6]). Despite the mean being 0.6, suggesting a medium effect size, the HPDI interval has an overlap with the ROPE, implying that there is no significant effect of the treatment on the user recall.
(a) Posterior distributions of the coefficients for the treatment and control conditions.
(b) Posterior distribution of the effect size of the treatment on user recall.
Figure 16. Figure 16(a) shows the posterior distributions of the coefficients for the treatment and control conditions. Specifically, the posterior distribution of the treatment condition is shifted to the right, indicating a positive influence of the treatment on Recall. Figure 16(b) shows the posterior distribution of the effect size of the treatment on the user recall centered around 0.6, suggesting a medium effect size. Yet, the HPDI has an overlap with the ROPE of [-0.1, +0.1], implying no significant effect of the treatment on the user recall.
8. Discussion
In this paper, we set out to design a new reading experience that scaffolds readers’ ability to find and understand information important to them within Terms of Service (ToS) contracts. Based on formative interviews and observations, we realized this vision in TermSight. Results from our evaluation of TermSight (N=20) suggest that TermSight made it easier to read, navigate, and contextualize information within ToS relevant to users. Especially interesting was how TermSight afforded multiple different reading strategies (e.g., summary-driven reading, §7.2.2), all of which helped participants overcome the key difficulties of contract reading we observed in our formative study (§3). However, similar to prior works on simplifying ToS or other complex documents  (Vu et al., 2007; Tabassum et al., 2018; Good et al., 2005; Korunovska et al., 2020; Robinson and Zhu, 2020; August et al., 2023), TermSight did not worsen or improve the comprehension or recall of the content. Below, we discuss the implications of our findings and the TermSight system.
8.1. Existing ToS are not designed to be read
Many existing ToS have static or interactive design elements, such as policy names, section headers, and overview summaries (Vu et al., 2007). However, participants in our formative study found these features vague, misleading, and ineffective for surfacing and interpreting relevant information (§3.3). The HCI literature often refers to these ineffective and deceptive designs as dark patterns in the user interface that manipulate behavior or obscure information to serve the interests of the service providers  (Mathur et al., 2021; Bösch et al., 2016; Kyi et al., 2023).
TermSight offers one possible design to overcome these dark patterns by offering contract-level guidance (e.g., Power Meter and document previews), document-level guidance (e.g., color-coded Summary Snippets), and phrase-level guidance (e.g., phrase identification). Participants in our usability study (§7) affirmed these design goals in both self-report and behavior metrics. Participants reported that features of TermSight helped them decide which policy to read (Power Meter, Document Preview) and what text to read within a policy (Summary Snippet). Moreover, participants highlighted how TermSight made the original text more approachable with Summary Snippets that broke down and simplified dense sections and Phrase Scope that clarified unfamiliar phrases and helped envision implications.
Future systems could take these designs further by adapting them to other types of contracts, such as leasing contracts, medical contracts, or employment contracts. For example, Summary Snippets may help readers surface and interpret relevant or predatory clauses in leases, while phrase scope may help patients envision the implications of clauses in the medical contract through patient-specific scenarios. Lastly, while we designed TermSight as a contract reading interface, TermSight can be readily extended as a browser extension that highlights relevant and important clauses based on user-defined personas to provide in-situ legal awareness and warnings when encountering contracts.
8.2. Envisioning consequences to legal text
Prior work has explored providing term definitions to aid comprehension of academic text (August et al., 2023; Head et al., 2021). Yet, participants in our study mentioned that definitions alone were not always sufficient for legal language. Participants noted that it can be difficult to contextualize abstract phrases in real-world scenarios or anticipate potential consequences—especially those they have not yet personally experienced. Our formative study further revealed how people’s intended usage of the service, personal values, and worries can vary. While prior work explored predicting unintended consequences for scientific endeavors and AI  (Wang et al., 2024; Pang et al., 2024), TermSight explored the potential value of presenting personalized scenarios to illustrate the possible implications of legal phrases before a user agrees to a ToS. Some participants even noted how the scenario was more helpful than the definition (§7.1.5). By grounding abstract terms in concrete, relatable examples, these scenarios may help users better reflect on the potential implications behind legal language. Future work could investigate additional ways to help consumers reflect on unintended consequences, such as by showing multiple scenarios or relevant legal cases.
8.3. Design for calibration and trust
LLM output can be imperfect due to hallucinations or confusing output  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024). It is therefore important to help users form appropriate trust and reliance for fostering effective Human-AI collaboration  (Schemmer et al., 2023; Cao et al., 2024). In this vein, prior reading interfaces integrating AI have introduced cognitive forcing functions (Buçinca et al., 2021), such as showing summaries only on demand (August et al., 2023) or at the end of the reading session (Chen et al., 2023), to encourage readers to first read the original document. Yet, it is unclear if these designs are always optimal for intelligent reading interfaces, since they can increase user frustration and may limit specific reading behaviors  (Buçinca et al., 2021). For example, these forcing functions, if applied to TermSight, might introduce significant friction for participants taking a summary-driven reading approach (§7.2.2).
TermSight took the approach of tightly linking the AI-generated summaries to the original text to allow users to compare and contrast the two. In TermSight, users can navigate between and compare the summary and the original text in one click. When given these affordances, participants took advantage of them. Some participants explicitly evaluated the Summary Snippets to calibrate their trust. Interestingly, one participant framed the evaluation of AI-generated output as his ‘due diligence’. Even those who took a summary-driven reading approach frequently interrogated the referenced original text as opposed to solely reading the summaries. Future work can explore how to help users calibrate trust with the system and evaluate AI output. This could include providing explanations of how the reading interface works in general, providing examples of AI imperfections, and providing explanations for individual features  (Schemmer et al., 2023).
8.4. Ethical Considerations and Social Implications
Features of TermSight (e.g., Summary Snippets, Classification of Power and Relevance, Term Definitions, and Scenarios) rely on the capabilities of LLMs. However, LLMs have been shown to produce imperfect outputs and factually incorrect hallucinations  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024). This introduces risks when interacting with LLM-generated text, as they are not part of the contract and are not legally binding. For example, incorrect summaries might cause readers to misinterpret the clauses in the contract that are binding. On one hand, LLMs have already shown strong capabilities in generating summaries (Zhang et al., 2024; Liu et al., 2024b), term definitions (Yin and Skiena, 2023; Jhirad et al., 2023), and scenarios (Jiang et al., 2024; Wang et al., 2024; Feng et al., 2024; Shaikh et al., 2024). Work is also actively investigating how to detect non-factual LLM output  (He et al., 2024) and how to guide LLMs to generate factually correct information  (Vu et al., 2024). As a result, LLM output would ideally become more factually correct over time. On the other hand, for future deployment of systems like TermSight, we still believe a validation mechanism for LLM output is necessary. For example, future implementation of TermSight’s features (e.g., term definitions and scenarios) can be accompanied by additional sources of information when available (e.g., news articles, court cases, or information from other ToS). Furthermore, TermSight could facilitate end-user auditing (Lam et al., 2022; Jahanbakhsh et al., 2022), allowing users to report inaccuracies. These user contributions could collectively help validate and refine model outputs.
8.5. Limitations and Future work
The generalizability of TermSight and our study findings is limited by the challenges associated with handling ToS documents and our sample of participants in the user study. TermSight currently relies on a set of precleaned Markdown or HTML files. Future works can investigate more advanced scraping and cleaning methods. In addition, our recruited participants were mostly college-educated and fluent in English. Marginalized communities, such as neurodivergent readers and individuals with limited English literacy, can face additional barriers when interpreting legal documents. For example, readers with ADHD might have a hard time staying focused when reading visually dense text such as ToS  (Stern and Shalev, 2013; Bental and Tirosh, 2007). Future works can investigate how to help marginalized populations make sense of legal contracts by evaluating TermSight with these populations of users or integrating additional supports such as translation.
9. Conclusion
ToS are service contracts outlining the rules and expectations for both the service provider and the consumer. Our formative study revealed ineffective and deceptive designs at all granularities of a ToS. To make ToS approachable for consumers, we designed and evaluated an intelligent reading interface: TermSight. Participants reported that TermSight reduced the difficulty of engaging with ToS and increased their willingness to do so. Participants highlighted that features of TermSight afforded them the ability to surface and interpret relevant information at all levels of ToS reading. They found the Summary Snippets with colors particularly helpful in breaking dense sections into snippets of simplified information, which helped them surface relevant information and power balance. The ‘duty to read’ originated when contracts were brief and based on mutual understanding and negotiation  (Kar and Radin, 2019; Benoliel and Becher, 2019). Modern ToS contracts, however, prioritize unilateral control over mutual agreement by design. TermSight presents a vision and serves as a technology probe into making legal contracts more accessible to the general public, providing consumers with the tools to make more informed legal decisions.
Appendix A Iterative Design
A total of 8 participants were recruited through Prolific to evaluate an early prototype of TermSight. This prototype closely resembled the current version of TermSight but lacked a document preview when hovering over the Power Meter. Additionally, the prototype used three saturation levels for relevance, resulting in a total of nine colors (three colors for power X three saturation levels for relevance). Participants could also toggle between two layouts for the Summary Snippets: condensed layout and in-context layout. The condensed layout matched the current design shown in Figure  1. In contrast, the in-context layout placed each Summary Snippet directly next to its corresponding Information Snippet. After a brief interface tutorial, participants were given 10 minutes to read a ToS using the interface followed by a semi-structured interview about their experience. Each session lasted for 40 minutes. Below we present the main findings and changes made.
General findings.
All participants described that reading with the prototype was easier and more engaging compared to typical ToS, which they tend to skip entirely or skim superficially. In contrast to their usual approach of skimming and trying to search for keywords that catch their eyes, all participants described how the Summary Snippets directly highlighted relevant information otherwise buried in dense sections. Additionally, participants expressed interest in reading the original text after using the Summary Snippets to surface relevant information and valued the ability to navigate between the summary and the original text. 6 participants found the scenarios provided by Phrase Scope to be particularly helpful as the definitions alone can be difficult to contextualize and envision implications.
Participants preferred condensed layout for quick navigation
7 participants preferred and primarily used the condensed layout of the Summary Snippets during the reading session. They noted that the purpose of the Summary Snippets was to support navigation, and the condensed layout made it easier to gain an overview and identify relevant information in a policy. In contrast, the in-context layout required significantly more scrolling, which participants found overwhelming. As a result, in the final version of TermSight, we used the condensed layout for the Summary Snippets (Figure  1).
Reduce visual complexity of the color scheme
All participants found the colors to be intuitive and effective in highlighting power and relevance, helping them decide what to read. However, 3 participants pointed out that while the power dimension was intuitive, the relevance dimension with 3 saturation levels was challenging to differentiate as there would be 9 colors in total. To reduce visual complexity, the final version of TermSight included only two saturation levels (High vs. Low) for relevance (Figure  2).
Complement Power Meter with Document Preview
Additionally, 4 participants suggested that while the Power Meter provided a general overview, they wanted a more concrete preview. This feedback led to the integration of a document preview feature in the final version of TermSight, which appears when users hover over the Power Meter (Figure  3).
Appendix B Additional Implementation Details of TermSight
In this section, we provide additional implementation details of TermSight.
B.1. Source Document and Pre-processing
[…]
B.2. Obtaining Summary Snippets and Information Snippets
The prompt used to obtain the Summary Snippets and Information Snippets is detailed in Figure  21. We constrained the summary snippets to be short because prior works have found that adding short summaries (10-20 words) under search results was more effective for navigational and information-seeking tasks compared to longer summaries, which can be harder to skim  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006). This length constraint was further optimized through prompt engineering. We noticed that setting summaries shorter than 12 words led to excessive fragmentation (i.e., too many summary snippets), while longer summaries tried to cover too much information, making them visually dense and less skimmable.
B.3. Classifying Information Snippets
In TermSight, two classifications were performed for each Information Snippet using GPT-4o with few-shot prompting (Figure  22 and  23). For the classification of Power, each Information Snippet was classified by the degree of control or agency it grants to the Service Provider or the User (Categories: Service, Neutral, User). For the classification of Relevance, Information Snippets were classified based on their relevance to the user persona (Categories: High, Low). User personas used for the study are included in Figure  17 and  18.
B.4. Phrase Scope
B.4.1. Identifying unfamiliar or vague phrases
For each document chunk, we employ two few-shot prompts to GPT-4o to identify potentially unfamiliar (Figure  24) or vague phrases (Figure  25). Few-shot examples for identifying unfamiliar phrases were selected based on phrases participants found challenging in our formative study, which included legal jargon, copyright licenses, and privacy-related terminology. Few-shot examples for identifying vague phrases were informed by vague phrases participants identified in the formative study, as well as prior works analyzing vagueness in Terms of Service (ToS) documents  (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). Both prompts were applied to each document chunk, producing two sets of identified phrases. Since a phrase may be both unfamiliar and vague, the union of these two sets was taken as the final set of identified phrases for a given chunk of text.
B.4.2. Generating phrase definitions and answers to user questions
We used a retrieval-augmented question answering approach  (Fok et al., 2024a) to generate the phrase definitions. Figure  26 shows the prompt used to generate the definitions after retrieving potentially relevant document chunks. The prompt explicitly instructs GPT-4o to produce a concise in-context definition of the input phrase and output a list of the retrieved chunks referenced in generating the definition. In the case when the definition is generated without referencing any of the retrieved chunks, it is still retained. TermSight intends to provide plausible definitions even when explicit definitions are absent from the document. This design decision is also supported by prior research demonstrating the capabilities of large language models in generating term definitions that are comparable to traditional dictionaries  (Yin and Skiena, 2023; Jhirad et al., 2023). When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked (Figure  28). If the answer is not specified in the retrieved chunks, the system generates a plausible answer based on the phrase context, similar to the definitions.
B.4.3. Generating scenarios
Our needfinding reveals the need to help readers not only interpret the definition of a phrase but to envision potential implications. As a result, we leveraged GPT-4o to generate customized scenarios of potential implications based on user persona. The prompt is specified in Figure  27.
Appendix C Evaluation of TermSight Output
We manually evaluated TermSight’s feature pipelines and overall performance. Our goal was not to assess advances in system performance, but to verify that the system reliably produced meaningful outputs that could support ToS reading.
C.1. Evaluating the Classification of Information Snippets
We sampled 116 Information Snippets with stratified sampling from all the sub-policies for each service included in the user study. For each Summary Snippet, we evaluated the classification of power and relevance. Of the 116 Information Snippets, 27 were classified by TermSight as favoring the user, 22 were classified as favoring the service, and 67 were classified as neutral. In addition, 59 were classified as of high relevance to the provided user persona and 57 were classified to be of low relevance.
Our evaluation revealed 3 instances where the power classifications were imperfect, mainly caused by the lack of context in the input information snippet. For example, the snippet (‘All Buy Now purchases in a ServiceY Show are final and binding’) was classified by TermSight as favoring the service. In the generated explanation, the assumption was that the statement meant that the user can not return or refund the purchase. However, users can return and get a refund if there are problems with the purchase (e.g., an item doesn’t match its description) as stated in the service’s return policy, making this snippet more neutral. The return policy was not included in the snippet context, as a result, this neutral statement was considered more service-oriented.
For the classification of relevance, there were 2 instances, out of 116, where irrelevant information snippets were classified as relevant to the persona. In both cases, the snippets were more relevant to sellers, not buyers, even though the persona given was that of a buyer. There were 11 instances where the information snippets classified as relevant were indirectly relevant to the input user personas. For example, sellers have to pay a fee to the platform after selling an item. Despite this information being more relevant to sellers, TermSight classified it as relevant to buyers as it exposed the potential hidden fees that sellers might include as part of their listing price.
C.2. Evaluating Term Definitions and Scenarios
We randomly sampled 113 generated definitions and scenarios using stratified sampling from all the sub-policies for each of the two ToS augmented for the user study. These were evaluated by members of the research team. Out of the 113 sampled phrases, 46 were identified by TermSight as vague phrases, 6 were identified as both jargon and vague phrases, and 61 were identified as jargon.
Our evaluation revealed that all the generated definitions were correct, but out of the 113 definitions, 4 were only general definitions of the phrase not specific to the ToS. On further inspection, we identified that this was because these vague phrases were not explicitly defined anywhere in the ToS. Additionally, their meanings are service-specific and cannot be extrapolated from common sense or inferred by large language models (LLMs). For example, for the phrase ‘aggregated anonymized statistics’, TermSight provided a general definition of what it might mean to aggregate and anonymize user data. However, the specific details—such as what user data are being aggregated—were not specified in the ToS. Rather than a limitation of TermSight, these imperfections highlight the ill-defined nature of the language used in ToS.
For the generated scenarios of phrases relevant to the user persona, we found one instance where the scenario was factually incorrect based on the input context. The input context and phrase stated that users do not gain ownership rights by downloading content from the service. However, the scenario claimed that users might lose ownership rights over the content they create by uploading it to the service. We also noticed that when the passages or phrases target a different audience (e.g., developers) than the user persona used to generate the scenarios (i.e., a lay user), the generated scenarios become less relevant or useful. We did not regenerate these imperfections to keep the user experience realistic to real-world settings where the LLM outputs are not guaranteed to be perfect. Participants in our user study were informed that AI output can be imperfect. This specific incorrect scenario was not accessed by any participants in the study.
Appendix D Bayesian Analysis Model Details
In this section, we provide additional details on the models used in the paper. All models used the DAG in Figure 9 to inform the model. Next, in § D.1, we describe the model used to estimate the effect of the treatment on the responses related to the experience of using the treatment interface. In § D.2, we describe the model used to estimate the effect of the treatment on the comprehension outcomes i.e., the quiz questions. Finally, in § D.3, we describe the model used to estimate the effect of the treatment on the experiment subjects’ ability to recall facts correctly.  Figure 19 and  Figure 20 show the forest plots of model coefficients for the experience and comprehension outcomes.
D.1. Modeling Experience Outcomes
We asked the experiment subjects to rate their experience of using the treatment or the baseline interface on a Likert scale from 1 to 5 (L=5 outcome values). There were N subjects and we asked each subject K (K=7) questions. Since the likert scale is ordinal, we modeled the experience outcomes using an ordinal regression model. The DAG in Figure 9 indicates that the experimental condition alone affects the experience outcomes. There were three randomizations: whether the subject was shown the treatment or control interface, and whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second. We modeled the experience outcomes using a cumulative logit model with a linear link function.
The model is given by:
[…]
In eq. 7, we further ensured that the samples are ordered i.e., $\kappa_{1}<\kappa_{2}<\dots<\kappa_{L-1}$. The priors are conservative. For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to outcome probability range of [0.04, 0.95].
D.2. Modeling Comprehension Outcomes
Like earlier, there were N subjects, and we asked each subject K (K=6) questions to test their comprehension of the ToS. Thus, the outcome is a binary variable with 1 indicating a correct answer and 0 indicating an incorrect answer. The DAG in Figure 9 indicates that the experimental condition alone affects the comprehension outcomes. There were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second. We modeled the comprehension outcomes using a logistic regression model. Notice that we created a joint variable $c_{j,k,l}$ for the coefficients corresponding to the question, service, and treatment/control. We did this because the questions were different across the two services used in the experiment. We included the treatment/control variable to be able to easily estimate the effect of the treatment on the comprehension outcomes per question.
The model is given by:
[…]
As in the previous model, the priors are conservative. For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to outcome probability range of [0.04, 0.95].
D.3. Modeling Recall Outcomes
There were N subjects, and we asked each subject a question on recall facts from the ToS. The outcome that we measure is the number of correctly recalled facts, making the outcome a count variable. Since many of the subjects could not recall any facts correctly, we modeled the recall outcomes using a Zero Inflated Poisson regression model. The DAG in Figure 9 indicates that the experimental condition alone affects the comprehension outcomes. As before, there were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second.
The model is given by:
[…]
As in the previous model, the priors are conservative. For example, on the Log scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse log to outcome range of [0.05, 21]. Notice that we model the number of correctly recalled facts, and thus is a conservative prior. Given that most subjects could not recall any facts, the zero inflation parameter is set to a weakly informative prior using a Beta distribution.
Appendix E Study Materials: Formative Study
Questions asked before the reading session:
  • How familiar are you with Terms of Service in general?
  • Have you previously read or wished that you read the Terms of Service? What were your reasons for wanting or not wanting to read the Terms of Service?
  • Have you used your assigned service before?
  • Have you read the ToS for your assigned service before?
Semi-structured interview questions asked after the reading session:
  • What were the challenges you faced when reading the Terms of Service?
  • How did you go about reading the ToS?
  • What information were you interested in in ToS? Both from your prior experience in reading ToS or this ToS?
  • Imagine if you have a magic wand that can transform the Terms of Service in whatever ways you want. How would you transform the Terms of Service?
Appendix F Study Materials: User Study
F.1. User Persona
Two personas were given to the participants during the user study for the social media service (Figure  17) and e-commerce service (Figure  18). The same personas were used for features of TermSight to classify relevance and generate personalized scenarios. The personas were designed based on information participants in the formative study described caring about.
User Persona for Social Media site: Content consumer who posts personal content Imagine you are a lay user of social media platforms. You are over 18 years old and located in the United States. Your usage of Social Media sites: You spend most of your time on the platform scrolling through feeds, liking posts, chatting with other users, and sharing personal content such as photos. Things you care about when using Social Media Sites: You care about Privacy, particularly what data are being collected and how your data can be used and shared. You care about what the service can do with user-generated content such as licenses over user content or advertising with user content. You care about potential liabilities when using ServiceX.
Figure 17. User Persona for Social Media site: Content consumer who posts personal content
User Persona for E-commerce site: Buyer who rarely posts reviews Imagine you are a lay user of E-commerce platforms. You are over 18 years old and located in the United States. Your usage of E-commerce sites: You typically engage with the E-commerce platform to buy new or used items from other users. You rarely post any reviews or content on the service. Things you care about when using E-commerce sites: You care about information related to making purchases, refunds, returns, user protection policies, termination, arbitration, and liabilities. You also care about Privacy, particularly what data are being collected and how your data can be used and shared.
Figure 18. User Persona for E-commerce site: Buyer who rarely posts reviews
F.2. Pre-survey Questions
Questions that were asked only once at the beginning of the interview:
  • For how many online platforms have you read their Terms and Conditions (T&C) before? [None (0), Few (1-3), Some (4-6), Many (7-9), A lot (¿10)]
  • How familiar are you with Terms and Conditions (T&C) for online platforms? (5-point likert rating)
5-point Likert rating questions that were asked before each of the two reading sessions:
  • How familiar are you with (E-commerce or Social Media) sites?
  • How well does the above user persona align with your personal usage of (E-commerce or Social Media) sites?
  • How well does the above user persona align with things you personally care about when using (E-commerce or Social Media) sites?
F.3. 5-point Likert Ratings of Reading Experience
Participants rated their reading experience after each of the two reading sessions.
  • E1: I’m interested in spending more time reading the service’s Terms and Conditions (T&C) with the current interface and wish to get a link after the study.
  • E2: How hard did you have to work to read the T&C?
  • E3: How easy was it for you to decide which sub-policies to read?
  • E4: How easy was it for you to decide what text to read within a sub-policy?
  • E5: How confident are you that you got all the relevant information from the T&C (including sub-policies)?
  • E6: How much do you feel like you understood the T&C?
  • E7: How much would you be willing to read the T&C of other services with this interface?
F.4. Semi-structured Interview Questions
Below are questions that were asked after each of the two reading sessions, except the last question which was only asked after the second reading session where participants would compare and contrast both interfaces.
  • Describe your experience reading the Terms of Service using the interface.
  • How did you read the ToS using the interface?
  • What were the challenges you faced when reading the Terms of Service? To what extent did the interface help you? (be specific about the features in the interface).
  • Which interface do you prefer? Why? Compare and contrast. (Asked only after the second reading session)
Question Correct Answer Relevant Passage
How can ServiceX use photographs you post for advertisements or promotions? ServiceX can use photos you post in ads without your permission and is not obligated to attribute you as the creator. [Social Media] “…you grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content and any name, username, voice, or likeness provided in connection with Your Content in all media formats and channels now known or later developed anywhere in the world…you irrevocably waive any claims and assertions of moral rights or attribution with respect to Your Content.” (User Agreement)
ServiceX licenses the content you post to third parties. Which of the following is a third-party licensee of user content on ServiceX allowed to do? Train AI models with user content without user consent. [Social Media] “For example, this license includes the right to use Your Content to train AI and machine learning models” (User Agreement) Our licensees cannot: Combine third-party data with ServiceX public content to target ads; Perform background checks; Track, alert, monitor, or investigate sensitive events (for example, protests or rallies) or sensitive organizations (for example, unions or activist groups)…” (Public Content Policy)
You bought a pair of shoes from ServiceY. Before accepting the delivery, you tried the shoes on but found the shoes small for your foot. What can you do to return the item and guarantee a refund? ServiceY does not accept “item doesn’t fit” as a valid reason for return. [E-commerce] “Trades, offline transactions, items that do not fit, or orders where the Buyer changed their mind are not covered by ServiceY Protect. These types of purchases are considered final sale and not eligible for return.” (Return Policy)
According to ServiceY’s authentication policy, when you buy an item over 500\$, ServiceY will first authenticate the item before shipping it to you. If ServiceY decides that the item is fake, what would happen? The item will be disposed. You will be refunded. [E-commerce] “If we reasonably believe an item to be counterfeit, we will provide a refund for the full cost of the item and any accompanying shipping costs to the buyer … if an item in question is deemed counterfeit, we shall ensure disposal of the item at our discretion.” (Counterfeit Items Policy)
Table 1. Example multiple choice comprehension questions: The questions were designed to be difficult and can’t be directly answered by only reading the Summary Snippets. These questions can also require knowledge about information from multiple positions of the same policy or information across different policies.
Phrase Type Original Passage Generated Definition Generated Scenario
Jargon “… When Your Content is created with or submitted to the Services, you grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and **sublicensable** license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content and any name, username, voice, or likeness provided …” Sublicensable refers to the ability to grant further licenses to third parties to use, copy, modify, or distribute your content. Imagine posting a photo on ServiceX, and later discovering it’s featured in a third-party app’s advertisement. Due to the sublicensable license, ServiceX had the right to allow another company to use your photo, without needing your further permission, raising concerns about control over your content.
Jargon “… The listed prices for Items do not include taxes, but the taxes will be displayed before a Buyer confirms the purchase. Use of **ServiceY Credit** (as defined below) may modify taxes that apply to a Buyer’s purchase … ” ServiceY Credit refers to non-redeemable promotional credits offered by ServiceY to be used exclusively for purchases on the Service. Imagine Jane, a savvy shopper, receives \$10 ServiceY Credit for a promotion. She buys a vintage lamp listed at \$50. At checkout, ServiceY Credit reduces her total to \$40, modifying applicable taxes. Jane saves money, but can’t withdraw or transfer the Credit—it only applies to ServiceY purchases.
Vague “To use certain features of our Services, you may be required to create a ServiceX account (an Account id) and provide us with a username, password, and **certain other information** about yourself as set forth in the Privacy Policy …” Certain other information refers to optional details such as a bio, gender, age, location, profile picture, or social link that you may provide when creating a ServiceX account. Imagine signing up for ServiceX, and you’re asked to provide a username, password, and ’certain other information’ like your age and location. Later, you find out that ServiceX uses this data to tailor ads specifically for you, raising concerns about how much they know about you and potential privacy risks.
Vague “… ServiceY reserves the right to discontinue providing Labels to any or all Users at any time and for **any reason** …” Any reason refers to ServiceY’s discretion to stop providing shipping labels without needing to specify a particular cause or justification. Imagine you sell vintage clothes online. ServiceY provides you with prepaid shipping labels. Suddenly, without explanation, they stop offering these labels to you. This means you’ll need to cover shipping costs yourself, impacting your profits. This demonstrates ServiceY’s right to discontinue services for ’any reason,’ affecting your business.
Table 2. Examples of the generated definitions and scenarios for potentially unfamiliar and vague phrases.
Figure 19. The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions. Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The results show a significant effect of the TermSight interface on the user experience for every question. Since the 94% HPDI for each pair (control, treatment) for every question do not overlap with each other, we should expect a significant effect of the treatment on the user experience. E1–E7 refer to the 7 experience questions specified in Appendix F.3.
Figure 20. The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions. The $x$-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05). Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap. This suggests no significant effect of the treatment on the user’s comprehension. Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
Prompt for Obtaining Summary Snippets and Information Snippets Summarize the input section of the terms of service into concise bullet points (less than 12 words) in plain language. When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point. For each bullet point summary, include the full-text reference to the original passage in {} and don’t use "..." to reduce text in the reference. When outputting a reference, don’t change anything in the original text such as spaces and newlines. There can be multiple sentence or paragraph references to a single summary. The references to summary should cover the original text
Example output format: {EXAMPLE OUTPUT}
Input: {INPUT TEXT CHUNK}
Figure 21. Prompt for obtaining Summary Snippets and Information Snippets
Prompt for Classifying Power Classify the input term from a Terms of Service agreement based on the power relationship and benefit between the service and the user. Use the following categories (Bad, Neutral, Good):
- Service: The term grants the service provider disproportionate power or control over the user. It may impose unfair restrictions, obligations, or liabilities on the user, or reduce the user’s rights and autonomy over their data or content.
- Neutral: The term outlines standard procedures, responsibilities, or conditions the user and service have. For example, users take responsibility for the content they post. It neither significantly favors the service provider nor the user, and does not substantially impact the user’s rights.
- User: The term empowers the user by offering clear protections, rights, or benefits, ensuring transparency, and limiting the service provider’s power.
Examples for each category: Service:
- The service can delete specific content without prior notice and without a reason.
- The service can license user content to third parties.
- The service tracks your personal data for advertising
Neutral:
- Users are responsible for the content they post
- Users agree not to use the service for illegal purposes
- Blocking first party cookies may limit your ability to use the service
User:
- You can opt out of targeted advertising
- The service does not sell your personal data
- The service will not allow third parties to access your personal information without a legal basis
Output format in JSON: {"Category": "Service/Neutral/User", "Explanation": "explanation of output" }
Input: {INPUT INFORMATION SNIPPET}
Figure 22. Prompt for classifying power balance of the Information Snippets
Prompt for Classifying Relevance For the input term from a Terms of Service, assign a relevance rating (High/Low) of the input term with respect to the user persona. Output "Low" for low-relevance terms.
Relevance rating:
[High]: The term is directly relevant to the user’s usage of the service or what the user cares about. The term applies to the user persona and is necessary for the user to know.
[Low]: The term is not relevant to the user’s usage of the service or what the user cares about. The term doesn’t apply to the user persona or is not necessary for the user to know.
User Persona: {INPUT USER PERSONA}
Output format in JSON: {"Relevance": "Low/High", "Explanation": "explanation of output" }
Input: {INPUT INFORMATION SNIPPET}
Figure 23. Prompt for classifying the relevance of Information Snippets to the user persona
Prompt for Identifying Unfamiliar Phrases You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of. Jargon refers to domain-specific terminologies that a lay user might not know about.
Example jargon:
- legal jargon: indemnity, arbitration, liability
- copyright licenses: sublicensable licenses, royalty-free licenses
- technical privacy terms: cookies, Ad identifiers, Authentication tokens
Return an empty array if the section does not contain jargon. The extracted word should exactly match the original input text with the same capitalization and sequence of words.
Output format in JSON: {"Jargon": []}
Input: {INPUT TEXT CHUNK}
Figure 24. Prompt for identifying potentially unfamiliar phrases for lay users in a text chunk.
Prompt for Identifying Vague Phrases You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service. A vague term refers to information that is vaguely abstracted without a clear definition provided in the section.
Example Vague terms: information, other, some, third parties, most, generally, personal data, others, general, many, various, might, services, certain information
Return an empty array if the section does not contain vague terms. The extracted word should exactly match the original input text with the same capitalization and sequence of words.
Output format in JSON: {"Vague": []}
Input: {INPUT TEXT CHUNK}
Figure 25. Prompt for identifying vague phrases in a text chunk.
Prompt for Generating Phrase Definitions Use information in the retrieved context to provide a definition of the user-selected phrase or term. Avoid using long sentences. For example, if the user-selected term is "information", define what the term "information" includes and refers to, such as: location data, interaction data, profile data, etc. The output definition should be specific and straight to the point, don’t include language that doesn’t contribute to the definition such as ’in the given context’. Output the string list of reference ids (["ref1", ...]) used to generate the definition under "References". If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for "References".
Examples: {EXAMPLES}
Output format in JSON: {"Definition": "", "References": ["ref1", "ref2", "ref3"]}
Retrieved Context: {RETRIEVED CONTEXT}
Question: What does {INPUT PHRASE} refers to?
Context around the user-selected phrase: {PHRASE CONTEXT}
Figure 26. Prompt for generating in context phrase definitions
Prompt for Generating Scenarios Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase. The scenario/example should be relevant to the below user persona using {an E-commerce platform of used items / a Social Media platform}.
User Persona: {INPUT USER PERSONA}
Output format in JSON: {"Story": ""}
User selected phrase: {INPUT PHRASE}
Context around the user-selected phrase: {PHRASE CONTEXT}
Definition of user selected phrase: {GENERATED DEFINITION}
Figure 27. Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase
Prompt for Generating Answers to User Questions You are an assistant for question-answering tasks. Use information in the retrieved context to answer the user’s question in less than 5 sentences. Output the string list of reference ids (["ref1", ...]) used to generate the definition under "References". If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for "References".
Examples: {EXAMPLES}
Output format in JSON: {"Answer": "", "References": ["ref1", "ref2", "ref3"]}
Retrieved Context: {RETRIEVED CONTEXT}
Question: {USER QUESTION}
User selected phrase: {INPUT PHRASE}
Context around the user-selected phrase: {PHRASE CONTEXT}
Figure 28. Prompt for generating answers to user questions

### v2

@@
 Abstract.
-Terms of Service (ToS) are ubiquitous, legally binding contracts that govern consumers’ digital interactions.
-However, ToS are not designed to be read: they are filled with pages of ambiguous and complex legal terminology that burden potential users.
-We introduce TermSight, an intelligent reading interface designed to make ToS more approachable.
-TermSight offers visual summaries that highlight the relevance and power balance of information in a ToS.
-TermSight also categorizes and simplifies information within the ToS into concise plain-language summaries.
-To aid in reading the original text, TermSight offers contextualized definitions and scenarios for unfamiliar phrases.
-Our within-subjects evaluation of TermSight (N=20) revealed that TermSight significantly reduced the difficulty of reading ToS and increased participants’ willingness to do so.
-We also observed emerging strategies that participants took when interacting with AI-powered features that highlight the diverse ways that TermSight assisted ToS reading.
-large language models, legal documents, reading support tools
-copyright: acmlicensedjournalyear: 2025ccs: Human-centered computing Interactive systems and tools
+Legal contracts govern much of our society, but their specialized language is difficult for non-experts to read.
+While AI has enabled simplification of complex language, legal contracts pose unique challenges because of their connection to readers’ values, ambiguity, and legally binding nature.
+Based on a formative study (N=20) using Terms of Service (ToS) as example contracts to study challenges in contract reading, we developed TermSight, an intelligent reading interface to probe the opportunities and challenges of designing augmentations for legal text.
+TermSight guides readers to relevant clauses with color-coded plain-language snippets of information and contextualizes ambiguous language with definitions and hypothetical scenarios.
+Importantly, TermSight’s features always foreground the original, legally-binding contract text (e.g., linking to associated clauses).
+Our within-subjects study (N=20) demonstrated the opportunities of TermSight in making ToS significantly easier to read and navigate while revealing the challenges of augmenting service contracts such as ToS.
+Legal Contracts, Augmented Reading, Large Language Models
+journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791387isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Applied computing Law
 Figure 1.
-The TermSight system provides guidance and reading support at multiple levels of ToS reading.
+TermSight provides multi-level support for reading Terms of Service (ToS).
 At the contract level, TermSight visualizes the relevance and power balance of content in each policy (1).
-At the document level, TermSight chunks, summarizes, and categorizes content into 1-sentence plain-language summaries (Summary Snippets) highlighted with colors that reflect power and relevance (2).
-Readers can click the Summary Snippet (2) or Highlight Bar (3) to navigate between the summary snippets on the left and the original text on the right.
-At the phrase level, TermSight offers phrase definitions and scenarios for unfamiliar and vague phrases (4).
+At the document level, TermSight chunks, summarizes, and categorizes content into one-sentence plain-language summaries (Summary Snippets) highlighted with colors that reflect power and relevance (2).
+Readers can click the Summary Snippet (2) or Highlight Bar (3) to navigate between the two.
+At the phrase level, TermSight offers phrase definitions and hypothetical scenarios for unfamiliar and ambiguous phrases (4).
+The TermSight system provides guidance and reading support at multiple levels of Terms of Service (ToS) reading.
+The interface contains three panels: a top navigation panel for switching between different policies, the original text on the right, and summaries on the left.
+(1) At the contract level, TermSight visualizes the distribution of content in each policy in the top navigation panel as a horizontal bar plot based on its relevance to the user and power balance (favoring tenant vs.
+landlord).
+(2) Within each policy, TermSight chunks, summarizes, and categorizes content into one-sentence plain-language summaries (Summary Snippets), highlighted with colors that reflect power (red = favor provider, yellow = neutral, green = favor user) and relevance (low saturation = low relevance, high saturation = high relevance).
+The summaries are shown in the left panel, with corresponding highlight bars aligned to the original text on the right.
+(3) Readers can click either a Summary Snippet or its Highlight Bar to navigate between the summary and the source text.
+(4) At the phrase level, TermSight provides definitions and scenarios for unfamiliar or ambiguous terms.
 1. Introduction
-Terms of Service (ToS)—the ubiquitous, rarely-read but legally binding contracts that govern the use of digital services—are widely incomprehensible to potential consumers.
-Despite consumers’ self-reported concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007), and their eventual regret when they encounter issues related to this information (e.g., the selling of their data)  (Good et al., 2005; Obar and Oeldorf-Hirsch, 2020), almost no one reads ToS content fully  (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020).
-There is good reason not to do so: ToS are not designed to be read.
-ToS are filled with legal jargon that can derail general-audience (also referred to as ‘lay-audience’) readers with confusing or under-defined terms (e.g., a ToS stating the service will collect certain ‘information’)  (Ibdah et al., 2021; Tang et al., 2021; Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022).
-ToS also contain more than one policy (e.g., a privacy policy) that in turn each contain reams of boilerplate or semi-boilerplate text.
-Within these pages of text, there might be information relevant to a reader (e.g., an arbitration clause or copyright policy) but such information is lost among the many pages of text  (Srinath et al., 2021; Obar and Oeldorf-Hirsch, 2020).
-Legal scholars have argued that these ambiguous, jargoned, and exceptionally long contracts are designed not to cultivate a shared understanding, but to mitigate liability and establish unilateral control (Kar and Radin, 2019; Benoliel and Becher, 2019; Karanicolas, 2021; Palka, 2023).
-Summarizing and simplifying text is a first step in making ToS more approachable, but resolves only part of the challenge these contracts pose.
-Prior interfaces have leveraged crowdsourcing (TOS, 2024; pri, 2024) and, more recently, AI-based approaches like large language models (LLMs) to extract and process text automatically from ToS (Pan et al., 2023).
-However, these approaches primarily focus on presenting information from pre-defined categories (e.g., user privacy) or on single policies within a ToS (e.g., a privacy policy).
-In reality, ToS contain many policies that cover different information (§3.3.1), and what information a reader needs is dependent on their planned usage of the site (e.g., a seller vs.
-buyer on an e-commerce site).
-We envision a new reading experience for ToS.
-Our goal is to enable readers to find information important to them anywhere within a ToS and scaffold their ability to understand and contextualize this information within their potential use of a service.
-We start by developing a more nuanced understanding of readers’ information needs and challenges when reading ToS through a formative observational study (N=20).
-We found that participants cared about different information based on their envisioned usage of a service (e.g., a content creator will have different needs than a passive consumer on a social media website) and their own personal values (e.g., people had different levels of comfort when it came to privacy and data sharing).
-When participants sought to resolve these information needs, they encountered barriers at every level of granularity within a ToS.
-Within the full ToS contract, participants struggled to navigate nested policies and decide which policies to read.
-Within a specific policy of the ToS (e.g., the privacy policy), participants struggled to surface and interpret relevant information from extended, visually dense, and difficult text.
-When reading the text itself, participants struggled to interpret jargon and phrases that were ambiguous or vague.
-Interestingly, existing affordances (e.g., policy names, section headers with table of contents, and summaries) were often not helpful and at times misleading.
-For example, participants often found that policy names and section headers were unclear and that overview summaries used deceptive friendly language to hide unfavorable clauses.
-We address the specific challenges surfaced by our formative work with a new intelligent reading interface, TermSight.
-TermSight provides multi-level reading guidance and support through contract-level overviews (Power Meter), document-level plain-language summaries (Summary Snippets), and phrase-level definitions with scenarios of potential implications (Phrase Scope).
-To evaluate TermSight, we conducted a counterbalanced within-subject study (N=20) comparing TermSight and a baseline HTML reader during a timed reading task.
-Participants found reading ToS to be significantly easier with TermSight and were more willing to read and spend time on ToS.
-Participants reported that features of TermSight provided guidance and reading support at all granularities of ToS, resolving challenges surfaced in our formative study.
-We also observed interesting and unexpected user behaviors centered around trade-offs between TermSight’s AI-generated features.
-In this paper, we make the following contributions:
+Legal contracts govern much of our society.
+We sign contracts for where we live, who we work with, and, increasingly, for our digital interactions online.
+In order to encompass many possible situations and actions, contracts contain highly specialized language, such as:
+“You grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content.” — Reddit Terms of Service  (85)
+On the one hand, this specialized language allows contracts to articulate the entitlements, obligations, and prohibitions (Sancheti et al., 2023) of the signing parties.
+On the other hand, this specialized language is exceptionally difficult for many people without a legal background to read  (Furth-Matzkin, 2017; Mueller, 1970; Amos et al., 2021).
+Recent AI systems, driven by language models (LMs), have expanded our ability to transform and augment text documents for different readers.
+Past reading systems have simplified scientific papers (Shin et al., 2024), medical research (August et al., 2023), and news articles (Chen et al., 2023) for general audience readers.
+These transformations have also begun to be applied to legal language, by, for example, generating summaries based on predefined categories of rights and responsibilities in contracts (Sancheti et al., 2023; Pan et al., 2023) or through explaining legal concepts (Jiang et al., 2024).
+However, legal text, and specifically contracts, pose unique challenges for simplification that current systems fail to address.
+First, contracts are inherently value-dependent: they are intended to encode and protect the values of the signing parties (Haapio and Passera, 2017; Fiesler et al., 2016; Sancheti et al., 2023), which may lead to varying information needs among individuals.
+Second, legal language is not just complex, but can be ambiguous (e.g., “We may share some of your personal information…”  (Bhatia et al., 2016)).
+While this language provides the linguistic flexibility to manage unforeseen circumstances  (Zhang, 2015), it also derails general audience readers  (Ibdah et al., 2021; Tang et al., 2021; Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022).
+Finally, the literal language of a contract is important: it is the only language that is legally binding  (Cheong et al., 2024).
+Consequently, any text transformations should support rather than replace reading the original text.
+In this paper, we explore augmentations that account for the unique challenges of simplifying legal contracts.
+We focus on Terms of Service (ToS), the ubiquitous but rarely-read contracts that govern the use of digital services (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020).
+Despite consumers’ concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007), these contracts remain difficult to engage with.
+While prior work often focuses on a single policy within the ToS (e.g., a privacy policy  (Nokhbeh Zaeem et al., 2022)), we explore challenges readers face when trying to engage with the full ToS contract.
+In a formative study (N=20) where participants read ToS, we found that these contracts represented the extremes of the difficulties that legal contracts pose more generally (i.e., value-dependent, ambiguous, and legally binding), validating this initial focus.
+The parts of the ToS that were important to participants depended heavily on their envisioned usage and personal values, and this value-dependent information was lost among the many nested policies.
+ToS also contained complex and ambiguous terms that confused readers.
+Finally, existing features to support contract reading (e.g., policy names and section headers) often did not accurately represent the actual text and were not legally binding.
+[^1] ToS often include clauses such as “Headings are used in these Terms for reference only and will not be considered when interpreting them”  (85).
+In response, we designed a novel intelligent reading interface for ToS contracts, TermSight.
+In contrast to prior work that provides overview summaries detached from the original text  (99; 83; S.
+Pan, T.
+Hoang, D.
+Zhang, Z.
+Xing, X.
+Xu, Q.
+Lu, and M.
+Staples (2023)), we explore augmentations that support contract reading while acknowledging the unique characteristics of legal text.
+TermSight provides visual overviews of policies within a ToS (Power Meter, Figure 1.1), guides readers to relevant clauses with color-coded plain-language snippets of information (Summary Snippets, Figure 1.2), and contextualizes unfamiliar or ambiguous language with definitions and scenarios of potential implications (Phrase Scope, Figure 1.4).
+Importantly, TermSight’s features always foreground and link to the original legal text.
+To explore the opportunities and challenges of designing contract reading interfaces, we conducted a counterbalanced within-subjects study (N=20) using TermSight and a baseline HTML reader.
+Compared to prior focus on simplified reading tasks (e.g., reading a single clause (Robinson and Zhu, 2020; Lee et al., 2025) or policy overview (Tabassum et al., 2018; Korunovska et al., 2020)), participants engaged with real ToS contracts with 10+ policies.
+In the study, participants found reading ToS to be significantly easier with TermSight while also being more willing to read the original ToS.
+Participants reported that TermSight provided multi-level guidance and support while letting them drill down into clauses within the ToS important to them.
+We also observed that participants often used TermSight’s generated text (e.g., Summary Snippets) as an entrance to, or a verification of, the original ToS text.
+At the same time, our findings revealed the broader limitations of technological solutions like TermSight on facilitating a deeper understanding of ToS due to the overwhelming amount of potentially relevant information.
+We end by discussing the implications of our findings for inspiring future contract interfaces and regulatory innovation.
+In summary, this paper makes the following contributions:
   1. (1)
-    We characterize the challenges readers face when trying to make sense of ToS contracts.
-Readers face challenges at all granularities of a ToS, from determining which policies to read to being able to resolve individual ambiguous terms within the text.
+    We characterize the challenges of contract reading, specifically ToS, including determining which policies to read, resolving ambiguous terms, and navigating misleading organization within the text.
   2. (2)
-    We design and develop an intelligent reading interface, TermSight, for helping readers make sense of ToS contracts at the contract, document, and phrase levels.
+    We develop a novel intelligent reading interface, TermSight, to explore the design opportunities and challenges for helping readers engage with ToS while focusing on the original, legally binding language.
   3. (3)
-    We collect empirical evidence from our user study (N=20) demonstrating the potential value of the interaction affordances of TermSight.
+    We collect empirical evidence from our user study demonstrating the values of TermSight’s features in finding relevant information, navigating ambiguous or complex language, and associating generated text with original text, while revealing the challenges of improving understanding of ToS.
 2. Background and Related Work
-2.1. Service Contracts and The Duty to Read
+2.1. Service Contracts
 A Terms of Service (ToS) is a type of standard form contract consisting of a body of policies and conditions outlined by a service provider, dictating the rules and expectations for both the service provider and the consumer.
-These contracts have evolved to be more legally comprehensive compared to earlier ToS  (Amos et al., 2021), encompassing a wide range of provisions such as copyright, privacy, returns, acceptable use, and service-specific terms.
-By clicking agree to a ToS, a legally binding contract is formed  (Dickens, 2007).
-U.S.
-contract law is premised on the doctrine of the ‘duty to read’ where contracting parties are legally responsible for reading and understanding the contract  (Benoliel and Becher, 2019).
-However, the ‘duty to read’ was established in the era when contracts were short and based on mutual negotiation  (Kar and Radin, 2019).
-Modern ToS often have important terms buried within a large body of text and documents  (Srinath et al., 2021).
+While prior work in HCI mainly focused on single policies within the ToS (e.g., a privacy policy  (Nokhbeh Zaeem et al., 2022)), these contracts encompass a wide range of provisions such as copyright, privacy, returns, acceptable use, and service-specific terms  (Amos et al., 2021).
+Important terms are often buried within this large body of documents  (Srinath et al., 2021).
+ToS can vary in content (Liu et al., 2021; Fiesler et al., 2016) and structure (Ibdah et al., 2021) across services.
 The language used to write ToS can also require advanced reading capability  (Srinath et al., 2021; Amos et al., 2021) and contain unfamiliar (Ibdah et al., 2021; Tang et al., 2021) or ambiguous terminology (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022).
-Due to the cognitive load associated with reading ToS, most consumers do not read modern ToS (Obar and Oeldorf-Hirsch, 2020; Bakos et al., 2014), despite the binding nature of ToS and consumers’ self-reports indicating concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007).
-ToS can also vary in content (Liu et al., 2021a; Fiesler et al., 2016) and structure (Ibdah et al., 2021) across services, leading to false consumer expectations.
-This widespread disengagement highlights what Kar and Radin described as a ”paradigm slip” in contract law, where agreements once premised on shared understanding and mutual consent are now reduced to unilateral boilerplate ”pseudo-contracts”, eroding fundamental principles of contract law and personal autonomy  (Kar and Radin, 2019; Palka, 2023; Benoliel and Becher, 2019).
-Our formative study complements prior findings by aggregating and uncovering consumers’ varying information needs and additional challenges encountered when navigating and reading policies within a ToS.
+Most consumers do not read modern ToS (Obar and Oeldorf-Hirsch, 2020; Bakos et al., 2014), despite the binding nature of ToS (Dickens, 2007) and consumers’ self-reports indicating concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007).
+This widespread disengagement highlights what Kar and Radin described as a “paradigm slip” in contract law, where agreements once premised on shared understanding and mutual consent are now reduced to unilateral boilerplate “pseudo-contracts”, eroding fundamental principles of contract law and personal autonomy  (Kar and Radin, 2019; Palka, 2023; Benoliel and Becher, 2019).
+Our formative study further supplements prior findings by revealing the challenges readers face when navigating the entire ToS and the presence of misleading navigational affordances (§3).
 2.2. Augmenting Terms of Service
-There are two major threads of work on improving the readability of ToS that we cover below: changing how ToS are designed or presented (§2.2.1) and providing summaries of ToS (§2.2.2).
-2.2.1. Stylistic Manipulations
-To help readers better understand the implications of using a service, prior work has investigated design recommendations for service providers to make documents within ToS more accessible  (Kay and Terry, 2010; Habib et al., 2021; Taber et al., 2020).
-Kay and Terry (2010) proposed Textured Agreements that use typographic manipulation, pull quotes, vignettes, and iconic symbols to improve reader attention and comprehension on privacy policy compared to plain text.
+Prior work has investigated new designs for readers navigating ToS  (Kay and Terry, 2010; Habib et al., 2021; Taber et al., 2020).
+Kay and Terry (2010) proposed Textured Agreements that use typographic manipulation, pull quotes, vignettes, and iconic symbols to improve reader attention and comprehension of privacy policy compared to plain text.
 Habib et al.
@@
 Taber et al.
-(2020) proposed crowdsourced sentiment highlighting of sentences in ToS, providing visual cues to guide readers toward sentences in ToS with negative sentiment.
+(2020) proposed crowdsourced sentiment highlighting of sentences in ToS.
 Design recommendations on styling user agreements have also been proposed  (Kay and Terry, 2010; Habib et al., 2021).
-While helpful in navigating contracts, these design changes do not simplify the content itself and do not take into consideration the fact that readers have different information needs when reading a ToS.
-2.2.2. Summarization
-Work has also proposed helping consumers gain an overview of specific policies within a ToS, often the privacy policy (PP), without requiring reading the original document.
-Platform for Privacy Preferences (P3P) is an early attempt to standardize privacy policies by allowing service providers to submit privacy policies in a machine-readable format  (Reagle and Cranor, 1999).
+Work has also explored helping consumers gain an overview of specific policies within a ToS, often the privacy policy (PP), without requiring reading the original document.
+Platform for Privacy Preferences (P3P) was an early attempt to standardize privacy policies by allowing service providers to submit privacy policies in a machine-readable format  (Reagle and Cranor, 1999).
 Tools such as Privacy Bird build atop the P3P format to provide warnings when a site’s PP does not match consumers’ privacy preferences  (Cranor et al., 2006).
-However, P3P is rarely adopted by service providers  (Cranor et al., 2008) and no longer supported  (Microsoft, 2016).
-Work has also proposed the use of a single-page summary for end-user license agreements (EULAs)  (Good et al., 2005, 2007), comic-based summaries for privacy policies  (Tabassum et al., 2018), and Privacy Policy Nutrition Labels (PPNL) that label privacy policies in a table format  (Kelley et al., 2009; Reinhardt et al., 2021) to help users gain an overview of a policy.
-Many earlier tools required buy-in from service providers to design better ToS.
-Other work, though, has focused on providing an overview summary of a ToS automatically or via crowdsourcing.
-For example, ToS;DR leverages volunteers to label and summarize terms in the privacy policy and the home page of ToS  (TOS, 2024).
-Works like CLAUDETTE automatically detect a predefined set of potentially unfair clauses from ToS  (Lippi et al., 2019; Guarino et al., 2021; Braun and Matthes, 2021; Braun et al., 2017).
+However, P3P was not widely adopted by service providers  (Cranor et al., 2008) and is no longer supported  (Microsoft, 12/15/2016).
+Work has also proposed the use of a single-page summary for end-user license agreements (EULAs)  (Good et al., 2005, 2007), comic-based summaries for privacy policies  (Tabassum et al., 2018), and Privacy Policy Nutrition Labels (PPNL) that label privacy policies in a table format  (Kelley et al., 2009; Reinhardt et al., 2021).
+Other work has focused on providing overview summaries automatically or via crowdsourcing.
+ToS;DR uses volunteers to label and summarize terms in the privacy policy and the home page of ToS  (99).
+Works like CLAUDETTE automatically detects a predefined set of potentially unfair clauses from ToS  (Lippi et al., 2019; Guarino et al., 2021; Braun and Matthes, 2021; Braun et al., 2017).
 PrivacyCheck (Zaeem et al., 2018; Nokhbeh Zaeem et al., 2020, 2022), PrivacyGuide (Tesfay et al., 2018), and Polisis  (Harkous et al., 2018) extract terms relevant to a list of privacy-related questions and criteria and generate a grading for each.
-Sathyendra et al.
-extract opt-out choices from privacy policies  (Sathyendra et al., 2017; Bannihatti Kumar et al., 2020).
-Relatedly, some works evaluate the completeness of privacy policies  (Costante et al., 2012) and the GDPR compliance of extracted clauses (Liu et al., 2021b).
-More recently,  Pan et al.
-(2023) investigated generating privacy nutrition labels from a privacy policy.
-Work has also attempted to generate an aggregated summary based on the extracted information from a privacy policy  (Tomuro et al., 2016; Keymanesh et al., 2020).
-TermSight extends this rich prior work by combining the benefit of plain language summaries with interactive affordances that enable readers to navigate the original ToS text, helping readers navigate and read the original body of policies within ToS beyond privacy policies.
-2.3. Intelligent Reading Interfaces
-TermSight takes inspiration from a rich body of work on intelligent reading interfaces that support readers in navigating documents and clarifying information.
-Prior work has explored helping readers navigate documents and locate relevant information in scholarly documents.
-For example, work has highlighted key information in research papers, such as objectives, results, and methods, to support skimming  (Fok et al., 2022).
-Also in research papers, ScholarPhi identifies position-sensitive definitions of terms and symbols, enabling readers to track their meaning across contexts  (Head et al., 2021).
-Talk to Papers allows readers to ask questions about research papers directly, while Qlarify supports clarification by letting readers expand text spans in paper abstracts to generate explanations  (Zhao and Lee, 2020; Fok et al., 2024a).
-In medical research papers, Paper Plain provides a list of guiding questions to help readers navigate important information in a medical paper (August et al., 2023).
+Past work has relied on either service providers or volunteers to make stylistic changes to contracts or automated tools to provide a summary external to the contract.
+In this paper, we explore new ways of combining these two threads to enable readers to navigate and read the original ToS made possible by the powerful text transformation capabilities of language models.
+2.3. Augmented Reading Interfaces
+Prior work has explored augmenting other types of documents, such as research papers (Lo et al., 2024) and news articles (Chen et al., 2023), with generative features.
+Work has provided faceted highlighting in research papers to support skimming  (Fok et al., 2023) and context-sensitive definitions of terms and symbols  (Head et al., 2021) to support reading.
+Also in research papers, Qlarify (Fok et al., 2024a) enabled readers to expand paper abstracts with generative explanations.
+Systems have also provided automated question answering for research papers (Zhao and Lee, 2020) and medical papers (August et al., 2023).
+Shin et al.
+(2024) transformed research papers into summaries of design implications.
+Newman et al.
+(2024) developed tools that synthesized multiple research papers to help researchers conduct literature reviews.
 In the business domain, Marco allows readers to search and ask questions across collections of business documents  (Fok et al., 2024b).
 In the news domain, Chen et al.
-(2023) introduced Marvista, a reading interface that identifies the most summative portions of news articles based on a reader’s time constraints or topic in the news article.
-Compared to prior reading interfaces that have focused on research papers, news articles, or business documents, TermSight focuses on helping general audience readers make sense of legal documents, namely, ToS.
-Legal documents, and specifically ToS, present challenges different from prior reading contexts.
-In contrast to other knowledge-intensive settings (e.g., scientific or medical text), general audience readers in the legal setting are faced with language not designed to communicate information.
-Readers then struggle to (1) determine what text to read within and across policies, (2) define and contextualize vague language, and (3) navigate existing affordances that are themselves ambiguous or incomplete.
-Prior intelligent reading interfaces and context often assume that there is a single document (August et al., 2023; Fok et al., 2022; Chen et al., 2023), readers have consistent information needs  (August et al., 2023), readers know what they need to read or ask  (Fok et al., 2022; Zhao and Lee, 2020; Fok et al., 2024b), or readers can make sense of term definitions  (Head et al., 2021; August et al., 2023; Fok et al., 2024a).
-These existing affordances, while powerful, do not fully support readers’ needs and challenges encountered when reading digital contracts such as ToS.
-TermSight takes inspiration from prior works to explore the design of reading interfaces to help readers surface and interpret relevant information from legal documents.
+(2023) introduced Marvista, a reading interface that identifies the most summative portions of news articles.
+In contrast to other documents explored previously, contracts—and specifically ToS—pose unique challenges to readers.
+Contracts directly connect with a reader’s personal values: one reader might prefer stronger privacy safeguards while another might care about licenses  (Haapio and Passera, 2017; Fiesler et al., 2016; Sancheti et al., 2023).
+Contract language is also meant to be interpreted; this means there are often ambiguous terms or incomplete information (e.g., the definition of ‘information’ when discussing personal information collection), with the assumption that the reader has the requisite knowledge to fill in missing information  (Bhatia et al., 2016).
+Finally, contract language itself is the only legally binding text (Cheong et al., 2024), meaning any text transformations must maintain and display a tight connection with the original text.
+We take inspiration from past augmented reading systems to investigate what features can support readers in the unique context of contract reading.
 3. Formative Study
-To better understand the challenges readers face when trying to make sense of the Term of Service and their information needs, we conducted a formative study to answer the following research questions:
+We first establish the challenges general-audience readers face when reading a legal contract.
+We focus on Terms of Service (ToS) because ToS are legally binding contracts that are signed routinely in digital interactions.
+While ToS are rarely read completely (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020), past legal work suggests that these contracts represent many of the broader issues faced in legal contracts today (Obar and Oeldorf-Hirsch, 2020; Kar and Radin, 2019) and therefore present an ideal setting to explore possible reading features.
+Our formative study was guided by the following research questions:
 RQ1: What information do readers want from a ToS?
@@
 3.1. Methodology
-3.1.1. Study Procedure:
-The entire study took 40 minutes and participants were compensated 10\$.
+3.1.1. Study Procedure
+The entire study took 40 minutes, and participants were compensated 10\$.
 The study was approved by the Institutional Review Board (IRB) office at our organization.
-Participants initially completed a demographics form and reported their past experiences interacting with ToS.
-Following the form, participants were asked to imagine as if they were a first-time user of the assigned service registering for the service and were provided 15 minutes to go through the ToS.
+Participants first reported their past experiences interacting with ToS.
+Next, participants were asked to imagine as if they were a first-time user registering for a randomly assigned service and were provided 15 minutes to go through the ToS while speaking aloud any challenges.
 Participants were free to navigate to different policies within the ToS.
-While reading, participants were instructed to speak aloud any challenges they faced or thoughts about the ToS.
-After reading the ToS, we conducted semi-structured interviews with participants about their experiences and challenges with reading ToS.
-Lastly, we asked the participants to envision how computational supports could aid their reading.
-All interview sessions were transcribed.
-The semi-structured interview questions are listed in Appendix E.
+After reading the ToS, we conducted semi-structured interviews about participants’ experiences and challenges with ToS reading, along with opportunities for computational support.
+The interview questions are listed in Appendix D.
 3.1.2. Analysis
@@
 The lead researcher thoroughly explored the data, noted interesting features, systematically coded the data, and iteratively compared the codes to generate the initial themes.
-Then, the lead researcher discussed with other members of the research team to make sure the themes fit the data and further developed the themes.
+Through weekly meetings over the course of two months, the lead researcher discussed with other members of the research team to make sure the themes fit the data and further developed the themes.
+This included iterating on the descriptions of the themes and merging related themes.
 3.1.3. Materials
-Participants were randomly assigned one of the 10 ToS in the study.
-We selected ToS from 10 services spanning social media, e-commerce, video streaming, and internet services (Social Media: Reddit, Facebook, Instagram, Twitter; E-commerce: Amazon, eBay; Video: Youtube, Netflix; General: Google, Yahoo).
-These 10 services were selected from the top 20 visited sites on semrush.com (2024/05) where we selected the top 4 social media services, top 2 E-commerce services, top 2 video platforms, and top 2 internet services.
-We over-sampled social media services because they are the most represented service category in the top 10 visited services.
-In the study, participants were first directed to the main ToS and were allowed to navigate across any other policies.
+Participants were randomly assigned one of the 10 services’ ToS [^2] Social Media: Reddit, Facebook, Instagram, Twitter; E-commerce: Amazon, eBay; Video: Youtube, Netflix; General: Google, Yahoo.
+The 10 services were selected from the top 20 visited sites on semrush.com (2024/05), where we sampled the top 4 social media services, top 2 E-commerce services, top 2 video platforms, and top 2 internet services.
+We oversampled social media services because they were the most represented service category in the top 10 visited services.
 3.1.4. Participants
-We recruited 20 participants through Prolific.
-Each participant was randomly assigned one of the 10 selected ToS.
+We recruited 20 participants through Prolific (Male: 6, Female: 14).
 All participants were over 18, fluent in English, and located in the US.
-14 participants self-identified as female and 6 as male.
-3 participants were between the ages of 18-25, 9 were between 26-35, 4 were between 36-45, 3 were between 46-55, and 1 was 67.
-3 participants had never read or skimmed a ToS before and 17 participants had read or skimmed at least 1 ToS before.
+Participants’ age ranged from 20 to 67 ($\mu=35.75,\sigma=12.24$).
+3 participants had never read or skimmed a ToS, and 17 participants had read or skimmed at least 1 ToS before.
 We did not require participants to have prior experience in reading the ToS.
-All the participants either had not read or did not remember reading the ToS they were assigned.
+All the participants had not read or did not remember reading the ToS they were assigned.
 3.2. RQ1: What information do readers want from a ToS?
-3.2.1. Participants have diverse information needs based on usage of the service and personal value
+3.2.1. Participants have diverse, value-dependent information needs when reading ToS
 All 20 participants described the need to know what control and rights the service had over the user, what the service was allowed to do, and how the user could be negatively impacted.
-Additionally, 12 participants highlighted the need to know their (as opposed to the service’s) rights and control.
-Furthermore, 4 participants expressed wanting to know what they were allowed or not allowed to do on the platform.
-Participants also noted more specific and diverse information that they cared about such as data collection and usage (11), intellectual property right over user content (8), purchasing and returns (7), account deletion by the service (5), arbitration and liabilities (3), content moderation (2), and exposure to offensive content or misinformation (1).
-10 participants explicitly mentioned how their information needs and priorities depended on usage and personal values.
+While all participants shared a desire to know their general rights, different participants also noted more specific information needs, such as data collection and usage (11), intellectual property rights (8), purchasing and returns (7), account deletion by the service (5), arbitration and liabilities (3), content moderation (2), and exposure to offensive content or misinformation (1).
+10 participants explicitly mentioned how their information needs from the ToS and priorities depended on service usage and personal values.
 For example, P1 described that “I don’t think any information is irrelevant.
@@
 Our findings revealed that participants encountered challenges at the contract (§3.3.1), document (§3.3.2), and phrase (§3.3.4) level.
-In addition, participants highlighted how existing navigational and reading affordances were often ineffective, misleading, and failed to support meaningful comprehension (§3.3.3).
+These challenges reflected the complexity and ambiguity of legal language.
+In addition, participants highlighted how existing navigational and reading affordances were often misleading and failed to support meaningful sensemaking of the actual ToS text (§3.3.3).
 3.3.1. Contract Level: Participants struggled to navigate nested policies and decide which policy to read.
-11 participants mentioned that it’s unclear to them what sub-policies are included when they are navigating the ToS because sub-policies are often hyperlinked throughout the documents.
-There often lack a centralized list of what policies are included as part of ToS.
-7 participants described the challenge of navigating across these nested policies as getting ‘lost in a whole tree of information’ (P18).
+Eleven participants mentioned that it’s unclear to them what sub-policies are included when they are navigating the ToS because sub-policies are often hyperlinked throughout the documents.
+There often lacks a centralized list of what policies are included as part of ToS.
+Seven participants described the challenge of navigating across these nested policies as getting ‘lost in a whole tree of information’ (P18).
 All 20 participants described relying on policy names to help decide whether to explore a policy.
-Yet, 17 participants described that the names of the policies alone did not provide a clear mental model of what was in the sub-policies and whether there was important information they should know about.
+Yet, 17 participants noted that the names of the policies alone did not provide a clear mental model of what was in the sub-policies and whether there was important information they should know about.
 For example, P14 mentioned that: “When it’s hyperlinked like ‘our rules and policies’ and ‘privacy policy’, it’s difficult to determine if I need to know something from the link.” 8 participants explicitly expressed their reluctance to click into the linked polices due to the lack of awareness of relevant information and the challenge of finding relevant information from a hyperlinked document.
 3.3.2. Document Level: Participants lacked guidance and struggled to surface relevant information from extended, visually dense, and obfuscating text within a policy.
-17 out of 20 participants described their reading behavior as skimming the ToS to see if any information caught their eye.
-6 participants noted that they were not sure what they should look for in the ToS and might have unintentionally missed important information.
-P12 further commented how it’s not always obvious if the information was relevant at first sight: “A lot of it seemed irrelevant until I started reading in a little bit more.”
-18 participants highlighted the overwhelming length of the document, especially when considering the nested policies.
-These participants expressed concerns about missing important terms buried in the text.
-For example, P7 pointed out that “When all the important information gets hidden behind walls of text, it doesn’t really incentivize anyone to read it, and it gets hidden.” Reflecting P7’s comment, 13 participants described the text as visually dense and difficult to parse when skimming.
-These participants further noted how important information can be positioned at the very bottom of an extended policy that was hard to discover.
-For example, P10 noticed how information about returns and cancellations of an E-commerce platform was positioned under the 13th section of the ToS with the title ‘Additional Terms’.
-Additionally, 3 participants described how the use of friendly language actually obfuscated the ToS and made skimming more challenging.
+Most (17/20) participants primarily skimmed the ToS.
+Six participants noted that they were not sure what they should look for in the ToS and might have unintentionally missed important information.
+P12 further commented how it’s not always obvious if the information was relevant at first sight: “A lot of it seemed irrelevant until I started reading in a little bit more.” Often, this inability to find information was connected to the overwhelming length of the document (18/20).
+Interestingly, many (13/20) participants also noted how text positioning seemed intentionally obfuscating: important information was positioned at the bottom of an extended policy.
+For example, P10 noticed how information about returns and cancellations of an E-commerce platform was positioned under the 13th section with the title ‘Additional Terms’.
+Similarly, three participants described how the use of friendly language tended to obfuscate information and made skimming more challenging.
 For example, P15 noticed language such as ‘We only use data to make [the service] a better place’ when reading a privacy policy.
 Yet, she later found out how user information can be used for targeted ads.
-As a result, P15 described that it’s easy for lay people to misinterpret the language and miss important terms: “If someone’s just skimming the terms of service and is not skeptical at all.
-Then, you could read over that and think that means they’re not using your data for targeted ads or stuff.”
 3.3.3. Document Level: Participants found existing affordances to be ineffective and misleading
-Some of the ToS participants read contained navigational and reading affordances like an interactive table of contents with section headers and overview summaries.
+Some of the ToS participants read contained navigational and reading affordances, such as interactive table of contents with section headers and overview summaries.
 However, 12 participants noted that, similar to policy names, section headings within policies were vague and not descriptive of the actual content.
-As a result, it was unclear what information was contained in a section by only looking at the section header.
 For example, when skimming the ToS, P10 skipped the ‘Content’ section but later realized that the section was about intellectual property rights over user-generated content: “It just says ‘content’ which is pretty vague when they’re talking about getting exclusive intellectual rights.
@@
 So people would skip right past it.
-I did the first time.”
-In addition to ineffective navigational affordances such as vague policy names (17/20) and section headers (12/20), 3 participants accessed a sub-policy with a summary at the top.
+I did the first time.” In addition to ineffective navigational affordances such as vague policy names (17/20) and section headers (12/20), three participants accessed a sub-policy with a summary at the top.
 All three participants (P4, P14, P15) described using the summaries to gain an overview at the start.
@@
 They don’t say that in the summary” (P4).
-They also expressed concern that readers might only glance at the summary and mistakenly think it covers the entire policy.
 3.3.4. Phrase Level: Participants struggled to contextualize unfamiliar or ambiguous terminology
 When reading individual sections in a policy, 12 participants pointed out unfamiliar legal terminology (e.g., Arbitration, Indemnity, or class lawsuit).
-These participants described how the terminology hindered their understanding and motivation to read the section.
-For example, P8 described skipping sections that were full of unfamiliar terms.
-Additionally, participants noted how challenging it was to interpret the meaning of a word or phrase in the context of their situation: “I can recognize these are legal terms, but I don’t necessarily know what that means for me” (P10).
-P6 also described how it’s hard to envision scenarios in which things go wrong beforehand from the ToS.
-8 participants noticed ambiguous or vague terminologies that made it challenging for them to understand the implications of signing.
-For example, P15 mentioned that “I really think the big difficulty is the vagueness of the language and the constancy of exceptions that are vague which gives them a lot more leeway.” More specifically, 5 participants noticed vague terms around the service’s information collection and sharing practices such as: ‘we share you data with 3rd parties’, ‘track information about your interaction’, and ‘retain certain information about you’.
-4 participants noted vague terms around how the service can use user-generated content.
-For example, P3 and P14 pointed out that it’s unclear what “use, distribute, modify, and copy user content” implicates and what are examples of what the service can do with user-generated content.
-3.4. Design Guidelines:
+These participants noted the difficulty to interpret the meaning of legal jargon in the context of their situation: “I can recognize these are legal terms, but I don’t necessarily know what that means for me” (P10).
+In a similar vein, eight participants noticed ambiguous terminology that made it challenging for them to understand the implications of signing (e.g., ‘we share your data with 3rd parties’, and ‘retain certain information about you’).
+P15 mentioned that “I really think the big difficulty is the vagueness of the language and the constancy of exceptions that are vague, which gives them a lot more leeway”.
+3.4. Design Guidelines
 We propose four design guidelines based on our observations that readers faced interrelated barriers at all levels of ToS reading:
@@
 DG2: Document level: Help readers identify potentially relevant information within a policy.
-DG3: Phrase level: Help readers read and interpret original text with technical language, jargon, and vague phrases.
+DG3: Phrase level: Help readers read and interpret original text with technical language, jargon, and ambiguous phrases.
 DG4: All summaries should be tightly coupled with the text used to generate them.
 When reading any generated overviews, readers should be able to retrieve the original document text in at most one click.
-4. The TermSight System:
-We reify these design guidelines in TermSight, an augmented reading interface that supports user sense-making of ToS with features at each level of granularity of a ToS:
+4. The TermSight System
+We reify these design guidelines in TermSight (Figure  1), an augmented reading interface for exploring opportunities to make ToS contracts more approachable with features at every granularity:
 Power Meter: A contract-level visualization of the relevance of information within a document and the distribution of power between the user and the platform that information represents.
-Summary Snippets: 1-sentence plain language summaries of information chunks at the document-level.
+Summary Snippets: One-sentence plain language summaries of information chunks at the document-level.
 Snippets are rendered with color and saturation that represent power and relevance to the user, similar to the Power Meter.
@@
 We adopted an iterative design approach in developing TermSight.
-8 participants evaluated an early prototype of TermSight, and their feedback informed the final design.
+Eight participants evaluated an early prototype of TermSight, and their feedback informed the final design.
 Additional details about the iterative design study, results, and design changes can be found in Appendix  A.
-4.1. Design Abstractions:
+4.1. Design Abstractions
 The design of TermSight is grounded in two design abstractions: 1) Information Snippets, and 2) Power and Relevance classification.
-These abstractions serve as foundational building blocks for TermSight’s features that support contract and document-level sensemaking as illustrated in Figure  6.
 4.1.1. Information Snippets
-Our formative study reveals that the information readers cared about in ToS is often buried within extended and visually dense text.
-Consequently, prior works that support document understanding through document-level summaries (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023) can fall short in helping readers surface pieces of information buried in large sections.
-To address this limitation, we use Information Snippets as a basic building block for TermSight.
+Our formative study revealed that the information readers cared about in ToS can be buried within visually dense text.
+To enable sense-making of finer-grained pieces of information as opposed to an entire section, we use Information Snippets as the unit of interaction in TermSight.
 An Information Snippet is a continuous span of original text that shares the same topic and can be summarized in a single sentence of up to 12 words (detailed in Section 5.2).
 These Information Snippets form the basis of the Summary Snippets (§4.3.1).
-Information Snippets are non-overlapping and together reconstruct the original document’s content (Examples in Figure  2).
-Figure  6 illustrates the process of how TermSight divides a paragraph into five Information Snippets, each paired with its Summary Snippet.
-The idea of designing interactions and sense-making around Information Snippets rather than entire sections was inspired by prior works that modularize and ‘objectify’ tools (Bier et al., 1993; Bederson et al., 1996), attributes (Xia et al., 2016), and AI agent’s memory (Huang et al., 2023) to enable interaction flexibility, interaction specificity, and direct manipulation.
-Figure 2.
-Design conceptualization of Power and Relevance: Power refers to the degree to which a snippet grants control to the service provider or the user.
-Relevance refers to whether or not the snippet is relevant to the user’s intended usage of the service and personal value.
-Examples of Information Snippets, TermSight classified under each category of power, are shown.
+Information Snippets are non-overlapping and together reconstruct the original document’s content.
+Figure  5 illustrates the process of how TermSight divides a paragraph into five Information Snippets, each paired with its Summary Snippet.
+The idea of designing interactions and sense-making around finer-grained Information Snippets was inspired by prior works that modularize and ‘objectify’ tools (Bier et al., 1993; Bederson et al., 1996), attributes (Xia et al., 2016), and AI agent’s memory (Huang et al., 2023) to enable interaction specificity and direct manipulation.
 4.1.2. Power and Relevance Classification
-We define Power and Relevance as two qualities of an information—and thus a summary—snippet.
+Our formative study revealed participants’ information needs regarding the rights and control held by the service and the user.
+These needs may also vary depending on the intended service usage.
+To provide information scent  (Pirolli and Card, 1999) for snippets of information that a reader may care about, we reify participants’ information needs by defining Power and Relevance as two qualities of an information—and thus a summary—snippet.
 Power is the degree to which a snippet’s text grants control to the Service Provider or the User (Categories: Service, Neutral, or User).
@@
 As shown in Figure  2, we visually encode Power and Relevance in TermSight with a combination of hue (Red, Yellow, Green) and saturation (High, Low).
+We constrained relevance to High and Low because our iterative design study revealed that more color–saturation pairs (i.e., 3 hues for Power x 3 saturations for Relevance) were difficult to interpret (Appendix  A).
 Figure 2.
-Power Meter: a visualization of the distribution of the power and relevance of Information Snippets within a document (a).
-Users can gain an overview of the policies within the ToS (b).
-When hovering over the Power Meter, the list of Summary Snippets of the corresponding policy is shown (c).
-Clicking the blue backquote would allow users to view the referenced Information Snippet (d).
-4.2. Power Meter: Policy level visualization of power and relevance
-While prior works have focused on a single policy such as the Privacy Policy  (Nokhbeh Zaeem et al., 2022; Tesfay et al., 2018) or the main ToS page (Lippi et al., 2019; Guarino et al., 2021), Terms of Services include many policies (e.g., content policy, community guidelines, return policy, and more).
-To help users form mental models of each sub-policy, TermSight provides information scent and preview to help users gain an overview of the policies within ToS to decide which policies to read (DG1).
-4.2.1. Providing information scent of a policy
-In TermSight, the policies that are part of the ToS agreement are placed in the top navigation panel where users can navigate to different policies.
-Each policy in the top navigation panel or hyperlinked within the document is accompanied by a Power Meter: a horizontal bar visualization representing the distribution of Information Snippets in a document, with colors denoting power and relevance.
-Shown in Figure 3, Power Meter provides information scent (Pirolli and Card, 1999) of the power distribution of Information Snippets within the policy and the fraction of Information Snippets that are potentially relevant to the user.
-4.2.2. Providing a preview of the document
-Hovering over a Power Meter reveals a preview popup containing a list of Summary Snippets, allowing users to skim and gain an overview of a policy’s Information Snippets without having to be redirected to a new page (Figure 3c).
-Each Summary Snippet in the popup includes a clickable quote icon that allows users to view the referenced Information Snippet (DG4).
-4.3. Summary Snippets: Plain language summaries of discrete information snippets
-To help users surface and interrogate snippets of information buried in a wall of text (DG2) and make original text more accessible (DG4), TermSight provides Summary Snippets accompanied with Highlight Bars for seamless navigation between the summary and original text.
+Design conceptualization of Power and Relevance.
+Power refers to the degree to which a snippet grants control to the service provider or the user.
+Relevance refers to whether or not the snippet is relevant to the user’s persona.
+The figure shows how TermSight defines and visualizes power and relevance using color.
+Power indicates the degree to which a snippet grants control to the service provider or the user (red = favor provider, yellow = neutral, green = favor user).
+Relevance indicates whether a snippet is pertinent to the user’s persona, which is based on intended service usage and personal values (low saturation = low relevance, high saturation = high relevance).
+The figure also provides examples.
+A provider-favoring clause: ‘This license includes the right for us to use your content to train machine learning models…’; A neutral clause: ‘Avoid posting illegal content’; and a user-favoring clause: ‘Our licensees cannot display public content deleted by the users…’.
+4.2. Power Meter: Visualization of the Distribution of Power and Relevance
+In TermSight, the policies that are part of the ToS are placed in the top navigation panel.
+To help users form mental models of each policy and decide which policies to read (DG1), each policy in the top navigation panel or hyperlinked within the document is accompanied by a Power Meter: a horizontal bar visualization representing the distribution of Information Snippets in a document, with colors denoting power and relevance (Figure 3).
+Our iterative design study further revealed the need to contextualize Power Meter with concrete previews (Appendix  A).
+Consequently, hovering over a Power Meter reveals a policy preview containing a list of Summary Snippets (Figure 3b).
+Each Summary Snippet in the popup includes a clickable icon (Figure 3c) that allows users to view the referenced Information Snippet (DG4).
+Figure 3.
+Power Meter visualizes the distribution of power and relevance of Information Snippets within a policy (a).
+On hover, a preview of Summary Snippets is shown (b) with options to view the referenced Information Snippets (c).
+The figure visualizes the Power Meter.
+(a) The Power Meter is a horizontal bar that visualizes the distribution of power and relevance of Information Snippets within a policy.
+(b) When a user hovers over the Power Meter, the system displays the list of Summary Snippets for that policy.
+(c) Each Summary Snippet is paired with a blue backquote icon that links to the referenced Information Snippet.
+4.3. Summary Snippets: Plain language Summaries of Information Snippets
+To help users surface relevant information within a document (DG2) and interrogate the original text (DG4), TermSight provides Summary Snippets accompanied by Highlight Bars for seamless navigation between the summary and the original text (Figure  1).
 4.3.1. Summary Snippets
-Unlike prior methods that rely on document-level overviews (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023), TermSight supports interaction and sensemaking for finer-grained snippets of information within dense sections.
-TermSight features Summary Snippets - one-sentence plain language summaries of Information Snippets (§4.1.1).
-Each Summary Snippet is rendered with color based on the Power and Relevance of the corresponding Information Snippet (§4.1.2).
-To enable users to interrogate original text and verify AI-generated summaries (DG4), clicking on a Summary Snippet jumps a user to the corresponding Information Snippet in the document (Figure  4a).
+In contrast with document-level  (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023), TermSight explores supporting interaction and sensemaking for finer-grained snippets of information within dense sections.
+TermSight features Summary Snippets: one-sentence plain language summaries of Information Snippets (§4.1.1).
+Summary Snippets are designed to be short (detailed in Section 5.2), which have been shown to be more effective for navigational tasks compared to longer snippets  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006).
+Each Summary Snippet is rendered with color based on the Power and Relevance of the corresponding Information Snippet to provide information scent (Pirolli and Card, 1999) for deciding which snippets to pay attention to (DG2).
+To enable users to interrogate original text and verify AI-generated summaries (DG4), clicking on a Summary Snippet jumps a user to the corresponding Information Snippet in the document.
 4.3.2. Highlight Bar
@@
 Highlight bars are positioned to the left of each Information Snippet in the original text, visually segmenting dense sections into Information Snippets.
-For paragraphs containing multiple Information Snippets, additional line breaks are added after each snippet to add visual separation.
+For paragraphs containing multiple Information Snippets, additional line breaks are added after each snippet for visual separation.
 Similar to the Summary Snippets, Highlight Bars are color-coded based on the Power and Relevance of the corresponding Information Snippet (§4.1.2).
-To support navigation from the original document to the summary, users can click on a Highlight Bar to jump to the corresponding Summary Snippet, as shown in Figure  4b.
+Users can click on a Highlight Bar to jump to the corresponding Summary Snippet.
 Figure 4.
-Viewport Scrolling: (a) By clicking the Summary Snippet, the viewport of the original passage will scroll to the referenced Information Snippet; (b) By clicking the Highlight bar, the left viewport will scroll to the corresponding Summary Snippet.
+Phrase Scope first identifies jargon or vague phrases in the contract (a).
+Clicking an identified phrase opens a tooltip with definition (b) and hypothetical scenario (c) while allowing users to ask clarification questions (d).
+Phrase Scope, a tooltip that provides a term definition and a scenario to contextualize the meaning of a phrase, while also allowing users to ask clarification questions.
+Users can access Phrase Scope by clicking on a system-identified phrase or highlighting any phrase they do not understand.
+The figure shows an example output for the phrase ‘royalty-free’.
+Definition: ‘Royalty-free refers to the license to use, copy, modify, or distribute your content without requiring licensing fees.’ Scenario: ‘Imagine you post a photo on ServiceX, and they use it in a global ad campaign without paying you.
+Because you agreed to a royalty-free license, ServiceX can freely use your content for commercial purposes without compensating you, reducing your control and potential earnings from your creation.’
+4.4. Phrase Scope: Phrase Identification, Definition, Scenario, and Clarification
+To help users engage with the original text (DG3), Phrase Scope first identifies phrases that may be unfamiliar or ambiguous to general audience readers and underlines them in blue (Figure 4).
+This provides a starting point for users to explore potentially unfamiliar phrases without relying on users to identify them.
+Users can also highlight any arbitrary span of text to request Phrase Scope.
+Clicking on the system or user-identified phrase opens Phrase Scope, a tooltip with three components.
+First, the definition of the phrase in the context of the ToS is offered, similar to prior work in other domains (Fok et al., 2024a).
+Beyond definitions, participants in the formative study described the challenge to contextualize abstract phrases in unforeseen scenarios.
+Inspired by prior work that encouraged scientists to reflect on the unintended consequences of their research  (Wang et al., 2024; Pang et al., 2024), Phrase Scope explores the opportunity to prompt reflection on the implications of a phrase by offering hypothetical scenarios personalized to the user persona.
+Lastly, users have an option to ask clarification questions.
+More examples of the generated definitions and scenarios can be found in Appendix Table  2.
+5. Implementation Details
+TermSight is rendered as a web interface implemented with Next.js.
+Below, we explain the input, output, processing, and overall performance of TermSight.
+Additional implementation details and prompts are specified in Appendix  B.
+5.1. System Input and Pre-processing
+TermSight assumes a set of HTML or markdown source files representing the policies included in the Terms of Service (ToS).
+We assumed clean source files because the focus of TermSight is not on developing web scraping technologies but rather on investigating meaningful navigational and reading affordances for ToS contracts.
+Each document is segmented into text ”chunks”, composed of one or more paragraphs, which are further vectorized and stored in a vector database (detailed in Appendix  B.1).
+For features that depend on user preferences, such as classifying the relevance of Information Snippets (§5.2) and generating personalized scenarios (§5.3), TermSight uses a text-based persona that includes users’ intended usage of the service (e.g., content consumers vs.
+content creators) and their values or concerns (e.g., privacy, copyright, etc).
+The user personas used for the study are further explained in §6.1.2 and displayed in Appendix  E.2.
+5.2. Obtaining Summary Snippets and Classifying Information Snippets
+For each chunk of text obtained from the document pre-processing pipeline, we prompted GPT-4o to generate a list of one-sentence summaries (Summary Snippets) each referenced to a span of the input text chunk (Information Snippet) as illustrated in Figure  5.
+While the length of the referenced text for each summary is unrestricted, each summary is constrained to a maximum of 12 words based on prompt engineering and prior work showing that short summaries can be more effective for navigation (e.g., 10-20 words)  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006).
+We noticed that setting summaries shorter than 12 words led to excessive fragmentation (i.e., too many Summary Snippets), while longer summaries tried to cover too much information, making them less skimmable.
+We then prompted GPT-4o to classify each Information Snippet along two dimensions: Power and Relevance (to user persona).
+Prompts can be found in Appendix  B.2 and  B.3.
 Figure 5.
-Phrase Scope: A tooltip that provides term definition and scenario that contextualizes the meaning of the phrase while allowing users to ask clarification questions.
-Users can hover over the references to view the referenced text in the ToS.
-4.4. Phrase Scope: Context-dependent definition, personalized scenario, and clarification question
-To help users read original text with unfamiliar or vague terminology (DG3), TermSight extends prior work focusing on offering dictionary definitions (e.g.,  (August et al., 2023)) by introducing Phrase Scope: a set of features that support the identification and interpretation of unfamiliar or vague phrases (Figure  5).
-First, TermSight identifies words and phrases that might be unfamiliar or vague to general audience readers and underlines these phrases in blue.
-This aims to provide a starting point for users to explore potentially unfamiliar phrases without relying on users to identify them.
-Clicking on the identified phrase opens Phrase Scope, a tooltip with the phrase’s definition in context, a personalized scenario to envision potential implications, and an option to ask clarification questions (examples in Table  2).
-In addition to the suggested phrases, TermSight allows users to highlight any arbitrary span of text to request Phrase Scope.
-5. Implementation details
-TermSight is rendered as a web interface implemented with Next.js. Below, we explain the input, output, processing, and overall performance of TermSight.
-5.1. System Input and Pre-processing:
-TermSight assumes a set of HTML or markdown source files representing the policies included in the Terms of Service (ToS).
-For our user study, the research team manually copied the text from documents in ToS on the internet into markdown files with annotated hyperlinks and headers.
-We assumed clean source files because the focus of TermSight is not on developing web scraping and cleaning technologies but rather on investigating meaningful navigational and reading affordances for ToS.
-Each document is segmented into text ”chunks”, composed of one or more paragraphs, which are further vectorized and stored in a vector database (detailed in Appendix  B.1).
-For the TermSight features that depend on user preferences, such as classifying the relevance of information snippets (§5.2) and generating personalized scenarios (§5.3.3), the system uses a text-based persona that includes users’ intended usage of the service (e.g., content consumers vs.
-content creators) and their values or concerns (e.g., privacy, copyright, etc.).
-The user personas used for the study are further explained in Section  6.1.2 and displayed in Appendix  F.1.
-Figure 6. A flowchart of the implementation for (1) obtaining Summary Snippets and Information Snippets and (2) classifying Power and Relevance for each Information Snippet.
-5.2. Obtaining Summary Snippets and Information Snippets:
-For each chunk of text obtained from the document pre-processing pipeline, we prompted GPT-4o to generate a list of 1-sentence summaries (Summary Snippets) each referenced to a span of the input text chunk (Information Snippet) as illustrated in Figure  6.
-While the length of the referenced text for each summary is unrestricted, each summary is constrained to a maximum of 12 words based on prior works (Cutrell and Guan, 2007; Sweeney and Crestani, 2006) and prompt engineering.
-Then, we prompted GPT-4o to classify each Information Snippet along two dimensions: Power and Relevance (to user persona).
-We detail our prompt design in Appendix  B.3.
-5.3. Phrase Scope
-Below, we detail the implementation of each feature in Phrase Scope.
-5.3.1. Identifying unfamiliar or vague phrases
-Prior works have demonstrated the capabilities of LLM in personalized jargon identification (Guo et al., 2024) and detecting phrases that can potentially be elaborated (Fok et al., 2024a).
-Inspired by this work, we employed GPT-4o with two few-shot prompts to identify potentially unfamiliar and vague phrases within the document chunks produced by the pre-processing pipeline (Prompts in Appendix  B.4.1).
-The input chunks are also referred to and used as the ”phrase context” for each identified phrase.
-5.3.2. Generating phrase definitions or responses to user questions
+A flowchart of the implementation for (1) obtaining a list of Summary Snippets each referenced to a span of the input text (i.e., Information Snippet) and (2) classifying Power and Relevance for each Information Snippet.
+The flowchart of TermSight’s implementation process.
+(1) The figure shows that a chunk of text is first summarized into a list of Summary Snippets, each referenced to a span of the input text chunk (Information Snippet).
+(2) After this step, TermSight classifies every Information Snippet with respect to Power and Relevance.
+5.3. Generating Phrase Scope
+Similar to prior work (Guo et al., 2024; Fok et al., 2024a), we prompt an LM to identify potentially unfamiliar and ambiguous phrases within document chunks produced by the pre-processing pipeline.
 To generate in-context definitions for the identified phrases, we used a retrieval-augmented question answering approach  (Fok et al., 2024a).
-We first query the vectorized document chunks (§5.1) to retrieve chunks that might define the phrase.
-The database query is framed as a question: “What does {input phrase} refer to in the sentence: {phrase context}.
-Here, phrase context refers to the chunk of text containing the input phrase as defined in §5.3.1.
-Then, the query is embedded using OpenAI’s text-embedding-3-small model, and the top 15 chunks from the vector database are retrieved based on cosine similarity.
+We vectorize the document chunks (§5.1) and query these chunks to retrieve relevant document context.
 These chunks are collectively referred to as ”retrieved chunks”.
-Using the phrase context and the retrieved chunks, we prompted GPT-4o to generate an in-context definition with references to the retrieved chunks.
+Then, we prompted GPT-4o to generate an in-context definition with references to the retrieved chunks.
 When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked.
-We detail our prompts in Appendix  B.4.2.
-5.3.3. Generating scenarios for phrases
-Scenarios intend to complement phrase definitions, helping users interpret and envision the potential implications of a phrase in context.
-In addition, customizing scenarios based on user personas intends to ensure that the generated examples are relevant to the user’s intended usage of the service (e.g., content creator or consumer) and personal value (e.g., privacy, copyright).
-Building on prior works that demonstrated the effectiveness of large language models (LLMs) in generating unintended consequences (Wang et al., 2024) and stories to teach legal concepts (Jiang et al., 2024), we employed GPT-4o with zero-shot prompting (Figure  27) to generate the scenarios with the phrase context, phrase definition, and user persona.
+Scenarios aim to prompt reflection on the potential implications of a phrase based on users’ intended usage of the service and personal value.
+We employed GPT-4o with zero-shot prompting to generate the scenarios tailored to the user persona.
+The prompts used for Phrase Scope are detailed in Appendix  B.4.
 5.4. Evaluation of System Output
-Before reporting on our user study, we randomly sampled and conducted a manual evaluation of TermSight’s core outputs.
+Before reporting on our user study, we randomly sampled and conducted a manual evaluation of TermSight’s core outputs (e.g., summary snippets, definitions, scenarios).
 Our goal was not to assess advances in system performance, but to verify that the system can produce meaningful outputs that could support ToS reading.
@@
 Similarly, we reviewed 113 generated definitions and scenarios each.
-One scenario was found to be factually incorrect due to hallucination.
+One scenario was found to be factually incorrect due to a hallucination.
+The input context and phrase stated that users do not gain ownership rights by downloading content from the service.
+However, the scenario claimed that users might lose ownership rights over the content they create by uploading it to the service.
 All the definitions were correct except for 4 overly general definitions for service-specific phrases not explicitly defined anywhere in the ToS.
-More details are provided in Appendix C.
+More details of the evaluation are provided in Appendix C.
 6. User Study
-We conducted a counterbalanced within-subject experiment with 20 participants to evaluate TermSight. We asked the following research questions:
+Similar to prior HCI studies  (Chen et al., 2023; Cao et al., 2023; Kambhamettu et al., 2024), the goal of our user study is not to prove TermSight as the definitive solution through a series of ablation experiments, but to use TermSight to explore the problem and solution space of designing contract reading interfaces.
+We asked the following research questions:
 RQ1: How did participants perceive TermSight and its features?
-RQ2: How did participants read with TermSight and its features?
-RQ3: How did TermSight influence comprehension and recall?
+RQ2: How did TermSight influence ToS comprehension and recall?
+RQ3: How did participants read with TermSight and its features?
 6.1. Study Design
-6.1.1. Study Conditions/Interface Variants
-To evaluate TermSight, we conducted a within-subject experiment where each participant used both TermSight and a baseline interface to read 2 services’ ToS contracts, one with each interface variant.
-The order of the interface variant and service type were counterbalanced to reduce ordering effects.
-The baseline interface intended to mimic a standard ToS.
-The baseline interface had the same layout as TermSight without the Power Meter, Summary Snippet, Highlight Bar, and Phrase Scope.
-In place of the Summary Snippets on the left was a table of contents.
-Participants could click a section header in the table of contents to navigate to the corresponding section.
-6.1.2. Materials: ToS and User Persona
+We conducted a within-subject study where each participant used both TermSight and a baseline interface to read two services’ ToS, once with each interface variant.
+The order of the interface and service type was counterbalanced to reduce ordering effects.
+6.1.1. Baseline
+Because there is no well-established standard ToS reader, we use the most commonly used interface as the baseline (i.e., an HTML reader).
+The baseline interface had the same layout as TermSight without the Power Meter, Summary Snippet, Highlight Bar, and Phrase Scope (Appendix Figure  25).
+In place of the Summary Snippets was a table of contents that can be used to navigate to different sections.
+Participants’ feedback in the user study confirmed that the baseline offered a strong frame of reference (§7.1).
+While reading interfaces exist for other domains (e.g.,  (August et al., 2023; Chen et al., 2023; Fok et al., 2023)), adopting them to contracts would result in novel systems unsuitable as baselines.
+For example, prior work has designed key questions (August et al., 2023) or faceted highlighting (Fok et al., 2023) specific to academic papers, which require co-design studies with lawyers to adapt them to contracts.
+Injecting additional features (e.g., overview summaries  (Tabassum et al., 2018), chat (Zhao and Lee, 2020)) into one or both conditions may also introduce confounds unrelated to our research questions.
+For example, by adding a policy or section-level summary in the baseline, the comparison with TermSight will be confounded by the design of summaries in the baseline (e.g., levels of complexity and detail), a variable that may impact different readers differently (August et al., 2024).
+6.1.2. Materials
+Terms of Service
 For the user study, we used the ToS of one social media site (Reddit) and one e-commerce site (Poshmark) because social media and e-commerce platforms are among the most visited digital services and are representative of the standard long-form contract used in prior studies  (Kay and Terry, 2010; Taber et al., 2020; Good et al., 2005; Tabassum et al., 2018; Obar and Oeldorf-Hirsch, 2020).
-For each service, we manually collected policies that are linked or hyperlinked as part of the ToS.
-We did not include policies that are location-specific (e.g., a California Privacy Notice).
-For Reddit, we also did not include policies for advertisers, brand&press, or governmental agencies as they were not relevant to a lay user and to keep the number of policies for both services similar.
+For each service, we collected policies that are linked as part of the ToS.
+We did not include location-specific policies (e.g., a California Privacy Notice).
+For Reddit, we also did not include policies for advertisers, publishers, or governmental agencies to keep the number of policies for both services similar.
 We collected a total of 14 policies for Reddit and 15 for Poshmark.
 The service names for Reddit and Poshmark were anonymized as ServiceX and ServiceY.
+User Persona
 Features of TermSight rely on a user persona to determine relevance.
-For the user study, we designed two personas based on users’ information needs identified in the formative study: a content consumer who posts personal content on social media sites (Figure  17) and a buyer who rarely posts reviews on E-commerce sites (Figure  18).
-Figure 7.
-The experiment procedure for the within-subject study.
-At the top, we show the procedure for the entire study.
-At the bottom, we show the process diagram for a reading session with either TermSight or Baseline.
-The orders for the reading session were counterbalanced in the study.
+For the user study, we designed two personas based on users’ information needs identified in the formative study: a content consumer who posts personal content on social media sites and a buyer who rarely posts reviews on e-commerce sites (Appendix  E.2).
+Participants in the user study validated and found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  13).
 6.1.3. Study Procedure
-The entire study took 90 minutes and participants were compensated 25\$.
-The study was approved by the Institutional Review Board (IRB) office at our organization.
-The study was composed of 2 reading sessions, 45 minutes each (Figure  7).
+The entire study took 90 minutes, and participants were compensated 25\$.
+The study was approved by the Institutional Review Board (IRB) at our organization.
+The study was composed of two reading sessions, 45 minutes each.
 After obtaining consent and before the reading sessions, we asked participants about their past experiences with ToS.
@@
 Then, an anonymized description of the service and user persona was given.
-Participants were given 10 minutes to read the ToS without having to speak aloud.
+Similar to prior studies (August et al., 2023), participants were given 10 minutes to read the ToS without having to speak aloud.
 Reminders were given at the 5 and 1-minute mark.
-After the reading session, participants completed seven 5-point Likert-style rating questions about their reading experiences followed by one free-response recall question and six multiple-choice comprehension questions.
+After the reading session, participants completed seven 5-point Likert-style rating questions about their reading experiences, followed by one free-response recall question and six multiple-choice comprehension questions.
 We also conducted semi-structured interviews with participants about their experiences and reading strategies for 10 minutes.
 The same procedure was repeated for the second reading session.
-The study materials are included in Appendix  F and the supplemental materials.
-Figure 8.
-Participants’ self-ratings of their general familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service.
-Most participants were familiar with both social media and e-commerce platforms in general and found the given persona to align with their personal usage of the service and personal value.
+The study materials are included in Appendix  E.
 6.2. Participants and Setting
@@
 Two participants encountered technical difficulties, and their data were removed from the analysis.
-Of the remaining 20 participants, 13 self-identify as female and 7 as male.
-1 participant had a high school degree, 3 had an associate degree, 12 had a bachelor’s degree, and 4 had a master’s degree.
-4 participants were between the ages of 18-25, 5 were between 26-35, 5 were between 36-45, and 6 were between 46-55.
-When asked about how many ToS they’ve read, 4 participants read none, 5 read between 1—3, 5 read between 4—6, 1 read between 7—9, and 5 read greater than 10.
-The median of participants’ self-rated familiarity with ToS was 3.5 ($\sigma=1.5$).
-Most participants were generally familiar with social media and e-commerce platforms and found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  8).
+The median of participants’ self-rated familiarity with ToS was 3.5 ($\sigma=1.5$) out of 5.
+Most participants were generally familiar with social media and e-commerce platforms and found their given user persona to highly align with their own profile (Figure  13).
 None of the participants had read or remembered the ToS for Reddit and Poshmark prior to the study.
+Additional participant demographics can be found in Appendix  E.3.
 6.3. Measures
-6.3.1. Perceived reading experiences
-We collected 5-point Likert-style ratings (1=“Not at all,” 5=“Very”) of participants’ reading experiences (Appendix  F.3).
-The experience measures included ease of reading, perceived understanding, and confidence in obtaining relevant information  (August et al., 2023).
-In addition, we included two 5-point Likert-style rating questions related to the ease of deciding which policy to read and what text to read within a policy, both of which are challenges identified in our formative study.
-Lastly, we included two 5-point Likert-style questions about their willingness to (1) read other ToS with the interface and (2) spend more time on their given ToS with the interface and get a link to the ToS after the study  (August et al., 2024).
-6.3.2. Reading comprehension and recall
-For reading comprehension, we wrote 10 multiple-choice questions for each service and selected 6 questions that were relevant for the given user persona for each service, similar to prior works on evaluating reading comprehension for ToS and privacy policy (Tabassum et al., 2018; Vu et al., 2007; August et al., 2023).
-The questions are different for the two services to minimize learning effects.
-Comprehension was measured as the number of questions participants got right out of 6.
-Examples of comprehension questions and their answers are listed in Table  1.
+6.3.1. Reading Experiences
+We collected seven 5-point Likert-style ratings (1=“Not at all,” 5=“Very”) of participants’ reading experiences (Appendix  E.5).
+The experience measures included ease of reading (adopted from perceived effort measure in NASA TLX)  (August et al., 2023; Hart and Staveland, 1988), perceived understanding  (August et al., 2023), and confidence in obtaining relevant information  (August et al., 2023).
+We added two questions specific to challenges identified in our formative study: the ease of deciding which policy to read and what text to read within a policy.
+Finally, we included two questions on participants’ willingness to read ToS with the interface and their willingness to spend more time on the ToS with the interface (August et al., 2024).
+6.3.2. Comprehension
+We wrote 10 multiple-choice questions for each service and selected 6 questions that are relevant for the given user persona (§6.1.2) for each service, similar to prior work on evaluating reading comprehension for ToS and privacy policies (Tabassum et al., 2018; Vu et al., 2007; August et al., 2023).
+We designed the comprehension questions to have one answer based on the original text, independent of features of TermSight (e.g., Summary Snippets).
+Consequently, participants need to read the original text to fully answer the questions.
+The questions were different for the two services to minimize learning effects.
+Comprehension was measured as the number of questions participants got right.
+We note that past studies with reading interfaces often fail to show differences in how people answer comprehension questions (Tabassum et al., 2018; Korunovska et al., 2020; Robinson and Zhu, 2020; August et al., 2023; Head et al., 2021).
+Similar to prior work  (August et al., 2023), our intention in including comprehension questions was to ensure that TermSight did not detract from overall comprehension.
+The questions can be found in the supplemental materials.
+6.3.3. Recall
 For recall, participants were asked to write down information in ToS they found interesting, surprising, or any detail they remembered: “What did you learn from reading the ToS?
@@
 Then, we labeled whether each reference was correct based on the ToS.
-Recall is measured as the number of correct facts participants wrote in the response  (Taber et al., 2020; Good et al., 2005).
+Recall is measured as the number of correct facts in the response  (Taber et al., 2020; Good et al., 2005).
 6.3.4. Feature Usage
@@
 In this section, we introduce our analysis framework for the user study.
-We used a causal framework (§6.4.1) to understand the effects of the treatment and a Bayesian analysis (§6.4.2) to estimate the treatment effect.
+We used a causal framework to understand the effects of the treatment and a Bayesian analysis to estimate the treatment effect (§6.4.1).
 Then, we discuss the thematic analysis (§6.4.2) used to analyze the qualitative responses from the semi-structured interviews.
-6.4.1. Structural Causal Framework
-We used a structural causal framework popularized by Pearl (2009) to better understand the effects of the treatment.
-A structural causal model framework comprises exogenous variables $U$, endogenous variables $V$ and a set of functions $F$ that represent the physical process that results in the values of the endogenous variables.
-For example, the question $v_{i}=f(v,u)$ implies that the values of $v$ and $u$ completely determine the value of $v_{i}$ through the function $f$.
-The structural causal model can be represented as a Directed Acyclic Graph (DAG) where the nodes represent the variables and the edges represent the causal relationships among the variables.
-Thus, in the preceding example, $v$ and $u$ are the parents of $v_{i}$ in the DAG.
-The structural causal model framework is non-parametric, but frequently used with generalized linear models to estimate the functions $f$.
-The structural causal framework has connections with structural equation models (Wright, 1934) used in the Economic Sciences, though the latter were developed with linear relationships in mind.
-Figure 9 shows the DAG for the study.
-[…]
 Figure 6.
-The causal graph (a DAG) has three types of nodes: Treatment ($T$) colored blue, pre-treatment covariates $D$ (demographics), $K$ (knowledge of Terms of Service) colored gray, and measured outcomes $Q$ (comprehension), $R$ (recall of correct facts), and $E$ (experience of the interface) colored orange.
-Since we use randomized assignment, including the interface type (treatment vs.
-control), type of ToS (social media vs.
-e-commerce website), and order of presentation (first vs.
+The causal graph (DAG) has three types of nodes: treatment ($T$); pre-treatment covariates $D$ (demographics), $K$ (knowledge of Terms of Service); and measured outcomes $Q$ (comprehension), $R$ (recall), and $E$ (user experience).
+Since we use randomized assignment for the interface type (treatment vs.
+control), ToS type (social media vs.
+e-commerce), and presentation order (first vs.
 second), the treatment $T$ is independent of the pre-treatment covariates $D$ and $K$.
-Thus, to estimate the causal effect of the treatment $T$ on the outcomes $Q$, $E$, and $R$, we can use a simple regression model, without conditioning on the pre-treatment covariates $D$ and $K$ since there is no backdoor path from the treatment $T$ to the outcome variables $Q$, $E$, and $R$.
+To estimate the causal effect of the treatment $T$ on the outcomes $Q$, $E$, and $R$, we can use a simple regression model, without conditioning on covariates $D$ and $K$, since there is no backdoor path from the treatment $T$ to the outcome variables.
+The causal graph (DAG) has three types of nodes: treatment (T) colored blue; pre-treatment covariates D (demographics), K (knowledge of Terms of Service) colored gray; and measured outcomes Q (comprehension), R (recall), and E (user experience) colored orange.
+Since we use randomized assignment for the interface type (treatment vs.
+control), ToS type (social media vs.
+e-commerce), and presentation order (first vs.
+second), the treatment T is independent of the pre-treatment covariates D and K.
+To estimate the causal effect of the treatment T on the outcomes Q, E, and R, we can use a simple regression model, without conditioning on covariates D and K, since there is no backdoor path from the treatment T to the outcome variables.
 6.4.1. Bayesian Analysis
-We used Bayesian inference to estimate the treatment effect of TermSight on the outcomes of interest.
+First, we used a structural causal framework popularized by Pearl (2009) to understand the effects of the treatment (Figure  6 shows the DAG for the study).
+More details about structural causal framework can be found in Appendix  F.1.
+Next, we used Bayesian inference to estimate the treatment effect of TermSight on the outcomes of interest.
 While the use of Bayesian estimates is growing in HCI (Kay et al., 2016; Koshy et al., 2023), we briefly justify its use over traditional non-Bayesian methods.
-Standard statistical methods including t-tests of significance can be powerful, but they require us to test for the conditions under which a test is valid (e.g., Normality for the t-test).
-Prior work in Biological and Psychological Sciences has shown that Null Hypothesis Statistical Testing (NHST) including $p$-values are problematic, and must be used with caution.
-They depend on the researchers’ intentions, when they stop collecting samples, and pre-selection of hypothesis (Ioannidis, 2005, 2012; Button et al., 2013).
-We believe that Bayesian techniques offer several advantages over non-Bayesian methods.
 First, as Kay et al.
-(2016) point out, a Bayesian framework leads to an accumulation of knowledge within HCI, where the posterior of the parameters in a prior work can serve as the prior in the current experiment.
-Second, a Bayesian model is transparent—the researcher will foreground all the assumptions in the analysis via their model, allowing for a principled critique of the model.
+(2016) points out, a Bayesian framework leads to an accumulation of knowledge within HCI, where the posterior of the parameters in a prior work can serve as the prior in the current experiment.
+Second, a Bayesian model is transparent—the researcher will foreground all the assumptions in the analysis via their model.
 Third, use of a Bayesian framework shifts the discussion from “did it work” to “effect size of the intervention” (Kay et al., 2016).
-While NHST techniques can compute the confidence interval, these intervals are susceptible to misinterpretation (Belia et al., 2005; Hoekstra et al., 2014) and importantly, underemphasized in favor of $p$-values.
-Finally, as  McElreath (2015) points out, a Bayesian model with use of maximum entropy priors for the parameters (e.g., the normal distribution) is the most conservative given the evidence to estimate of the effect of the treatment.
+While NHST techniques can compute the confidence interval, these intervals are susceptible to misinterpretation (Belia et al., 2005; Hoekstra et al., 2014) and, importantly, underemphasized in favor of $p$-values.
+Finally, as  McElreath (2015) points out, a Bayesian model with the use of maximum entropy priors for the parameters (e.g., the normal distribution) is the most conservative given the evidence to estimate the effect of the treatment.
 6.4.2. Thematic Analysis
@@
 7.1. RQ1: How did participants perceive TermSight and its features?
+17 participants highlighted that the baseline interface was easier to read than most of the ToS they had seen in the past, which often lacked the top navigation panel or a table of contents.
 When asked about which version of the interface they prefer, 19 participants preferred TermSight.
 One participant preferred the baseline interface, citing difficulty navigating TermSight’s two scrollable columns on a small screen with a trackpad that lacks scroll functionality.
-Below we present our quantitative findings on the experience outcomes (§7.1.1), followed by qualitative findings organized around recurring themes about features of TermSight.
-Figure 10.
-5-Point Likert Ratings of participants’ reading experiences: Participants found it to be easier to find and understand relevant information.
-Participants were also more willing to spend more time on the ToS with TermSight and read other ToS with TermSight.
-The 7 experience questions are specified in Appendix F.3.
-7.1.1. Quantitative Findings: TermSight improved user experience measures
-Figure 10 shows participants’ ratings of the 7 reading experience measures for both interfaces.
-Our Bayesian model analysis (model details in Appendix D.1) reveals a significant effect of the TermSight interface on the reading experience measures.
-The forest plots shown in Figure 19 in the Appendix demonstrate that the 94% Highest Posterior Density Interval (HPDI [^1] In Bayesian analysis, the 94% HPDI refers to that smallest interval (this interval is unique) of the posterior distribution that has 94% of the probability mass.
+Below we present our quantitative findings on the experience outcomes (§7.1.1), followed by qualitative findings organized around recurring themes about features of TermSight (§7.1.2 – §7.1.5).
+7.1.1. TermSight improved all seven user experience measures
+Figure  8 shows participants’ ratings of the 7 reading experience measures for both interfaces.
+Our Bayesian analysis (model details in Appendix F.2) reveals a significant effect of the TermSight interface on the reading experience measures.
+Participants found reading ToS with TermSight to require significantly less effort.
+When navigating ToS, deciding which policy to read and what text to read within a policy was significantly easier with TermSight.
+Moreover, participants were significantly more willing to read ToS and spend more time reading ToS with TermSight.
+The forest plots shown in Appendix Figure  14 demonstrate that the 94% Highest Posterior Density Interval (HPDI [^3] In Bayesian analysis, the 94% HPDI refers to that smallest interval of the posterior distribution that has 94% of the probability mass.
 It is common in Bayesian analysis to use intervals distinct (97% and 89% HPDI intervals are also used) from the typical 95% value to avoid the confusion of the frequentist confidence intervals.) for each treatment–control pair for every user experience measure does not overlap with each other.
 This indicates a significant positive effect of the treatment on all 7 user experience measures.
-In Figure 11, we show the contrast between the treatment and control conditions averaged across 7 user experience measures as well as the posterior of the effect size.
-The posterior distribution (Figure 11(b)) of the effect size (Cohen’s $d$) of the treatment on the user experience is centered around 2.7, with an HPDI of [2.3, 3.2], indicating a large effect size [^2] This is the effect size measured not on the outcome Likert scale (1–5) but on latent scale of the model coefficients..
-Because the HPDI doesn’t overlap with the ROPE, the effect of TermSight on user experience is significant.
-Below, we further present recurring themes related to participants’ perceptions and feedback on features of TermSight.
-(a) Posterior distributions of the coefficients for the treatment and control conditions.
-(b) Posterior distribution of the effect size of the treatment on the user experience.
 Figure 7.
-Figure 11(a) shows the posterior distributions of the coefficients for the treatment and control conditions.
-Note that the posterior distribution of the treatment condition is shifted to the right, indicating a positive effect of the treatment on the user experience.
-Figure 11(b) shows the posterior distribution of the effect size of the treatment on the user experience centered around 2.7, suggesting a large effect size.
-Additionally, the HPDI doesn’t overlap with the ROPE, suggesting a significant effect of the treatment on user experience.
-7.1.2. Power Meter provided information scent of the policies and reduced navigation cost
-All 20 participants described that the Power Meter helped them gain an intuition of the content of a policy and provided guidance.
-Participants described how the saturated colors, especially red and green, “piqued interest”(P4).
-For example, P12 described how Power Meter provided guidance and supported decision-making on which policy to pay attention to.
+The posterior distribution of the contrast between treatment and control on user experience is shifted to the right, with the posterior distribution of the effect size centered around 2.7 without overlapping with ROPE, suggesting a significant and large effect size.
+Figure 7 shows the posterior distribution of the contrast between treatment and control on user experience.
+Note that the posterior distribution of the contrast is shifted to the right, with the posterior distribution of the effect size centered around 2.7 without overlapping with ROPE, suggesting a significant and large effect size.
+In Figure  7, we show the contrast between the treatment and control conditions averaged across 7 user experience measures as well as the posterior of the effect size.
+The posterior distribution of the effect size (Cohen’s $d$) of the treatment on user experience is centered around 2.7, indicating a large effect size [^4] This is the effect size measured not on the outcome Likert scale (1–5) but on the latent scale of the model coefficients..
+Because the HPDI doesn’t overlap with the ROPE [^5] We use a Region of Practical Equivalence (ROPE) of [-0.1, 0.1].
+ROPE is a region where the difference between the treatment and control conditions is considered practically equivalent.
+This prevents us from focusing on small differences that are not practically meaningful., the effect of TermSight on user experience is significant.
+To further disentangle how individual features of TermSight contributed to the improvement in user experience, we present recurring themes related to participants’ perceptions and feedback on features of TermSight.
+Figure 8.
+Participants’ ratings of reading experiences.
+Participants found it to be easier to find and understand relevant information with TermSight.
+Moreover, participants were more willing to read and spend more time on ToS with TermSight.
+The figure shows seven barplots of 5-point Likert Ratings of participants’ reading experiences with TermSight and the baseline interface.
+Participants rated TermSight to be better for all seven experience measures.
+For example, participants found it to be easier to find and understand relevant information.
+Participants were also more willing to spend more time on the ToS with TermSight and read other ToS with TermSight.
+7.1.2. Power Meter reduced navigation cost
+All 20 participants described that the Power Meter helped them gain an intuition of the content within each policy and supported decision-making on which policy to pay attention to.
 “The color system was really guiding a lot of my decisions on which policies to read.
-The vibrant red was always what I was gonna try to get to 1st and the things that I cared the most about.
+The vibrant red was always what I try to get to first and the things that I cared the most about.
 So, finding which terms I’m agreeing to that I have the least amount of agency.” (P12)
@@
 It was more like a guessing game.” (P6)
-8 participants described how the document preview that appears when hovering over the Power Meter complemented and provided a more detailed overview of the policy to help them decide what to read.
-Additionally, 9 participants mentioned how the document preview allowed them to access information from linked pages or policies without “clicking on the link and possibly losing where I am on” (P1).
-7.1.3. Summary Snippet and Color provided guidance and helped surface power and relevance within documents
-18 Participants described how the color of the Summary Snippets helped surface where power lies in the contract, which is challenging to ”see” when presented in plain text with the baseline interface.
-“The color coding gave me a better understanding of who’s benefiting more, how is it benefiting them, and how it affects me.
-Versus [in the baseline interface] it feels like it only really benefits the company itself.” (P18)
+Additionally, eight participants described how the document preview reduced navigation cost by allowing them to access information from other policies without “clicking on the link and possibly losing where I am on” (P1).
+7.1.3. Summary Snippets and Color provided guidance and helped surface power and relevance within documents
+Eighteen participants described how the color of the Summary Snippets helped surface where power lies within a document, which is challenging to “see” otherwise.
+“The color coding gave me a better understanding of who’s benefiting more, how it’s benefiting them, and how it affects me.
+Versus [in the baseline], it feels like it only really benefits the company itself.” (P18)
 Moreover, all participants described how the color and summary allowed them to make decisions and prioritize what to focus on strategically.
@@
 It would take me a lot longer to filter for information otherwise.” (P6)
-While participants described focusing on the more saturated colors, especially red and green, 6 participants also described skimming the less saturated ones to validate and make sure they didn’t miss information.
-7.1.4. Summary Snippet simplified and broke down dense text to support sensemaking
-All participants highlighted how the Summary Snippets “broke down the entire section into a few bullet-pointed summaries” (P11), which helped them absorb the information and surface relevant pieces of information hidden in dense text.
-9 participants further described how the Summary Snippets simplified the text and served as a scaffold to encourage reading the original text.
+While participants described focusing on the more saturated colors, especially red and green, six participants complimented how TermSight made AI decisions (i.e., classification of power and relevance) transparent by keeping all the Summary Snippets visible and described skimming the less saturated ones to validate and make sure they didn’t miss information.
+7.1.4. Summary Snippets simplified and broke down dense text to support sensemaking
+Summary Snippets “broke down the entire section into a few bullet-pointed summaries” (P11), which helped participants surface and absorb relevant information hidden in dense text (20/20).
+Nine participants further described how the Summary Snippets simplified the text and served as a scaffold to encourage reading the original text.
 “Not only does it [TermSight] bring me up to things that I would normally miss.
@@
 … It is really guiding me through this entire document, and it’s simplifying it.” (P2)
-Similarly, 4 participants mentioned how the highlight bars and line breaks added for each information snippet made it easier to consume the original text without having to “mentally separate the concepts” (P7).
 P20 described the experience of using Summary Snippets to make sense of the big ideas of the section while being able to dive deeper as forming “a web of thought and ideas: what everything kind of means together and piece by piece”.
-Not only do the Summary Snippets serve as an entry point to the original text, 4 participants described how the Summary Snippets helped them check their understanding after reading the original text.
-7.1.5. Term definition and scenario help consume original text and envision implications
-14 participants described that the underlined phrases attracted their attention, made them feel like these were phrases that “should be known and understood” (P18), and prompted them to check on their understanding of the phrase.
-In addition, participants commented on how the definitions and scenarios made legal language more approachable (14/20).
+Not only do the Summary Snippets serve as an entry point to the original text, four participants explicitly highlighted how the Summary Snippets helped them check their understanding after reading the original text.
+7.1.5. Term definitions and scenarios helped consume original text and envision implications
+Fourteen participants noted that the system identified phrases attracted their attention and prompted them to check on their understanding of the phrase that they may not have noticed otherwise.
+Moreover, the definitions and scenarios made legal language more approachable for participants (14/20).
 “It really just made it a lot easier to understand, because most of the time they write it in legal terms that most people don’t know.
@@
 So this broke it down easier for most people like the Layperson.” (P17)
-12 participants highlighted how the scenarios helped interpret the meaning and envision potential implications of the phrase in the context of the user’s intended usage of the service and was “easier to resonate”(P17).
+12 participants highlighted how the scenarios helped envision potential implications of the phrase in the context of the user’s intended usage of the service and was “easier to resonate” (P17).
 “I like that it provides a real-world example of when and how [the phrase] would be relevant to someone.
@@
 Moreover, one participant (P1) noted how the scenarios helped her come up with more questions, which she asked with the Ask function.
-Through this process, TermSight helped P1 to quickly obtain information from correlated policies.
-7.2. RQ2: How did participants read with TermSight and its features?
+Through this process, TermSight helped P1 to quickly obtain information from related policies.
+7.2. RQ2: How did TermSight influence comprehension and recall?
+(a) Posterior distribution of contrast (between treatment and control) on comprehension.
+The figure shows the posterior distributions of the contrast between the treatment and control conditions on comprehension outcome.
+The posterior distribution is nearly evenly split around zero and overlaps with ROPE, indicating no effect of the treatment on the comprehension quiz questions.
+(b) Posterior distribution of contrast (between treatment and control) and effect size on recall.
+On the left, the figure shows the posterior distributions of the contrast between the treatment and control conditions for the recall outcomes.
+The posterior distribution of the treatment condition is shifted to the right, indicating a positive influence of the treatment on Recall.
+On the right, the figure shows the posterior distributions of the effect size of treatment centered around 0.6, suggesting a medium effect size.
+Yet, the HPDI has an overlap with the ROPE of [-0.1, +0.1], implying no significant effect of the treatment on the user recall.
 Figure 9.
-Usage of features during the reading session.
+In (a), the distribution centers around zero and overlaps with ROPE, indicating no effect of treatment on comprehension.
+(b) illustrates a positive influence of treatment on recall with the posterior distribution of the effect size centered around 0.6, suggesting a medium effect size.
+Because the HPDI overlaps with ROPE, there is no significant effect of treatment on recall.
+In (a), the distribution centers around zero and overlaps with ROPE, indicating no effect of treatment on comprehension.
+(b) illustrates a positive influence of treatment on recall with the posterior distribution of the effect size centered around 0.6, suggesting a medium effect size.
+Because the HPDI overlaps with ROPE, there is no significant effect of treatment on recall.
+7.2.1. Comprehension
+Out of 6 questions, participants on average scored 2.00 ($\sigma=1.03$) when using TermSight and 2.15 ($\sigma=1.23$) when using the baseline interface.
+Our Bayesian analysis for the comprehension quiz (model details in Appendix  F.3) reveals no significant differences in the comprehension scores across interface conditions.
+As shown in Figure  9(a), the posterior distribution of the contrast between treatment and control is centered around zero.
+When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair as shown by the forest plots in Appendix Figure 15, suggesting no significant differences.
+While we aimed to design comprehension questions with one single answer based on the original text, we acknowledge that legal contracts and problems may contain room for interpretation  (Marotta-Wurgler and Stein, 2025).
+Therefore, we performed a supplemental, more conservative analysis by removing any question where the applicability of the original text to the question may contain room for interpretation, potentially leading to different answer choices (removed 2 out of 6 for each service type).
+The findings remained consistent, showing no significant difference in comprehension.
+More details of the analysis can be found in Appendix G.
+7.2.2. Recall
+Participants on average recalled 1.54 ($\sigma=1.96$) correct facts using TermSight and 1.30 ($\sigma=1.72$) correct facts using the baseline interface.
+Our Bayesian analysis for the recall task (model details in Appendix F.4) reveals no significant differences in the recall scores across interface conditions.
+Figure  9(b) shows the posterior distribution of the effect size of the treatment on recall ($\mu=0.6$, HPDI = [-0.35, 1.6]).
+Despite the mean being 0.6, suggesting a medium effect size, the HPDI interval has an overlap with the ROPE, implying no significant effect of the treatment on recall.
+7.3. RQ3: How did participants read with TermSight and its features?
+Figure 10.
+Feature usage.
 Each dot represents the number of times a feature is used by one participant during a reading session.
 Participants used most features of TermSight.
-In contrast to the baseline, participants when using TermSight more frequently clicked the Summary Snippets as opposed to the section titles to navigate to the original text.
+In contrast to the baseline, participants, when using TermSight, more frequently clicked the Summary Snippets as opposed to the section headers to navigate to the original text.
+Box plots for the usage of features during the reading session for TermSight and the Baseline interface.
+Participants used most features of TermSight.
+In contrast to the baseline, participants, when using TermSight, more frequently clicked the Summary Snippets as opposed to the section titles to navigate to the original text.
+7.3.1. Feature Usage
+In the baseline interface, participants on average navigated to 7 policies ($\sigma=5$).
+Participants described trying to skim everything from top to bottom and often skipped difficult sections.
+11 participants clicked the section heading in the table of contents at least once to navigate to targeted sections.
+As shown in Figure  10 and  11, participants used most features of TermSight throughout the session.
+When using TermSight, participants on average navigated to 6 policies ($\sigma=3$).
+19 participants hovered over a Power Meter to gain an overview of documents in ToS for more than 1 second.
+13% (17/135) of the total usage was for previewing documents hyperlinked inline, demonstrating that the Power Meter provided value also for inline hyperlinks.
+Summary Snippets were the most frequently used feature.
+19 participants clicked the Summary Snippets to navigate and read the original text at least once ($\mu=17,\sigma=13$).
+In contrast, participants rarely clicked the section headers to navigate to the original text (4 clicks in total across 3 participants).
+This demonstrates that Summary Snippets were potentially more useful as navigational features to surface and interrogate relevant information from the original text compared to section headers.
+In addition, 10 participants clicked the Highlight Bar to read the Summary Snippet after reading the original text to check their understanding.
+For Phrase Scope, 16 participants accessed phrase definitions and 12 participants viewed hypothetical scenarios.
+Out of 77 instances when Phrase Scope was accessed, 72 were suggested by TermSight.
+This demonstrates the value of suggesting difficult or ambiguous phrases as opposed to solely relying on users to notice them.
 Figure 11.
@@
 The Summary Snippet is the most frequently used feature throughout.
-7.2.1. Feature Usage
-In the baseline interface, participants on average navigated to 7 policies ($\sigma=5$).
-Participants’ reading behaviors were similar to our findings from the formative study (§3.3.2).
-Participants described how they were trying to skim everything from top to bottom.
-When they encounter sections that were hard to read, they tended to skip.
-Only 11 participants clicked the section heading in the table of contents at least once to navigate to targeted sections.
-As shown in Figure  12 and  13, participants used most features of TermSight throughout the session, demonstrating that TermSight provided value throughout the session.
-When using TermSight, participants on average navigated to 6 policies ($\sigma=3$).
-19 participants used the document preview of Power Meter to gain an overview of documents in ToS for more than 1 second.
-13% (17/135) of the total usage was for previewing documents hyperlinked inline, demonstrating that the Power Meter provided value also for inline hyperlinks.
-Summary Snippet was the most frequently used feature, 19 participants clicked the Summary Snippet to navigate and read the original text at least once.
-On average, they clicked 17 Summary Snippets ($\sigma=13$).
-In contrast, participants rarely clicked the section headings to navigate to the original text (4 clicks in total across 3 participants).
-This demonstrates that Summary Snippets were potentially more useful as navigational affordances to surface and interrogate relevant information from the original text compared to section headers.
-We also noticed how 10 participants clicked the Highlight Bar to read the summary snippet after reading the original text, suggesting how these participants were back-checking their understanding using the summary snippets.
-For Phrase Scope, 16 participants viewed the term definitions and 12 participants viewed the term scenarios.
-Participants on average used term definition 3 times ($\sigma=4.3$) and term scenario 1 time ($\sigma=2.7$).
-Out of the 77 instances when Phrase Scope was accessed, 72 were suggested by TermSight.
-Participants rarely manually highlighted phrases to activate phrase scope, asked questions, or checked the references provided for the definitions.
-While only 1 participant asked questions using Phrase Scope, P1 asked 8 questions in a 10-minute reading session.
-While TermSight provided multiple reading affordances, participants were in control of deciding whether or not to rely on these features.
-As opposed to a consistent preference for one interaction strategy over others, we observed diverse ways that participants relied or did not rely on TermSight (Figure  14).
-Below, we provide case studies of these non-exclusive emergent reading strategies.
+Five barplots of the minute-by-minute usage of features of TermSight during the ten-minute reading task (i.e., Accessed Document Preview, Click Summary Snippet, Click Highlight Bar, Accessed Phrase Definition, and Accessed Phrase Scenario).
+Participants used these features throughout the reading session, rather than being limited to the beginning or end.
+The Summary Snippet and Document Preview are the most frequently used features throughout.
 Figure 12.
-Scrolling behavior of participants that employed each non-exclusive reading strategy.
+Scrolling behavior of participants who employed each non-exclusive reading strategy.
 TermSight features two scrollable columns: Summary Snippets on the left (red lines) and Original Text on the right (blue lines).
@@
 (b) Original text-driven reading occurs when participants first read the original text and use the Summary Snippets as a supplement (9/20).
+(c) The calibration of trust mainly occurs at the beginning of the session, where participants would compare and contrast the original text to the Summary Snippets with color (6/20).
+Visualization of the scrolling behavior of participants who employed each of the three non-exclusive reading strategies.
+(a) Summary-driven reading occurs when participants primarily read and scroll the Summary Snippets, occasionally clicking on them to interrogate the original text (19/20).
+(b) Original text-driven reading occurs when participants first read the original text and use the Summary Snippets as a supplement (9/20).
 (c) The calibration of trust mainly occurs at the beginning of the session where participants would compare and contrast the original text to the Summary Snippets with color (6/20).
-7.2.2. Summary driven reading
+7.3.2. Reading Strategies for Accessing Original Text
+We observed diverse ways in which participants used TermSight to read ToS (Figure  12).
+Interestingly, these reading strategies centered around the use of the original text.
+In some cases, participants used TermSight’s features as entries into the original text, while in others, the features were used as verification of the original text.
+In still others, participants used features in tandem with the text to calibrate trust in TermSight.
+Below, we provide case studies of these strategies.
+Summary driven reading
 19 out of 20 participants at times relied on the Summary Snippets as a primary entry point for engaging with the text.
-One notable example of this behavior was P3, who mainly read the Summary Snippets and clicked 11 of them to interrogate the original text (Figure  14a).
+One notable example of this behavior was P3, who mainly read the Summary Snippets and clicked 11 of them to interrogate the original text (Figure  12a).
 P3 described how colors were driving his attention to read.
@@
 I clicked on them because I felt like there was more information for me [behind the Summary], and I want to learn more about it.” (P3)
-7.2.3. Original text driven reading
-At other times participants centered their interactions around the original text.
-9 participants went through at least one policy focusing mainly on the original text.
-These participants described how they wanted to first read the actual text before relying on AI-generated summaries due to the potential imperfections of AI.
-These participants explained, “I didn’t want my opinion to be affected by AI” (P5) and “sometimes summaries don’t have all of the information from my other experiences” (P18).
-When reading the Privacy Policy, P8 focused on the original text as shown by the frequent scrolling of the original text in Figure  14b.
+Original text driven reading
+On the other hand, nine participants went through at least one policy focusing mainly on the original text.
+These participants described how they wanted to first read the actual text before relying on AI-generated summaries due to worries about the potential imperfections of AI.
+When reading the privacy policy, P8 focused on the original text as shown by the frequent scrolling of the original text in Figure  12b.
 In addition, P8 clicked the Highlight Bar 11 times to refer back to the Summary Snippets.
@@
 I think it [AI] would try to label it the best it can and I think it did, to be honest.” (P8)
-7.2.4. Calibration of trust
-6 participants described that they went through a calibration process with AI-generated features such as the Summary and Color at the start of the reading session.
+Calibration of trust
+Six participants described that they went through a calibration process with AI-generated features such as the Summary Snippets and color coding.
 For example, P6 described how his reading behavior was different during the calibration process and after.
-P6 described reading the original text on the right side to evaluate the colors and summaries at the beginning of the reading session (original text driven reading).
-Afterward, P6 relied on the summaries a lot more by scrolling and skimming the left panel and clicking on the summary snippets to dive to the original text as seen in Figure  14c (summary driven reading).
-“So like early on in the reading, I kind of read the whole thing on the right first.
+At the beginning of the reading session, P6 focused on reading the original text to evaluate the colors and summaries (original text driven reading).
+Afterward, P6 relied on the summaries a lot more by scrolling and skimming the left panel and clicking on the Summary Snippets to dive to the original text, as seen in Figure  12c (summary driven reading).
+“Early on in the reading, I read the whole thing on the right first.
 Then, I looked at the summaries, and once I realized that it did do a good job summarizing it with colors.
-I trusted it enough to trust it for the rest of the time.” (P6)
-7.2.5. Switching strategy based on time and document.
-ToS includes multiple documents, and in the study, there were 14-15 policies with varying lengths.
-We noticed how participants’ reading behavior can vary based on document length (P5, P19).
-For example, P19 mainly read the summary for longer policies such as User Agreement and Privacy Policy (Summary driven reading).
-For shorter policies such as Buyer Policy, P19 only read the original text (Original text driven reading).
-We also noticed how time can change participants’ reading strategy (P7, P8).
-For example, P7 spent 8.5 minutes and was very interrogative when reading the User Agreement policy.
-When she realized that she was running out of time, she skimmed through the summary snippets for 5 more policies without interrogating the original text.
-7.3. RQ3: How did TermSight influence comprehension and recall?
-Below we present the results of the comprehension quiz (§7.3.1) and recall task (§7.3.2).
-7.3.1. Comprehension
-For the comprehension quiz, out of 6 questions, participants on average scored 2.00 ($\sigma=1.03$) when using TermSight and 2.15 ($\sigma=1.23$) when using the baseline interface.
-Our Bayesian model analysis for the comprehension quiz (model details in Appendix  D.2) reveals no significant differences in the comprehension scores across interface conditions.
-As shown in Figure 15, the posterior distributions for both conditions were nearly identical and centered around zero.
-When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair as shown by the forest plots in Figure 20 in the Appendix, suggesting no significant differences.
-Figure 15.
-The figure shows the posterior distributions of the coefficients for the treatment and control conditions.
-Both posterior distributions are nearly evenly split around zero, indicating no effect of the treatment on the comprehension quiz questions.
-7.3.2. Recall
-For the recall task, participants using TermSight recalled 39 correct facts and 3 incorrect facts in total.
-On the other hand, participants using the baseline interface totally recalled 26 correct facts and 7 incorrect facts.
-Our Bayesian model analysis for the recall task (model details in § D.3) reveals no significant differences in the recall scores across interface conditions.
-Figure 16(b) shows the posterior distribution of the effect size of the treatment on recall ($\mu=0.6$, HPDI = [-0.39, 1.6]).
-Despite the mean being 0.6, suggesting a medium effect size, the HPDI interval has an overlap with the ROPE, implying that there is no significant effect of the treatment on the user recall.
-(a) Posterior distributions of the coefficients for the treatment and control conditions.
-(b) Posterior distribution of the effect size of the treatment on user recall.
-Figure 16.
-Figure 16(a) shows the posterior distributions of the coefficients for the treatment and control conditions.
-Specifically, the posterior distribution of the treatment condition is shifted to the right, indicating a positive influence of the treatment on Recall.
-Figure 16(b) shows the posterior distribution of the effect size of the treatment on the user recall centered around 0.6, suggesting a medium effect size.
-Yet, the HPDI has an overlap with the ROPE of [-0.1, +0.1], implying no significant effect of the treatment on the user recall.
+I trusted it enough for the rest of the time.” (P6)
 8. Discussion
-In this paper, we set out to design a new reading experience that scaffolds readers’ ability to find and understand information important to them within Terms of Service (ToS) contracts.
-Based on formative interviews and observations, we realized this vision in TermSight.
-Results from our evaluation of TermSight (N=20) suggest that TermSight made it easier to read, navigate, and contextualize information within ToS relevant to users.
-Especially interesting was how TermSight afforded multiple different reading strategies (e.g., summary-driven reading, §7.2.2), all of which helped participants overcome the key difficulties of contract reading we observed in our formative study (§3).
-However, similar to prior works on simplifying ToS or other complex documents  (Vu et al., 2007; Tabassum et al., 2018; Good et al., 2005; Korunovska et al., 2020; Robinson and Zhu, 2020; August et al., 2023), TermSight did not worsen or improve the comprehension or recall of the content.
-Below, we discuss the implications of our findings and the TermSight system.
-8.1. Existing ToS are not designed to be read
-Many existing ToS have static or interactive design elements, such as policy names, section headers, and overview summaries (Vu et al., 2007).
-However, participants in our formative study found these features vague, misleading, and ineffective for surfacing and interpreting relevant information (§3.3).
-The HCI literature often refers to these ineffective and deceptive designs as dark patterns in the user interface that manipulate behavior or obscure information to serve the interests of the service providers  (Mathur et al., 2021; Bösch et al., 2016; Kyi et al., 2023).
-TermSight offers one possible design to overcome these dark patterns by offering contract-level guidance (e.g., Power Meter and document previews), document-level guidance (e.g., color-coded Summary Snippets), and phrase-level guidance (e.g., phrase identification).
-Participants in our usability study (§7) affirmed these design goals in both self-report and behavior metrics.
-Participants reported that features of TermSight helped them decide which policy to read (Power Meter, Document Preview) and what text to read within a policy (Summary Snippet).
-Moreover, participants highlighted how TermSight made the original text more approachable with Summary Snippets that broke down and simplified dense sections and Phrase Scope that clarified unfamiliar phrases and helped envision implications.
-Future systems could take these designs further by adapting them to other types of contracts, such as leasing contracts, medical contracts, or employment contracts.
-For example, Summary Snippets may help readers surface and interpret relevant or predatory clauses in leases, while phrase scope may help patients envision the implications of clauses in the medical contract through patient-specific scenarios.
-Lastly, while we designed TermSight as a contract reading interface, TermSight can be readily extended as a browser extension that highlights relevant and important clauses based on user-defined personas to provide in-situ legal awareness and warnings when encountering contracts.
-8.2. Envisioning consequences to legal text
-Prior work has explored providing term definitions to aid comprehension of academic text (August et al., 2023; Head et al., 2021).
-Yet, participants in our study mentioned that definitions alone were not always sufficient for legal language.
-Participants noted that it can be difficult to contextualize abstract phrases in real-world scenarios or anticipate potential consequences—especially those they have not yet personally experienced.
-Our formative study further revealed how people’s intended usage of the service, personal values, and worries can vary.
-While prior work explored predicting unintended consequences for scientific endeavors and AI  (Wang et al., 2024; Pang et al., 2024), TermSight explored the potential value of presenting personalized scenarios to illustrate the possible implications of legal phrases before a user agrees to a ToS.
+In this paper, we explored how AI-powered systems might support ToS contract reading given the unique challenges that legal text poses (i.e., value-dependent, ambiguous, and legally binding).
+Our formative study revealed the barriers ToS presents, which informed the design of an augmented reading interface, TermSight.
+Results from our user study (N=20) suggest that TermSight’s features made ToS contracts easier to read and navigate, without an increase or decrease in comprehension.
+Below, we discuss the implications of our findings around the barriers ToS contracts present (§8.1), the design implications for future contract reading interfaces (§8.2), and the limitations of technology in mitigating the societal difficulties with service contracts (§8.3).
+8.1. From Simplifying Legal Text to Navigating Dark Patterns
+Past work on augmenting legal text has focused on simplification (e.g.,  (Manor and Li, 2019)).
+In the context of ToS contracts, our formative study revealed that the barriers people face extend well beyond complex text.
+Participants struggled to navigate multiple policies within a ToS, surface relevant information from dense legal language, and interpret ambiguous terminology.
+Surprisingly, existing navigational affordances—such as policy names, section headers, or overview summaries—often obfuscated rather than clarified the original text, a design reminiscent of dark patterns in the HCI literature  (Mathur et al., 2021; Bösch et al., 2016; Kyi et al., 2023).
+These barriers may not be unique to ToS, but reflect broader challenges for reading legal contracts taken to their extreme.
+For example, leases, insurance policies, and employment contracts similarly feature multiple addenda; long, dense text; and ambiguous language that obscures consumer rights  (Mueller, 1970; Furth-Matzkin, 2017).
+Future work could expand our findings by conducting large-scale analyses of dark patterns in contracts, an area largely underexplored by the HCI community.
+In addition, future contract reading interfaces should consider going beyond simplifying legal text to helping users navigate contracts while combating the dark patterns of design.
+TermSight serves as an initial exploration for inspiring future contract reading tools in this direction.
+8.2. Design Implications and Opportunities
+TermSight is a preliminary exploration of designing contract reading interfaces. Below, we discuss the design implications of our findings and opportunities for further exploration.
+8.2.1. Providing guidance at multiple levels of granularity
+While prior work focuses on simplifying legal text (e.g.,  (Manor and Li, 2019)), TermSight demonstrated the potential values of guiding readers at the contract-level (e.g., Power Meter and document previews), document-level (e.g., color-coded Summary Snippets), and phrase-level (e.g., phrase identification).
+Participants in our user study affirmed these designs in both self-report and interaction behaviors (§7).
+For example, participants reported how Power Meter transformed deciding which policy to read from a “guessing game” to a more deliberate decision-making process.
+On the other hand, Summary Snippets with color and Phrase Scope drew participants’ attention to snippets of information and phrases buried in dense text that they might otherwise miss.
+Consequently, finding which policy to read and what text to read within a policy was significantly easier for participants with TermSight.
+The specificity of TermSight’s guidance (i.e., relevance classification) relies on the user persona.
+For the user study, we designed two personas to be able to control for their quality (§6.1.2).
+These personas also aligned with participants’ profiles in the user study (Figure  13).
+Future work could explore allowing users to further customize their personas to receive personalized guidance and overcome the barriers of doing so (e.g., forming information bubbles  (Sharma et al., 2024)).
+8.2.2. Prompting reflection on unintended consequences
+Prior work has explored providing term definitions to aid reading academic papers (August et al., 2023; Head et al., 2021).
+However, legal contracts are unique because contracts are agreements for the unforeseen future rather than a description of the past  (Hart, 2017).
+Reflecting this characteristic of contract, participants in our formative study mentioned that definitions alone may not always be sufficient for legal language because it can be difficult to contextualize abstract phrases in unforeseen future scenarios.
+In the user study, participants described the value of Phrase Scope for presenting scenarios to prompt reflection on possible implications of signing a contract, similar to how prior work has encouraged scientists to reflect on the unintended consequences of their research (Wang et al., 2024; Pang et al., 2024).
 Some participants even noted how the scenario was more helpful than the definition (§7.1.5).
-By grounding abstract terms in concrete, relatable examples, these scenarios may help users better reflect on the potential implications behind legal language.
-Future work could investigate additional ways to help consumers reflect on unintended consequences, such as by showing multiple scenarios or relevant legal cases.
-8.3. Design for calibration and trust
-LLM output can be imperfect due to hallucinations or confusing output  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024).
-It is therefore important to help users form appropriate trust and reliance for fostering effective Human-AI collaboration  (Schemmer et al., 2023; Cao et al., 2024).
-In this vein, prior reading interfaces integrating AI have introduced cognitive forcing functions (Buçinca et al., 2021), such as showing summaries only on demand (August et al., 2023) or at the end of the reading session (Chen et al., 2023), to encourage readers to first read the original document.
-Yet, it is unclear if these designs are always optimal for intelligent reading interfaces, since they can increase user frustration and may limit specific reading behaviors  (Buçinca et al., 2021).
-For example, these forcing functions, if applied to TermSight, might introduce significant friction for participants taking a summary-driven reading approach (§7.2.2).
-TermSight took the approach of tightly linking the AI-generated summaries to the original text to allow users to compare and contrast the two.
-In TermSight, users can navigate between and compare the summary and the original text in one click.
-When given these affordances, participants took advantage of them.
-Some participants explicitly evaluated the Summary Snippets to calibrate their trust.
-Interestingly, one participant framed the evaluation of AI-generated output as his ‘due diligence’.
-Even those who took a summary-driven reading approach frequently interrogated the referenced original text as opposed to solely reading the summaries.
-Future work can explore how to help users calibrate trust with the system and evaluate AI output.
-This could include providing explanations of how the reading interface works in general, providing examples of AI imperfections, and providing explanations for individual features  (Schemmer et al., 2023).
-8.4. Ethical Considerations and Social Implications
-Features of TermSight (e.g., Summary Snippets, Classification of Power and Relevance, Term Definitions, and Scenarios) rely on the capabilities of LLMs.
-However, LLMs have been shown to produce imperfect outputs and factually incorrect hallucinations  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024).
-This introduces risks when interacting with LLM-generated text, as they are not part of the contract and are not legally binding.
-For example, incorrect summaries might cause readers to misinterpret the clauses in the contract that are binding.
-On one hand, LLMs have already shown strong capabilities in generating summaries (Zhang et al., 2024; Liu et al., 2024b), term definitions (Yin and Skiena, 2023; Jhirad et al., 2023), and scenarios (Jiang et al., 2024; Wang et al., 2024; Feng et al., 2024; Shaikh et al., 2024).
-Work is also actively investigating how to detect non-factual LLM output  (He et al., 2024) and how to guide LLMs to generate factually correct information  (Vu et al., 2024).
-As a result, LLM output would ideally become more factually correct over time.
-On the other hand, for future deployment of systems like TermSight, we still believe a validation mechanism for LLM output is necessary.
-For example, future implementation of TermSight’s features (e.g., term definitions and scenarios) can be accompanied by additional sources of information when available (e.g., news articles, court cases, or information from other ToS).
-Furthermore, TermSight could facilitate end-user auditing (Lam et al., 2022; Jahanbakhsh et al., 2022), allowing users to report inaccuracies.
-These user contributions could collectively help validate and refine model outputs.
+Future work could further explore how to design these scenarios and investigate their influence on user perception of the contractual clauses, such as by manipulating the linguistic framing (Sharma et al., 2023) or showing information from legal cases and news articles.
+8.2.3. Centering augmentations around the original text
+Legal contracts differ fundamentally from documents like academic papers or news articles because the language itself carries legal power (Chen-Wishart, 2012).
+Consequently, LM-powered text transformations cannot substitute the original text.
+TermSight explored the approach of tightly linking generated text with the original text through visual placement (i.e, side-by-side views) and deep linking (i.e., linking to narrowly attributed text like a single clause), allowing readers to directly compare language and navigate to the original text in at most one click.
+When given these features, participants took advantage of them to engage with the original text (§7.3.2).
+Some participants dove into the original text by finding relevant Summary Snippets, while others verified their understanding of the original text with the Summary Snippets.
+Some participants even explicitly evaluated the Summary Snippets with the original text to calibrate trust in the system.
+An exciting avenue for future work would be to go beyond supporting original text reading to encouraging it.
+Features that nudge readers toward the original text where appropriate, such as revealing summaries only after reading  (Chen et al., 2023), are an interesting start, but little work has evaluated these features’ effects in situations where the original text might be discouraging to engage with.
+8.3. Technological Limitations and Legal Implications
+Compared to simplified reading tasks (e.g., reading a single clause (Robinson and Zhu, 2020; Lee et al., 2025) or policy overview (Tabassum et al., 2018; Korunovska et al., 2020)), our user study involved realistic ToS with over 10 policies.
+TermSight made ToS significantly easier to read and navigate.
+However, TermSight did not worsen or improve comprehension, reminiscent of prior findings that simplified language often did not improve comprehension of privacy policies  (Tabassum et al., 2018; Korunovska et al., 2020; Robinson and Zhu, 2020).
+There may be multiple explanations behind this finding.
+For example, Summary Snippets may have distracted attention away from the original text, the 12-word limit of Summary Snippets may have constrained the conceptual coverage of the original text, and participants’ calibration behaviors (i.e., confirming summary information with the original text) may have distracted them from the reading task.
+However, our findings may have also revealed the broader limitations of technological solutions in mitigating the societal difficulties with service contracts such as ToS.
+Although TermSight offered multi-level guidance on relevance (e.g., Power Meter, Summary Snippets), the sheer volume of potentially relevant information in ToS remains a dominant challenge.
+For example, in the social media ToS, 172/759 (23%) discrete Information Snippets were classified by TermSight as being relevant to the given user persona (§6.1.2), including 6 snippets needed to answer the comprehension questions.
+Within this subset, being able to notice and recall snippets of information to correctly answer the comprehension questions may be difficult.
+While future deployment of TermSight could allow users to customize the personas for more targeted guidance, this may risk introducing information bubbles  (Adamic and Glance, 2005; Conover et al., 2021; Garrett, 2009; Sharma et al., 2024).
+Given that the formation of a binding contract under U.S.
+law is premised on the doctrine of the ‘duty to read’ [^6] Contracting parties are legally responsible for reading and understanding the contract before signing (Benoliel and Becher, 2019)., our findings raise fundamental questions about the practicality of this duty for ToS.
+Instead of relying solely on technological fixes, policy and legal innovations are needed.
+For example, policy makers and legal scholars could explore alternative paradigms of contracting that operationalize ToS into shorter, context-specific contracts to reduce the upfront burden to consent.
 8.4. Limitations and Future Work
-The generalizability of TermSight and our study findings is limited by the challenges associated with handling ToS documents and our sample of participants in the user study.
-TermSight currently relies on a set of precleaned Markdown or HTML files.
-Future works can investigate more advanced scraping and cleaning methods.
-In addition, our recruited participants were mostly college-educated and fluent in English.
-Marginalized communities, such as neurodivergent readers and individuals with limited English literacy, can face additional barriers when interpreting legal documents.
+The nature of this work is exploratory.
+Similar to prior HCI studies (e.g.,  (Chen et al., 2023; Cao et al., 2023; Kambhamettu et al., 2024)), we designed and studied TermSight to explore the challenges and opportunities of designing contract reading interfaces, rather than proposing TermSight as the definitive solution.
+Future work can pursue finer-grained exploration through controlled ablation studies.
+Examples may include investigating the influence of Power Meter on ToS reading frequency in the wild, assessing how hypothetical scenarios affect user perception of clauses (e.g., trust  (Shneiderman, 2000)), or investigating how AI affordances (e.g., Summary Snippets) support or distract from original text reading (e.g, (Nielsen et al., 2023)) through eye-tracking.
+Because participants were paid to read ToS in a timed reading task, their motivation and behavior may not reflect real-world usage.
+Moreover, a within-subject design may lead to ordering effects or fatigue.
+In the study, we counterbalanced the conditions and didn’t observe a significant effect of ordering on user experience, comprehension, and recall (Appendix Figure  16).
+Yet, it may still have influences on participants’ interaction.
+Future work can expand our study by conducting between-subject or field experiments.
+In addition, our participants were recruited from Prolific, mostly college-educated, and fluent in English, which may limit the generalization of our findings to the broader population.
+Marginalized communities, including neurodivergent readers and individuals with limited English literacy, can face additional barriers when interpreting legal documents.
 For example, readers with ADHD might have a hard time staying focused when reading visually dense text such as ToS  (Stern and Shalev, 2013; Bental and Tirosh, 2007).
-Future works can investigate how to help marginalized populations make sense of legal contracts by evaluating TermSight with these populations of users or integrating additional supports such as translation.
+Future work can investigate how to help alternative reader populations make sense of legal contracts by evaluating TermSight with these populations of users or integrating additional supports (e.g., translation).
+Future systems could adapt our design of TermSight to other contracts.
+For example, Summary Snippets may help readers surface and interpret relevant or predatory clauses in leases, while Phrase Scope may help patients envision the unintended consequences of signing a medical contract through patient-specific scenarios.
+Finally, features of TermSight rely on the capabilities of LMs, which may produce imperfect outputs  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024).
+While work is actively investigating how to guide LMs to generate factually correct information  (Vu et al., 2024), we believe a validation mechanism for LLM output is necessary for future deployment of systems like TermSight, such as by facilitating end-user (Lam et al., 2022; Jahanbakhsh et al., 2022) or expert auditing of AI output (Jiang et al., 2024).
 9. Conclusion
-ToS are service contracts outlining the rules and expectations for both the service provider and the consumer.
-Our formative study revealed ineffective and deceptive designs at all granularities of a ToS.
-To make ToS approachable for consumers, we designed and evaluated an intelligent reading interface: TermSight.
+We regularly sign legal contracts for where we live, who we work with, and how we interact with digital services.
+Yet, these contracts have drifted from their premise of facilitating mutual understanding to complex documents that discourage reading.
+In this work, we explored the opportunities and limitations of intelligent reading support for legal contracts, using ToS contracts as a case study.
+Our formative study revealed ineffective and deceptive designs at all levels of a ToS.
+To make service contracts approachable, we designed and evaluated an intelligent reading interface: TermSight.
 Participants reported that TermSight reduced the difficulty of engaging with ToS and increased their willingness to do so.
-Participants highlighted that features of TermSight afforded them the ability to surface and interpret relevant information at all levels of ToS reading.
-They found the Summary Snippets with colors particularly helpful in breaking dense sections into snippets of simplified information, which helped them surface relevant information and power balance.
-The ‘duty to read’ originated when contracts were brief and based on mutual understanding and negotiation  (Kar and Radin, 2019; Benoliel and Becher, 2019).
-Modern ToS contracts, however, prioritize unilateral control over mutual agreement by design.
-TermSight presents a vision and serves as a technology probe into making legal contracts more accessible to the general public, providing consumers with the tools to make more informed legal decisions.
-Appendix A Iterative Design
+Additionally, features of TermSight enabled participants the ability to surface and approach relevant information at all levels of contract reading.
+Taken together, TermSight presents one avenue for making legal contracts more approachable to the general public.
+However, facilitating a deeper understanding of ToS remains an open challenge due to the overwhelming amount of information in ToS.
+Appendix A Iterative Design Study
 A total of 8 participants were recruited through Prolific to evaluate an early prototype of TermSight.
 This prototype closely resembled the current version of TermSight but lacked a document preview when hovering over the Power Meter.
-Additionally, the prototype used three saturation levels for relevance, resulting in a total of nine colors (three colors for power X three saturation levels for relevance).
+Additionally, the prototype used three saturation levels for relevance, resulting in a total of nine colors (3 hues for power x 3 saturation levels for relevance).
 Participants could also toggle between two layouts for the Summary Snippets: condensed layout and in-context layout.
@@
 In contrast, the in-context layout placed each Summary Snippet directly next to its corresponding Information Snippet.
-After a brief interface tutorial, participants were given 10 minutes to read a ToS using the interface followed by a semi-structured interview about their experience.
+After a brief interface tutorial, participants were given 10 minutes to read a ToS using the interface, followed by a semi-structured interview about their experience.
 Each session lasted for 40 minutes.
-Below we present the main findings and changes made.
-General findings.
-All participants described that reading with the prototype was easier and more engaging compared to typical ToS, which they tend to skip entirely or skim superficially.
-In contrast to their usual approach of skimming and trying to search for keywords that catch their eyes, all participants described how the Summary Snippets directly highlighted relevant information otherwise buried in dense sections.
-Additionally, participants expressed interest in reading the original text after using the Summary Snippets to surface relevant information and valued the ability to navigate between the summary and the original text.
-6 participants found the scenarios provided by Phrase Scope to be particularly helpful as the definitions alone can be difficult to contextualize and envision implications.
+Below, we present the main findings and changes made.
 Participants preferred condensed layout for quick navigation
@@
 In contrast, the in-context layout required significantly more scrolling, which participants found overwhelming.
-As a result, in the final version of TermSight, we used the condensed layout for the Summary Snippets (Figure  1).
+As a result, in the final version of TermSight, we used the condensed layout (Figure  1).
 Reduce visual complexity of the color scheme
 All participants found the colors to be intuitive and effective in highlighting power and relevance, helping them decide what to read.
-However, 3 participants pointed out that while the power dimension was intuitive, the relevance dimension with 3 saturation levels was challenging to differentiate as there would be 9 colors in total.
+However, 3 participants pointed out that while the power dimension was intuitive, the relevance dimension with 3 saturation levels was challenging to differentiate, as there would be 9 colors in total.
 To reduce visual complexity, the final version of TermSight included only two saturation levels (High vs.
@@
 This feedback led to the integration of a document preview feature in the final version of TermSight, which appears when users hover over the Power Meter (Figure  3).
-Appendix B Additional Implementation Details of TermSight
-In this section, we provide additional implementation details of TermSight.
+Appendix B Additional Implementation Details
 B.1. Source Document and Pre-processing
@@
 B.2. Obtaining Summary Snippets and Information Snippets
-The prompt used to obtain the Summary Snippets and Information Snippets is detailed in Figure  21.
+The prompt used to obtain the Summary Snippets and Information Snippets is detailed in Figure  17.
 We constrained the Summary Snippets to be short because prior works have found that adding short summaries (10-20 words) under search results was more effective for navigational and information-seeking tasks compared to longer summaries, which can be harder to skim  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006).
-This length constraint was further optimized through prompt engineering.
-We noticed that setting summaries shorter than 12 words led to excessive fragmentation (i.e., too many summary snippets), while longer summaries tried to cover too much information, making them visually dense and less skimmable.
 B.3. Classifying Information Snippets
-In TermSight, two classifications were performed for each Information Snippet using GPT-4o with few-shot prompting (Figure  22 and  23).
+In TermSight, two classifications were performed for each Information Snippet using GPT-4o with few-shot prompting (Figure  18 and  19).
 For the classification of Power, each Information Snippet was classified by the degree of control or agency it grants to the Service Provider or the User (Categories: Service, Neutral, User).
 For the classification of Relevance, Information Snippets were classified based on their relevance to the user persona (Categories: High, Low).
-User personas used for the study are included in Figure  17 and  18.
+User personas used for the study are included in Appendix  E.2.
 B.4. Phrase Scope
-B.4.1. Identifying unfamiliar or vague phrases
-For each document chunk, we employ two few-shot prompts to GPT-4o to identify potentially unfamiliar (Figure  24) or vague phrases (Figure  25).
-Few-shot examples for identifying unfamiliar phrases were selected based on phrases participants found challenging in our formative study, which included legal jargon, copyright licenses, and privacy-related terminology.
-Few-shot examples for identifying vague phrases were informed by vague phrases participants identified in the formative study, as well as prior works analyzing vagueness in Terms of Service (ToS) documents  (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022).
+B.4.1. Identifying unfamiliar or ambiguous phrases
+For each document chunk, we employ two few-shot prompts to GPT-4o to identify potentially unfamiliar (Figure  20) or vague phrases (Figure  21).
+Few-shot examples for identifying unfamiliar and ambiguous phrases were selected based on phrases participants found challenging in our formative study.
 Both prompts were applied to each document chunk, producing two sets of identified phrases.
@@
 B.4.2. Generating phrase definitions and answers to user questions
-We used a retrieval-augmented question answering approach  (Fok et al., 2024a) to generate the phrase definitions.
-Figure  26 shows the prompt used to generate the definitions after retrieving potentially relevant document chunks.
-The prompt explicitly instructs GPT-4o to produce a concise in-context definition of the input phrase and output a list of the retrieved chunks referenced in generating the definition.
-In the case when the definition is generated without referencing any of the retrieved chunks, it is still retained.
-TermSight intends to provide plausible definitions even when explicit definitions are absent from the document.
-This design decision is also supported by prior research demonstrating the capabilities of large language models in generating term definitions that are comparable to traditional dictionaries  (Yin and Skiena, 2023; Jhirad et al., 2023).
-When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked (Figure  28).
-If the answer is not specified in the retrieved chunks, the system generates a plausible answer based on the phrase context, similar to the definitions.
+We first retrieve potentially relevant document chunks from the vector database.
+The database query is framed as a question: “What does {input phrase} refer to in the sentence: {phrase context}.
+Here, phrase context refers to the chunk of text containing the input phrase.
+Then, the query is embedded using OpenAI’s text-embedding-3-small model, and the top 15 chunks from the vector database are retrieved based on cosine similarity.
+Figure  22 shows the prompt used to generate the definitions after retrieving potentially relevant document chunks.
+When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked (Figure  24).
 B.4.3. Generating scenarios
-Our needfinding reveals the need to help readers not only interpret the definition of a phrase but to envision potential implications.
-As a result, we leveraged GPT-4o to generate customized scenarios of potential implications based on user persona.
-The prompt is specified in Figure  27.
+We leveraged GPT-4o to generate customized scenarios of potential implications based on user persona. The prompt is specified in Figure  23.
+Classification Input Snippet Output Imperfection
+Power “All Buy Now purchases in a ServiceY Show are final and binding.” Service Lack of enough context in the input. Users do have the option to return. It’s a neutral clause.
+Power “Use public content for any illegal, deceptive, unethical, false, misleading, or improper purpose, including the infringement of third-party intellectual property rights.” Neutral Lack of enough context in the input.
+The clause describes what third party licensee cannot do with user content.
+It’s a user-benefiting clause.
+Relevance “Your earnings are based on the listing price and actual earnings will vary based on the final order price, Seller discounts, and any other applicable taxes and discounts.” High Lack of enough context in the input.
+This clause is for sellers and is less relevant to the input buyer persona.
+Relevance “ServiceY cannot guarantee that a ServiceY consignment listing will be sold or that a certain sales amount will be earned for individual items or an entire shipment.” High Lack of enough context in the input.
+This clause is for sellers and is less relevant to the input buyer persona.
+Table 1. Examples of imperfect classifications of Power and Relevance.
+Examples of imperfect classifications of Power and Relevance. The four columns describe the type of classification (Relevance or Power), model input, model output, and the imperfection.
 Appendix C Evaluation of TermSight Output
-We manually evaluated TermSight’s feature pipelines and overall performance.
-Our goal was not to assess advances in system performance, but to verify that the system reliably produced meaningful outputs that could support ToS reading.
-C.1. Evaluating the Classification of Information Snippets
-We sampled 116 Information Snippets with stratified sampling from all the sub-policies for each service included in the user study.
-For each Summary Snippet, we evaluated the classification of power and relevance.
-Of the 116 Information Snippets, 27 were classified by TermSight as favoring the user, 22 were classified as favoring the service, and 67 were classified as neutral.
-In addition, 59 were classified as of high relevance to the provided user persona and 57 were classified to be of low relevance.
-Our evaluation revealed 3 instances where the power classifications were imperfect, mainly caused by the lack of context in the input information snippet.
+C.1. Classification of Information Snippets
+Out of the sampled 116 snippets, our evaluation revealed 3 instances where the power classifications were imperfect, mainly caused by the lack of context in the input Information Snippet.
 For example, the snippet (‘All Buy Now purchases in a ServiceY Show are final and binding’) was classified by TermSight as favoring the service.
@@
 However, users can return and get a refund if there are problems with the purchase (e.g., an item doesn’t match its description) as stated in the service’s return policy, making this snippet more neutral.
-The return policy was not included in the snippet context, as a result, this neutral statement was considered more service-oriented.
+The return policy was not included in the snippet context; as a result, this neutral statement was considered more service-oriented.
+We didn’t observe any clause favoring the service provider being misclassified as neutral or favoring the user.
 For the classification of relevance, there were 2 instances, out of 116, where irrelevant Information Snippets were classified as relevant to the persona.
@@
 For example, sellers have to pay a fee to the platform after selling an item.
-Despite this information being more relevant to sellers, TermSight classified it as relevant to buyers as it exposed the potential hidden fees that sellers might include as part of their listing price.
-C.2. Evaluating Term Definitions and Scenarios
-We randomly sampled 113 generated definitions and scenarios using stratified sampling from all the sub-policies for each of the two ToS augmented for the user study.
-These were evaluated by members of the research team.
-Out of the 113 sampled phrases, 46 were identified by TermSight as vague phrases, 6 were identified as both jargon and vague phrases, and 61 were identified as jargon.
-Our evaluation revealed that all the generated definitions were correct, but out of the 113 definitions, 4 were only general definitions of the phrase not specific to the ToS.
+Despite this information being more relevant to sellers, TermSight classified it as relevant to buyers, as it exposed the potential hidden fees that sellers might include as part of their listing price.
+We show examples of imperfect classification of Power and Relevance in Table  1.
+C.2. Term Definitions and Scenarios
+Our evaluation revealed that all the generated definitions were correct, but out of the 113 definitions, 4 were only general definitions of the phrase, not specific to the ToS.
 On further inspection, we identified that this was because these vague phrases were not explicitly defined anywhere in the ToS.
@@
 For example, for the phrase ‘aggregated anonymized statistics’, TermSight provided a general definition of what it might mean to aggregate and anonymize user data.
-However, the specific details—such as what user data are being aggregated—were not specified in the ToS.
+However, the specific details—such as what user data is being aggregated—were not specified in the ToS.
 Rather than a limitation of TermSight, these imperfections highlight the ill-defined nature of the language used in ToS.
@@
 We also noticed that when the passages or phrases target a different audience (e.g., developers) than the user persona used to generate the scenarios (i.e., a lay user), the generated scenarios become less relevant or useful.
-We did not regenerate these imperfections to keep the user experience realistic to real-world settings where the LLM outputs are not guaranteed to be perfect.
+We did not regenerate these imperfections to keep the user experience realistic in real-world settings where the LLM outputs are not guaranteed to be perfect.
 Participants in our user study were informed that AI output can be imperfect.
 This specific incorrect scenario was not accessed by any participants in the study.
-Appendix D Bayesian Analysis Model Details
-In this section, we provide additional details on the models used in the paper.
-All models used the DAG in Figure 9 to inform the model.
-Next, in § D.1, we describe the model used to estimate the effect of the treatment on the responses related to the experience of using the treatment interface.
-In § D.2, we describe the model used to estimate the effect of the treatment on the comprehension outcomes i.e., the quiz questions.
-Finally, in § D.3, we describe the model used to estimate the effect of the treatment on the experiment subjects’ ability to recall facts correctly.
-Figure 19 and  Figure 20 show the forest plots of model coefficients for the experience and comprehension outcomes.
-D.1. Modeling Experience Outcomes
-We asked the experiment subjects to rate their experience of using the treatment or the baseline interface on a Likert scale from 1 to 5 (L=5 outcome values).
-There were N subjects and we asked each subject K (K=7) questions.
-Since the likert scale is ordinal, we modeled the experience outcomes using an ordinal regression model.
-The DAG in Figure 9 indicates that the experimental condition alone affects the experience outcomes.
-There were three randomizations: whether the subject was shown the treatment or control interface, and whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second.
-We modeled the experience outcomes using a cumulative logit model with a linear link function.
-The model is given by:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-In eq.
-7, we further ensured that the samples are ordered i.e., $\kappa_{1}<\kappa_{2}<\dots<\kappa_{L-1}$.
-The priors are conservative.
-For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to outcome probability range of [0.04, 0.95].
-D.2. Modeling Comprehension Outcomes
-Like earlier, there were N subjects, and we asked each subject K (K=6) questions to test their comprehension of the ToS.
-Thus, the outcome is a binary variable with 1 indicating a correct answer and 0 indicating an incorrect answer.
-The DAG in Figure 9 indicates that the experimental condition alone affects the comprehension outcomes.
-There were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second.
-We modeled the comprehension outcomes using a logistic regression model.
-Notice that we created a joint variable $c_{j,k,l}$ for the coefficients corresponding to the question, service, and treatment/control.
-We did this because the questions were different across the two services used in the experiment.
-We included the treatment/control variable to be able to easily estimate the effect of the treatment on the comprehension outcomes per question.
-The model is given by:
-[…]
-[…]
-[…]
-[…]
-[…]
-As in the previous model, the priors are conservative.
-For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to outcome probability range of [0.04, 0.95].
-D.3. Modeling Recall Outcomes
-There were N subjects, and we asked each subject a question on recall facts from the ToS.
-The outcome that we measure is the number of correctly recalled facts, making the outcome a count variable.
-Since many of the subjects could not recall any facts correctly, we modeled the recall outcomes using a Zero Inflated Poisson regression model.
-The DAG in Figure 9 indicates that the experimental condition alone affects the comprehension outcomes.
-As before, there were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second.
-The model is given by:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-As in the previous model, the priors are conservative.
-For example, on the Log scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse log to outcome range of [0.05, 21].
-Notice that we model the number of correctly recalled facts, and thus is a conservative prior.
-Given that most subjects could not recall any facts, the zero inflation parameter is set to a weakly informative prior using a Beta distribution.
-Appendix E Study Materials: Formative Study
+Appendix D Study Materials: Formative Study
 Questions asked before the reading session:
@@
   • How did you go about reading the ToS?
-  • What information were you interested in in ToS? Both from your prior experience in reading ToS or this ToS?
-  • Imagine if you have a magic wand that can transform the Terms of Service in whatever ways you want. How would you transform the Terms of Service?
-Appendix F Study Materials: User Study
-F.1. User Persona
-Two personas were given to the participants during the user study for the social media service (Figure  17) and e-commerce service (Figure  18).
+  • What information were you interested in in ToS? Both from your prior experience in reading ToS and this ToS?
+  • Imagine if you have a magic wand that can transform the ToS in whatever ways you want. How would you transform the Terms of Service?
+Appendix E Study Materials: User Study
+E.1. Baseline Interface
+The baseline ToS reading interface used in the user study can be found in Figure  25.
+E.2. User Persona
+Two personas were given to the participants during the user study for the social media service (content consumer who posts personal content) and the e-commerce service (buyer who rarely posts reviews).
 The same personas were used for features of TermSight to classify relevance and generate personalized scenarios.
-The personas were designed based on information participants in the formative study described caring about.
-User Persona for Social Media site: Content consumer who posts personal content Imagine you are a lay user of social media platforms.
+The personas were designed based on information that participants in the formative study described caring about.
+Persona: Content consumer who posts personal content Imagine you are a lay user of social media platforms.
 You are over 18 years old and located in the United States.
 Your usage of Social Media sites: You spend most of your time on the platform scrolling through feeds, liking posts, chatting with other users, and sharing personal content such as photos.
-Things you care about when using Social Media Sites: You care about Privacy, particularly what data are being collected and how your data can be used and shared.
-You care about what the service can do with user-generated content such as licenses over user content or advertising with user content.
+Things you care about when using Social Media Sites: You care about Privacy, particularly what data is being collected and how your data can be used and shared.
+You care about what the service can do with user-generated content, such as licenses over user content or advertising with user content.
 You care about potential liabilities when using ServiceX.
-Figure 17. User Persona for Social Media site: Content consumer who posts personal content
-User Persona for E-commerce site: Buyer who rarely posts reviews Imagine you are a lay user of E-commerce platforms.
+Persona: Buyer who rarely posts reviews Imagine you are a lay user of E-commerce platforms.
 You are over 18 years old and located in the United States.
@@
 Things you care about when using E-commerce sites: You care about information related to making purchases, refunds, returns, user protection policies, termination, arbitration, and liabilities.
-You also care about Privacy, particularly what data are being collected and how your data can be used and shared.
-Figure 18. User Persona for E-commerce site: Buyer who rarely posts reviews
-F.2. Pre-survey Questions
-Questions that were asked only once at the beginning of the interview:
-  • For how many online platforms have you read their Terms and Conditions (T&C) before? [None (0), Few (1-3), Some (4-6), Many (7-9), A lot (¿10)]
-  • How familiar are you with Terms and Conditions (T&C) for online platforms? (5-point likert rating)
+You also care about Privacy, particularly what data is being collected and how your data can be used and shared.
+E.3. Participant Demographics
+Of the 20 participants, 13 self-identify as female and 7 as male.
+1 participant had a high school degree, 3 had an associate degree, 12 had a bachelor’s degree, and 4 had a master’s degree.
+4 participants were between the ages of 18-25, 5 were between 26-35, 5 were between 36-45, and 6 were between 46-55.
+When asked about how many ToS they’ve read, 4 participants read none, 5 read between 1–3, 5 read between 4–6, 1 read between 7–9, and 5 read greater than 10.
+Participants in the user study also found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  13).
+E.4. Pre-survey Questions
+Questions that were asked once at the beginning of the interview:
+  • For how many online platforms have you read their Terms of Service (ToS) before? [None (0), Few (1-3), Some (4-6), Many (7-9), A lot (¿10)]
+  • How familiar are you with Terms of Service (ToS) for online platforms? (5-point Likert rating)
 5-point Likert rating questions that were asked before each of the two reading sessions:
@@
   • How well does the above user persona align with things you personally care about when using (E-commerce or Social Media) sites?
-F.3. 5-point Likert Ratings of Reading Experience
+Figure 13.
+Participants’ self-ratings of their familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service.
+Most participants were familiar with both social media and e-commerce platforms and found the given persona to align with their personal usage of the service and personal value.
+Participants’ self-ratings of their general familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service.
+Most participants were familiar with both social media and e-commerce platforms in general and found the given persona to align with their personal usage of the service and personal value.
+E.5. 5-point Likert Ratings of Reading Experience
 Participants rated their reading experience after each of the two reading sessions.
-  • E1: I’m interested in spending more time reading the service’s Terms and Conditions (T&C) with the current interface and wish to get a link after the study.
-  • E2: How hard did you have to work to read the T&C?
+  • E1: I’m interested in spending more time reading the service’s Terms of Service (ToS) with the current interface and wish to get a link after the study.
+  • E2: How hard did you have to work to read the ToS?
   • E3: How easy was it for you to decide which sub-policies to read?
   • E4: How easy was it for you to decide what text to read within a sub-policy?
-  • E5: How confident are you that you got all the relevant information from the T&C (including sub-policies)?
-  • E6: How much do you feel like you understood the T&C?
-  • E7: How much would you be willing to read the T&C of other services with this interface?
-F.4. Semi-structured Interview Questions
-Below are questions that were asked after each of the two reading sessions, except the last question which was only asked after the second reading session where participants would compare and contrast both interfaces.
-  • Describe your experience reading the Terms of Service using the interface.
+  • E5: How confident are you that you got all the relevant information from the ToS (including sub-policies)?
+  • E6: How much do you feel like you understood the ToS?
+  • E7: How much would you be willing to read the ToS of other services with this interface?
+E.6. Semi-structured Interview Questions
+Below are questions that were asked after each of the two reading sessions, except the last question.
+The last question was only asked after the second reading session, where participants would compare and contrast both interfaces.
+  • Describe your experience reading the ToS using the interface.
   • How did you read the ToS using the interface?
@@
   • Which interface do you prefer? Why? Compare and contrast. (Asked only after the second reading session)
-Question Correct Answer Relevant Passage
-How can ServiceX use photographs you post for advertisements or promotions?
-ServiceX can use photos you post in ads without your permission and is not obligated to attribute you as the creator.
-[Social Media] “…you grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content and any name, username, voice, or likeness provided in connection with Your Content in all media formats and channels now known or later developed anywhere in the world…you irrevocably waive any claims and assertions of moral rights or attribution with respect to Your Content.” (User Agreement)
-ServiceX licenses the content you post to third parties.
-Which of the following is a third-party licensee of user content on ServiceX allowed to do?
-Train AI models with user content without user consent.
-[Social Media] “For example, this license includes the right to use Your Content to train AI and machine learning models” (User Agreement) Our licensees cannot: Combine third-party data with ServiceX public content to target ads; Perform background checks; Track, alert, monitor, or investigate sensitive events (for example, protests or rallies) or sensitive organizations (for example, unions or activist groups)…” (Public Content Policy)
-You bought a pair of shoes from ServiceY.
-Before accepting the delivery, you tried the shoes on but found the shoes small for your foot.
-What can you do to return the item and guarantee a refund?
-ServiceY does not accept “item doesn’t fit” as a valid reason for return.
-[E-commerce] “Trades, offline transactions, items that do not fit, or orders where the Buyer changed their mind are not covered by ServiceY Protect.
-These types of purchases are considered final sale and not eligible for return.” (Return Policy)
-According to ServiceY’s authentication policy, when you buy an item over 500\$, ServiceY will first authenticate the item before shipping it to you.
-If ServiceY decides that the item is fake, what would happen?
-The item will be disposed.
-You will be refunded.
-[E-commerce] “If we reasonably believe an item to be counterfeit, we will provide a refund for the full cost of the item and any accompanying shipping costs to the buyer … if an item in question is deemed counterfeit, we shall ensure disposal of the item at our discretion.” (Counterfeit Items Policy)
-Table 1.
-Example multiple choice comprehension questions: The questions were designed to be difficult and can’t be directly answered by only reading the Summary Snippets.
-These questions can also require knowledge about information from multiple positions of the same policy or information across different policies.
+Appendix F Bayesian Analysis Model Details and Additional Analysis
+In this section, we provide details on the Bayesian models.
+We first explain the concepts of structural causal framework and DAG in §F.1.
+All models were informed by the DAG in Figure  6.
+Next, in §F.2, we describe the model used to estimate the effect of the treatment on the user experience.
+In §F.3, we describe the model used to estimate the effect of the treatment on the comprehension outcomes, i.e., the quiz questions.
+Finally, in §F.4, we describe the model used to estimate the effect of the treatment on the experiment subjects’ ability to recall facts correctly.
+Figure  14 and  15 show the forest plots of model coefficients for the experience and comprehension outcomes.
+Figure  16 presents forest plots illustrating the effect of ordering on user experience, comprehension, and recall.
+F.1. Structural Causal Framework
+We used a structural causal framework popularized by Pearl (2009) to better understand the effects of the treatment.
+A structural causal model framework comprises exogenous variables $U$, endogenous variables $V$, and a set of functions $F$ that represent the physical process that results in the values of the endogenous variables.
+For example, the question $v_{i}=f(v,u)$ implies that the values of $v$ and $u$ completely determine the value of $v_{i}$ through the function $f$.
+The structural causal model can be represented as a Directed Acyclic Graph (DAG) where the nodes represent the variables and the edges represent the causal relationships among the variables.
+Thus, in the preceding example, $v$ and $u$ are the parents of $v_{i}$ in the DAG.
+The structural causal model framework is non-parametric, but frequently used with generalized linear models to estimate the functions $f$.
+The structural causal framework has connections with structural equation models (Wright, 1934) used in the Economic Sciences, though the latter were developed with linear relationships in mind.
+Figure  6 shows the DAG for the study.
+F.2. Modeling Experience Outcomes
+We asked the experiment subjects to rate their experience of using the treatment or the baseline interface on a Likert scale from 1 to 5 (L=5 outcome values).
+There were N subjects, and we asked each subject K (K=7) questions.
+Since the Likert scale is ordinal, we modeled the experience outcomes using an ordinal regression model.
+The DAG in Figure  6 indicates that the experimental condition alone affects the experience outcomes.
+There were three randomizations: whether the subject was shown the treatment or control interface (i.e., interface type), whether the subject was shown the social media ToS or the e-commerce ToS (i.e., service type), and whether the treatment was shown first or second (i.e., presentation order).
+We modeled the experience outcomes using a cumulative logit model with a linear link function:
+(1) $Y_{i,j,k,l,m}$ $\sim\mathrm{OrderedLogistic}(\kappa,\phi_{i,j,k,l,m})\quad$
+(2) $\phi_{i,j,k,l,m}$ $=\alpha_{i}+c_{j,k}+s_{l}+o_{m}\quad$
+Below, we specify the priors for $\alpha_{i}$ (intercept, one per participant), $c_{j,k}$ (joint variable for the interface type and user experience item), $s_{l}$ (service type), $o_{m}$ (presentation order), and $\kappa_{i}$ (cutpoints).
+(3) $\alpha_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,N\}\quad$
+(4) $c_{j,k}$ $\sim\mathrm{Normal}(0,1),j\in\{1,2\},k\in\{1,\dots,K\}\quad$
+(5) $s_{l}$ $\sim\mathrm{Normal}(0,1),l\in\{1,2\}\quad$
+(6) $o_{m}$ $\sim\mathrm{Normal}(0,1),m\in\{1,2\}\quad$
+(7) $\kappa_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,L-1\}\quad$
+In eq (7), we further ensured that the samples are ordered, i.e., $\kappa_{1}<\kappa_{2}<\dots<\kappa_{L-1}$.
+The priors are conservative.
+For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to the outcome probability range of [0.04, 0.95].
+F.3. Modeling Comprehension Outcomes
+There were N subjects, and we asked each subject K (K=6) questions to test their comprehension.
+The outcome is a binary variable with 1 indicating a correct answer and 0 indicating an incorrect answer.
+The DAG in Figure  6 indicates that the experimental condition alone affects the comprehension outcomes.
+There were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the e-commerce ToS, and whether the treatment was shown first or second.
+We modeled the comprehension outcomes using a logistic regression model.
+Notice that we created a joint variable $c_{j,k,l}$ for the coefficients corresponding to the interface type, question, and service type.
+We did this because the comprehension questions were different across the two services used in the experiment.
+We included the interface variable (i.e., treatment vs.
+control) to be able to easily estimate the effect of the treatment on the comprehension outcomes per question.
+The model is given by:
+(8) $Y_{i,j,k,l,m}$ $\sim\mathrm{Binomial}(p_{i,j,k,l,m})\quad$
+(9) $\mathrm{Logistic}(p_{i,j,k,l,m})$ $=\alpha_{i}+c_{j,k,l}+o_{m}\quad$
+Below, we specify the priors for $\alpha_{i}$ (intercept, one per participant), $c_{j,k,l}$ (joint variable for the interface type, comprehension question, and service type), and $o_{m}$ (presentation order).
+(10) $\alpha_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,N\}\quad$
+(11) $c_{j,k,l}$ $\sim\mathrm{Normal}(0,1),j\in\{1,2\},k\in\{1,\dots,K\},l\in\{1,2\}\quad$
+(12) $o_{m}$ $\sim\mathrm{Normal}(0,1),m\in\{1,2\}\quad$
+As in the previous model, the priors are conservative.
+For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to the outcome probability range of [0.04, 0.95].
+F.4. Modeling Recall Outcomes
+There were N subjects, and we asked each subject to recall facts from the ToS.
+The outcome that we measure is the number of correctly recalled facts, making the outcome a count variable.
+Since many of the subjects could not recall any facts correctly, we modeled the recall outcomes using a Zero Inflated Poisson regression model.
+The DAG in Figure  6 indicates that the experimental condition alone affects the comprehension outcomes.
+As before, there were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second.
+The model is given by:
+(13) $Y_{i,j,k,l}$ $\sim\mathrm{ZeroInflatedPoisson}(\lambda_{i,j,k,l},\phi)\quad$
+(14) $\mathrm{Log}(\lambda_{i,j,k,l})$ $=\alpha_{i}+c_{j}+s_{k}+o_{l}\quad$
+Below, we specify the priors for $\alpha_{i}$ (intercept, one per participant), $c_{j}$ (interface type), $s_{k}$ (service type), $o_{l}$ (presentation order), and $\phi$ (probability of zero inflation).
+(15) $\alpha_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,N\}\quad$
+(16) $c_{j}$ $\sim\mathrm{Normal}(0,1),j\in\{1,2\}\quad$
+(17) $s_{k}$ $\sim\mathrm{Normal}(0,1),k\in\{1,2\}\quad$
+(18) $o_{l}$ $\sim\mathrm{Normal}(0,1),l\in\{1,2\}\quad$
+(19) $\phi$ $\sim\mathrm{Beta}(2,2)\quad$
+As in the previous model, the priors are conservative.
+For example, on the Log scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse log to the outcome range of [0.05, 21].
+Notice that we model the number of correctly recalled facts, and thus is a conservative prior.
+Given that most subjects could not recall any facts, the zero inflation parameter is set to a weakly informative prior using a Beta distribution.
+Appendix G Supplemental Analysis for Comprehension Outcomes
+We aimed to design the comprehension questions to have a single answer that best matches the clauses in the original text.
+However, we acknowledge that legal problems and contractual clauses may leave room for interpretation  (Marotta-Wurgler and Stein, 2025).
+As a result, we run an additional, more conservative analysis by removing any question that may contain room for interpretation on whether the original text would apply to the situation being asked in the question, leading to different answer choices.
+We removed 2 out of 6 questions for each service type (Social Media: Q1, Q4; Shopping: Q2, Q3).
+For all other questions, there is a single correct answer choice directly matching the original text, while all the other options either conflict with or were never mentioned in the original text.
+Below is one example of the questions removed from this analysis:
+Social Media Q1: How can ServiceX use photographs you post for advertisements or promotions?
+Original Answer: ServiceX can use photos you post in ads without your permission and is not obligated to attribute you as the creator.
+Relevant Clause: “…you grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content and any name, username, voice, or likeness provided in connection with Your Content in all media formats and channels now known or later developed anywhere in the world…you irrevocably waive any claims and assertions of moral rights or attribution with respect to Your Content.” (User Agreement)
+Possible Room for Interpretation: Though the original answer best matches the relevant clause.
+The clause does not explicitly mention advertisement.
+This may leave room for interpretation on whether the clause applies to ads.
+We use the same Bayesian model as §F.3, except that there are 4 comprehension questions for each service that are analyzed (i.e., K=4).
+Out of the 4 questions for each service, participants on average scored 1.5 ($\sigma=1.0$) when using TermSight and 1.6 ($\sigma=1.1$) when using the baseline interface.
+Our Bayesian model analysis reveals no significant differences in the comprehension scores across interface conditions.
+The posterior distribution of the contrast between treatment and control on comprehension is centered around zero and overlaps with ROPE.
+When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair, suggesting no significant differences for every question.
+These findings matches with our original findings.
+The figures for this additional analysis can be found in the supplemental materials (e.g., forest plots).
+Figure 14.
+The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions.
+Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
+The inner, thicker line represents the 50% HPDI.
+The results show a significant effect of the TermSight interface on the user experience for every question.
+Since the 94% HPDI for each pair (control, treatment) for every question does not overlap with each other, we should expect a significant effect of the treatment on the user experience.
+E1–E7 refer to the 7 experience questions specified in Appendix E.5.
+The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the 7 experience questions.
+Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
+The inner, thicker line represents the 50% HPDI.
+There are a total of 14 lines (7 pairs of 2) in the plot.
+The results show a significant effect of the TermSight interface on the user experience for every question.
+Since the 94% HPDI for each pair (control, treatment) for every question does not overlap with each other, we should expect a significant effect of the treatment on the user experience.
+The 7 experience questions are specified in Appendix E.5.
+Figure 15.
+The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the comprehension questions.
+The $x$-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05).
+Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
+The inner, thicker line represents the 50% HPDI.
+The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap.
+This suggests no significant effect of the treatment on the user’s comprehension.
+Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
+The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the comprehension questions.
+There are a total of 24 lines (12 pairs of 2).
+The x-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05).
+Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
+The inner, thicker line represents the 50% HPDI.
+The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap.
+This suggests no significant effect of the treatment on the user’s comprehension for every question.
+Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
+Figure 16.
+The figure shows the forest plots of coefficients in the model for user experience, comprehension, and recall with respect to the ordering.
+Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
+The inner, thicker line represents the 50% HPDI.
+Since the 94% HPDI for each ordering pair (first, second) overlaps, there is no significant effect of ordering on user experience, comprehension, and recall.
+The figure shows the forest plots of coefficients in the model for user experience, comprehension, and recall with respect to the ordering.
+Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
+The inner, thicker line represents the 50% HPDI.
+Since the 94% HPDI for each ordering pair (first, second) overlaps, there is no significant effect of ordering for user experience, comprehension, and recall.
+Prompt for Obtaining Summary Snippets and Information Snippets
+Summarize the input section of the Terms of Service into concise bullet points (less than 12 words) in plain language.
+When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point.
+For each bullet point summary, include the full-text reference to the original passage in {} and don’t use ”…” to reduce text in the reference.
+When outputting a reference, don’t change anything in the original text, such as spaces and newlines.
+There can be multiple sentences or paragraphs that reference a single summary.
+The references to summary should cover the original text.
+Example output format: {EXAMPLE OUTPUT} Input: {INPUT TEXT CHUNK}
+Figure 17. Prompt for obtaining Summary Snippets and Information Snippets.
+Prompt for obtaining Summary Snippets and Information Snippets.
+Summarize the input section of the terms of service into concise bullet points (less than 12 words) in plain language.
+When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point.
+For each bullet point summary, include the full-text reference to the original passage in and don’t use ”…” to reduce text in the reference.
+When outputting a reference, don’t change anything in the original text, such as spaces and newlines.
+There can be multiple sentences or paragraphs that reference a single summary.
+The references to the summary should cover the original text.
+Example output format: EXAMPLE OUTPUT Input: INPUT TEXT CHUNK.
+Prompt for Classifying Power
+Classify the input term from a Terms of Service agreement based on the power relationship and benefit between the service and the user. Use the following categories (Service, Neutral, User):
+- Service: The term grants the service provider disproportionate power or control over the user.
+It may impose unfair restrictions, obligations, or liabilities on the user, or reduce the user’s rights and autonomy over their data or content.
+- Neutral: The term outlines standard procedures, responsibilities, or conditions the user and service have.
+For example, users take responsibility for the content they post.
+It neither significantly favors the service provider nor the user, and does not substantially impact the user’s rights.
+- User: The term empowers the user by offering clear protections, rights, or benefits, ensuring transparency, and limiting the service provider’s power.
+Examples for each category: Service:
+- The service can delete specific content without prior notice and without a reason.
+- The service can license user content to third parties.
+- The service tracks your personal data for advertising
+Neutral:
+- Users are responsible for the content they post
+- Users agree not to use the service for illegal purposes
+- Blocking first-party cookies may limit your ability to use the service
+User:
+- You can opt out of targeted advertising
+- The service does not sell your personal data
+- The service will not allow third parties to access your personal information without a legal basis
+Output format in JSON: {”Category”: ”Service/Neutral/User”, ”Explanation”: ”explanation of output” }
+Input: {INPUT INFORMATION SNIPPET}
+Figure 18. Prompt for classifying the power balance of the Information Snippets.
+[…]
+Prompt for Classifying Relevance
+For the input term from a Terms of Service, output a relevance rating (High/Low) of the input term with respect to the user persona.
+[High]: The term is directly relevant to the user’s usage of the service or what the user cares about. The term applies to the user persona and is necessary for the user to know.
+[Low]: The term is not relevant to the user’s usage of the service or what the user cares about. The term doesn’t apply to the user persona or is not necessary for the user to know.
+User Persona: {INPUT USER PERSONA}
+Output format in JSON: {”Relevance”: ”Low/High”, ”Explanation”: ”explanation of output” }
+Input: {INPUT INFORMATION SNIPPET}
+Figure 19. Prompt for classifying the relevance of Information Snippets to the user persona.
+Prompt for classifying the relevance of Information Snippets to the user persona.
+For the input term from a Terms of Service, assign a relevance rating (High/Low) of the input term with respect to the user persona.
+Relevance rating: [High]: The term is directly relevant to the user’s usage of the service or what the user cares about.
+The term applies to the user persona and is necessary for the user to know.
+[Low]: The term is not relevant to the user’s usage of the service or what the user cares about.
+The term doesn’t apply to the user persona or is not necessary for the user to know.
+User Persona: INPUT USER PERSONA Output format in JSON: ”Relevance”: ”Low/High”, ”Explanation”: ”explanation of output” Input: INPUT INFORMATION SNIPPET
+Prompt for Identifying Unfamiliar Phrases
+You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of.
+Jargon refers to domain-specific terminologies that a lay user might not know about.
+Example jargon:
+- legal jargon: indemnity, arbitration, liability
+- copyright licenses: sublicensable licenses, royalty-free licenses
+- technical privacy terms: cookies, Ad identifiers, Authentication tokens
+Return an empty array if the section does not contain jargon. The extracted word should exactly match the original input text with the same capitalization and sequence of words.
+Output format in JSON: {”Jargon”: []}
+Input: {INPUT TEXT CHUNK}
+Figure 20. Prompt for identifying potentially unfamiliar phrases for lay users in a text chunk.
+Prompt for identifying potentially unfamiliar phrases for lay users in a text chunk.
+You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of.
+Jargon refers to domain-specific terminologies that a lay user might not know about.
+Example jargon: - legal jargon: indemnity, arbitration, liability - copyright licenses: sublicensable licenses, royalty-free licenses - technical privacy terms: cookies, Ad identifiers, Authentication tokens Return an empty array if the section does not contain jargon.
+The extracted word should exactly match the original input text with the same capitalization and sequence of words.
+Output format in JSON: ”Jargon”: [] Input: INPUT TEXT CHUNK
+Prompt for Identifying Vague Phrases
+You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service.
+A vague term refers to information that is vaguely abstracted without a clear definition provided in the section.
+Example Vague terms: information, other, some, third parties, most, generally, personal data, others, general, many, various, might, services, certain information
+Return an empty array if the section does not contain vague terms. The extracted word should exactly match the original input text with the same capitalization and sequence of words.
+Output format in JSON: {”Vague”: []}
+Input: {INPUT TEXT CHUNK}
+Figure 21. Prompt for identifying ambiguous phrases in a text chunk.
+Prompt for identifying ambiguous phrases in a text chunk.
+You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service.
+A vague term refers to information that is vaguely abstracted without a clear definition provided in the section.
+Example Vague terms: information, other, some, third parties, most, generally, personal data, others, general, many, various, might, services, certain information Return an empty array if the section does not contain vague terms.
+The extracted word should exactly match the original input text with the same capitalization and sequence of words.
+Output format in JSON: ”Vague”: [] Input: INPUT TEXT CHUNK
+Prompt for Generating Phrase Definitions
+Use information in the retrieved context to provide a definition of the user-selected phrase or term.
+Avoid using long sentences.
+For example, if the user-selected term is ”information”, define what the term ”information” includes and refers to, such as: location data, interaction data, profile data, etc.
+The output definition should be specific and straight to the point; don’t include language that doesn’t contribute to the definition, such as ’in the given context’.
+Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”.
+If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”.
+Examples: {EXAMPLES}
+Output format in JSON: {”Definition”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”]}
+Retrieved Context: {RETRIEVED CONTEXT}
+Question: What does {INPUT PHRASE} refer to?
+Context around the user-selected phrase: {PHRASE CONTEXT}
+Figure 22. Prompt for generating in-context phrase definitions.
+Prompt for generating in context phrase definitions.
+Use information in the retrieved context to provide a definition of the user-selected phrase or term.
+Avoid using long sentences.
+For example, if the user-selected term is ”information”, define what the term ”information” includes and refers to, such as: location data, interaction data, profile data, etc.
+The output definition should be specific and straight to the point; don’t include language that doesn’t contribute to the definition, such as ’in the given context’.
+Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”.
+If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”.
+Examples: EXAMPLES Output format in JSON: ”Definition”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”] Retrieved Context: RETRIEVED CONTEXT Question: What does INPUT PHRASE refer to?
+Context around the user-selected phrase: PHRASE CONTEXT
+Prompt for Generating Scenarios
+Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase.
+The scenario/example should be relevant to the below user persona using {an E-commerce platform of used items / a Social Media platform}.
+User Persona: {INPUT USER PERSONA}
+Output format in JSON: {”Story”: ””}
+User selected phrase: {INPUT PHRASE}
+Context around the user-selected phrase: {PHRASE CONTEXT}
+Definition of user selected phrase: {GENERATED DEFINITION}
+Figure 23. Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase.
+Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase.
+Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase.
+The scenario/example should be relevant to the below user persona using an E-commerce platform of used items / a Social Media platform.
+User Persona: INPUT USER PERSONA Output format in JSON: ”Story”: ”” User selected phrase: INPUT PHRASE Context around the user-selected phrase: PHRASE CONTEXT Definition of user selected phrase: GENERATED DEFINITION
+Prompt for Generating Answers to User Questions.
+You are an assistant for question-answering tasks.
+Use information in the retrieved context to answer the user’s question in less than 5 sentences.
+Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”.
+If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”.
+Examples: {EXAMPLES}
+Output format in JSON: {”Answer”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”]}
+Retrieved Context: {RETRIEVED CONTEXT}
+Question: {USER QUESTION}
+User selected phrase: {INPUT PHRASE}
+Context around the user-selected phrase: {PHRASE CONTEXT}
+Figure 24. Prompt for generating answers to user questions.
+Prompt for generating answers to user questions.
+You are an assistant for question-answering tasks.
+Use information in the retrieved context to answer the user’s question in less than 5 sentences.
+Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”.
+If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”.
+Examples: EXAMPLES Output format in JSON: ”Answer”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”] Retrieved Context: RETRIEVED CONTEXT Question: USER QUESTION User selected phrase: INPUT PHRASE Context around the user-selected phrase: PHRASE CONTEXT
 Phrase Type Original Passage Generated Definition Generated Scenario
@@
 Table 2. Examples of the generated definitions and scenarios for potentially unfamiliar and vague phrases.
+Four examples of the Phrase Scope’s generated definitions and scenarios for potentially unfamiliar and vague phrases.
+The four columns describe the type of phrase (Jargon or Vague phrase), the original passage that contains the phrase, the generated definition, and the generated scenario.
 Figure 25.
-The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions.
-Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
-The inner, thicker line represents the 50% HPDI.
-The results show a significant effect of the TermSight interface on the user experience for every question.
-Since the 94% HPDI for each pair (control, treatment) for every question do not overlap with each other, we should expect a significant effect of the treatment on the user experience.
-E1–E7 refer to the 7 experience questions specified in Appendix F.3.
-Figure 20.
-The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions.
-The $x$-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05).
-Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient.
-The inner, thicker line represents the 50% HPDI.
-The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap.
-This suggests no significant effect of the treatment on the user’s comprehension.
-Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
-Prompt for Obtaining Summary Snippets and Information Snippets Summarize the input section of the terms of service into concise bullet points (less than 12 words) in plain language.
-When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point.
-For each bullet point summary, include the full-text reference to the original passage in {} and don’t use "..." to reduce text in the reference.
-When outputting a reference, don’t change anything in the original text such as spaces and newlines.
-There can be multiple sentence or paragraph references to a single summary.
-The references to summary should cover the original text
-Example output format: {EXAMPLE OUTPUT}
-Input: {INPUT TEXT CHUNK}
-Figure 21. Prompt for obtaining Summary Snippets and Information Snippets
-Prompt for Classifying Power Classify the input term from a Terms of Service agreement based on the power relationship and benefit between the service and the user.
-Use the following categories (Bad, Neutral, Good):
-- Service: The term grants the service provider disproportionate power or control over the user.
-It may impose unfair restrictions, obligations, or liabilities on the user, or reduce the user’s rights and autonomy over their data or content.
-- Neutral: The term outlines standard procedures, responsibilities, or conditions the user and service have.
-For example, users take responsibility for the content they post.
-It neither significantly favors the service provider nor the user, and does not substantially impact the user’s rights.
-- User: The term empowers the user by offering clear protections, rights, or benefits, ensuring transparency, and limiting the service provider’s power.
-Examples for each category: Service:
-- The service can delete specific content without prior notice and without a reason.
-- The service can license user content to third parties.
-- The service tracks your personal data for advertising
-Neutral:
-- Users are responsible for the content they post
-- Users agree not to use the service for illegal purposes
-- Blocking first party cookies may limit your ability to use the service
-User:
-- You can opt out of targeted advertising
-- The service does not sell your personal data
-- The service will not allow third parties to access your personal information without a legal basis
-Output format in JSON: {"Category": "Service/Neutral/User", "Explanation": "explanation of output" }
-Input: {INPUT INFORMATION SNIPPET}
-Figure 22. Prompt for classifying power balance of the Information Snippets
-Prompt for Classifying Relevance For the input term from a Terms of Service, assign a relevance rating (High/Low) of the input term with respect to the user persona.
-Output "Low" for low-relevance terms.
-Relevance rating:
-[High]: The term is directly relevant to the user’s usage of the service or what the user cares about. The term applies to the user persona and is necessary for the user to know.
-[Low]: The term is not relevant to the user’s usage of the service or what the user cares about. The term doesn’t apply to the user persona or is not necessary for the user to know.
-User Persona: {INPUT USER PERSONA}
-Output format in JSON: {"Relevance": "Low/High", "Explanation": "explanation of output" }
-Input: {INPUT INFORMATION SNIPPET}
-Figure 23. Prompt for classifying the relevance of Information Snippets to the user persona
-Prompt for Identifying Unfamiliar Phrases You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of.
-Jargon refers to domain-specific terminologies that a lay user might not know about.
-Example jargon:
-- legal jargon: indemnity, arbitration, liability
-- copyright licenses: sublicensable licenses, royalty-free licenses
-- technical privacy terms: cookies, Ad identifiers, Authentication tokens
-Return an empty array if the section does not contain jargon. The extracted word should exactly match the original input text with the same capitalization and sequence of words.
-Output format in JSON: {"Jargon": []}
-Input: {INPUT TEXT CHUNK}
-Figure 24. Prompt for identifying potentially unfamiliar phrases for lay users in a text chunk.
-Prompt for Identifying Vague Phrases You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service.
-A vague term refers to information that is vaguely abstracted without a clear definition provided in the section.
-Example Vague terms: information, other, some, third parties, most, generally, personal data, others, general, many, various, might, services, certain information
-Return an empty array if the section does not contain vague terms. The extracted word should exactly match the original input text with the same capitalization and sequence of words.
-Output format in JSON: {"Vague": []}
-Input: {INPUT TEXT CHUNK}
-Figure 25. Prompt for identifying vague phrases in a text chunk.
-Prompt for Generating Phrase Definitions Use information in the retrieved context to provide a definition of the user-selected phrase or term.
-Avoid using long sentences.
-For example, if the user-selected term is "information", define what the term "information" includes and refers to, such as: location data, interaction data, profile data, etc.
-The output definition should be specific and straight to the point, don’t include language that doesn’t contribute to the definition such as ’in the given context’.
-Output the string list of reference ids (["ref1", ...]) used to generate the definition under "References".
-If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for "References".
-Examples: {EXAMPLES}
-Output format in JSON: {"Definition": "", "References": ["ref1", "ref2", "ref3"]}
-Retrieved Context: {RETRIEVED CONTEXT}
-Question: What does {INPUT PHRASE} refers to?
-Context around the user-selected phrase: {PHRASE CONTEXT}
-Figure 26. Prompt for generating in context phrase definitions
-Prompt for Generating Scenarios Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase.
-The scenario/example should be relevant to the below user persona using {an E-commerce platform of used items / a Social Media platform}.
-User Persona: {INPUT USER PERSONA}
-Output format in JSON: {"Story": ""}
-User selected phrase: {INPUT PHRASE}
-Context around the user-selected phrase: {PHRASE CONTEXT}
-Definition of user selected phrase: {GENERATED DEFINITION}
-Figure 27. Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase
-Prompt for Generating Answers to User Questions You are an assistant for question-answering tasks.
-Use information in the retrieved context to answer the user’s question in less than 5 sentences.
-Output the string list of reference ids (["ref1", ...]) used to generate the definition under "References".
-If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for "References".
-Examples: {EXAMPLES}
-Output format in JSON: {"Answer": "", "References": ["ref1", "ref2", "ref3"]}
-Retrieved Context: {RETRIEVED CONTEXT}
-Question: {USER QUESTION}
-User selected phrase: {INPUT PHRASE}
-Context around the user-selected phrase: {PHRASE CONTEXT}
-Figure 28. Prompt for generating answers to user questions
+The baseline reading interface.
+Participants can navigate to different policies using the top navigational panel.
+In place of the Summary Snippets on the left is a table of contents.
+Participants can click a section header in the table of contents to navigate to the corresponding section.
+The baseline system used in the user study.
+The interface contains three panels: a top navigation panel for switching between different policies, the original text on the right, and the table of contents on the left.
+Participants could click a section header in the table of contents to navigate to the corresponding section.